A Complete Bibliography of Publications in HOPOS: Journal of the International Society

for the History of Philosophy of Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

26 November 2024 Version 1.28

Title word cross-reference

$105.00 [Cre20, Hat20b, MP18, Ver20b]. $109.00 [Ste19b]. $109.99 [Sin19]. $110.00 [Lav20, Pea19b]. $140.00 [Del19]. $149.00 [Cap20]. $155.00 [Dor20]. $175.00 [Rag19]. $180.00 [Del19]. $202.00 [Gre20]. $25.00 [Cam20]. $250.00 [Roq19]. $27.92 [Bsc19]. $318.00 [Pal15]. $35.00 [MP18]. $40.00 [Rus19]. $45.00 [Gro20, Ste19a]. $47.96 [BKV^{+}22]. $48.00 [Rap20]. $49.95 [Lan19, Pea19a]. $50.00 [Geo19]. $51.95 [Rag19]. $54.95 [Büt20]. $58.00 [Roq19]. $65.00 [Bal19]. $75.00 [Cam20]. $84.99 [Sin19]. $85.00 [Bal19, Dun24, Lav20, Ste19b]. $89.95 [Mil20]. $95.00 [Tub20]. $99.00 [Rus19]. < [Fee18]. > [Fee18].

0 [Büt20, Dor20, Hat20b, Pea19a]. 0-8229-4522-3 [Pea19a].

1 [Cre20, Dun16]. 1662 [And15]. 1685 [Jes11]. 1689 [McC15]. 1700 [Lær18]. 1760 [Jes11]. 1820s [McC24]. 1830s [Ver24]. 1850s [Ver24]. 18th [Pal15, Zam24b, Hol13]. 1913 [Atk24]. 1919 [Car24]. 1930 [Fee18]. 1930s [McC24]. 1935 [Car24]. 1950s [Ric12]. 1960s [Chi08, Ric12, Dro11]. 1980s [PS23].

2 [Dun17, Geo19, Lav20, Ver20b]. 2019 [Kis24].

3 [Cap20, Dob20a, Gre20, Pea19a, Tub20].

4 [Mil20, Rap20]. 46 [Bre14]. 46-51 [Bre14].

5 [BKV^{+}22, Gro20]. 51 [Bre14]. 52a [dCH11].

6. [Cam20].

7 [DP19].

8 [Dik22, Dun24].

978 [BKV^{+}22, Büt20, Cam20, Cap20, Cre20, DP19, Dob20a, Dor20, Dun24, Geo19, Gre20, Gro20, Hat20b, Lav20, Mil20, Rap20, Tub20, Ver20b]. 978-0-19-060804-0 [Hat20b]. 978-0-19-284677-8 [Dun24]. 978-0-226-08437-4 [Rap20]. 978-0-226-44998-2 [Geo19]. 978-0-226-47478-6. [Cam20]. 978-0-691-17171-5 [Gro20]. 978-0-812-25066-4 [Mil20]. 978-0-815-37349-0 [Rag19]. 978-0-815-37985-0 [Dor20]. 978-1-032-02036-5 [BKV^{+}22]. 978-1-107-15250-2 [Ver20b]. 978-1-107-17815-1 [Cre20]. 978-1-137-47250-2 [Lav20]. 978-1-138-50313-7 [DP19]. 978-1-138-90089-9 [Rag19]. 978-1-4214-2178-0 [Büt20]. 978-14-38-47367-3 [Tub20]. 978-3-030-02128-3 [Cap20]. 978-3-110-56782-3 [Dob20a]. 978-3-319-96046-3 [Bal19]. 978-3-319-96047-0 [Bal19]. 978-90-04-32591-3 [Gre20].

= [Rig66].

A. [Bok15, Doy24]. Aaron [Kor18]. Abduction [Plu11]. Abel [Bre21b]. Absence [PG19]. Absolute [Bre21b]. Abstraction [Chi20, CP21]. Academic [Lan19, Mai14, Mih24, And15]. according [Bei22]. Account [Det18]. Accounts [Vat18]. Accusation [Ade15]. Achinstein [Dun14]. across [CD13, Gar15b]. Action [DS21, Pal15, Hol13]. Active [Rus21, SDR^{+}21]. Ad [Duc14]. Adam [AES^{+}24, Cam20, Cap20, Lau24]. Adamson [Koe24]. Add [Kri12]. Adrian [Ryn14]. Adriana [Pod14]. Advancement [Ada11, Poo10]. after [Rus09]. Again [Dom13]. Against [Dun16, Ant22, Bor11, Toa13, vS22]. Age [Lær18, Lev15, Mag12, Pea19b, McE18, Nic11]. agency [Fis17, Bsc19]. Aït [DeS12]. Aït-Touati [DeS12]. al [Çim19, IW18, Koe24]. al-Bitruji [Çim19]. al-Haytham [IW18]. Al-Razi [Koe24]. Alan [Doy24, BV11, Nas16]. Albany [Tub20]. Alex [Pod14]. Alexander [Hof13, Mun12]. Alfred [Hat12a, Kuu13]. Algebra [Toa21, Wüt14, Cao10]. Alisa [Hep11]. Allan [Boy13, ST16]. Allhoff [Mir11]. Almost [Rei14a, EKD^{+}13]. Alois [Bia14, Mee24]. Alspector [Mir11]. Alspector-Kelly [Mir11]. Alternative [Iva17, Stu15a]. American [Kas24, Leg24, Pea14, Rei14b, Hat15a, HMJ15, Jew12, Kle15a, KK60, Mis13, Mor17, Neu20, Ric12, Tub21, Ver20a]. among [Ari11, Gre91, Hon18, Roz13]. Amos [Ste19b]. Anaclastic [Dik22]. Analogy [Vat18]. Analysis [Gav24, Jal24, Ric10, Kob14]. Analytic [Gav24, Kis24, Lau24, Pre17, Red24, Ueb21a, Kor18]. Analytical [Esa19, Dun24]. Analyticity [Mat24]. analytics [Bro16, Gol17]. Anastasios [CL12, dP16]. Anatomy [New14, WG12, Koo10, Wax13]. Anaxagoras [Mil14a, Gra13]. Ancient [Lev15, Meh14, Sch14a, Bet10, KM11, Lær18]. Anderson [Gie24]. André [Fal18, Mos14, Pal15]. Andrea [Dob20a, Zam18a]. Andrew [Kin16, GBV^{+}22, Rei14b]. Angélique [Ara19]. Angus [FS12]. Anik [Car16a]. Animal [Gib15, Wil20, Lov16]. Animals [De 19]. Animate [Sha21]. Animation [Omo21, Rut21]. Animi [Jon13, Cor11]. Animism [Pec21, Sha21, dC21]. Anne [Rus21]. Annibale [Ela13]. Annie [Bre18]. Annus [Feh21]. Anomaly [Bsc15]. Anschluss [Ryc13, Sie10]. Anstey [Hil12]. Answer [Red24, Bru05]. anthology [MAKA09, Mir11]. Anthropology [Ste20]. Anti [Joh10, Gie12]. Anti-Darwinism [Joh10, Gie12]. Anticipated [Red24]. Antimetaphysics [Ueb11]. Antinomy [Bei22, How22]. Antipsychologism [Edg20a]. Antiquaria [Feo11, Tos09]. antiques [Tos09]. antiquity [Ger10, Rem13]. Antirealism [Wra21]. Antoine [Mel17]. Antoine-Mahut [Mel17]. Apostle [Mur24]. Apperception [Hat20a]. Appraisal [ES17, Fal21, Tur24]. Appreciating [BS22]. Arabatzis [Obe14]. Arabic [dCH11]. Aratus [Gee13, Lat14]. Archaeology [Kri12]. Archeus [Dem21]. archives [Tos09]. archivi [Feo11, Tos09]. Arena [Rom12]. Argentina [LNN12, Pod14]. Argument [BKV^{+}22, Ngu23, SDR^{+}21]. Ariew [Det16, Roz13]. Aristides [Bog13]. Aristotelian [McC15, Mon15, Mos14, Blo12, God10, Omo21, Sga12, Sga17]. Aristotelis [Pic16, Zup12, SBT10]. Aristotle [Bal19, Boy13, Cro16, Fal17, Gol16, Hat15b, Pop16, Bro16, BZ14, Con24a, De 14, De 19, DeW17a, Ebr15, Gab24, Gil14, Gol17, GLB12, GS17, Len11, Len14, Mar14, McC24]. Arithmetic [Dun16, Dun17]. Arne [Rad13]. Art [Are12, Hol10, RH10, Bee24]. Arthur [Don24, Slo24, Stu11b]. Articles [Don24]. Articulation [Kin18b]. Artificial [Sim11, BV07]. Arts [FMB12, Pow12, Edd15, Row15]. Asian [Mil15a]. assayer [GD90]. Assessment [Bia16]. Association [Hat20a]. Astrology [Ito14, Bon13]. Astronomical [Çim19, Lat14, Gee13]. Astronomy [Fol22, Lat11, ABS10, Gra13, Mil14a]. Atkins [Vio24]. Atlantic [Irv12]. Atoms [Eat21, Hud20]. Atti [Rig66]. Aubin [Lat11]. Aubrey [Ada11, Poo10]. Aufbau} [Dew17b]. Auguste [Pea19a, Bou13, Bre14, Esa19, Gui14]. Aurelia [Don24]. Austro [Mor17].

Author [FA11]. Authors [SB20]. Autonomy [Chi20]. autres [Bre15]. Aux [Ler10a, Nem12]. Averroe [Ant11]. Averroes [Gla09, Çim19]. Avicenna [Zam24a]. Avogadro [Cok20]. Axiom [The20]. Axiomatics [Toa21].

B [Gor21, HES^{+}24, Lag14, Tub20, Wra16]. BAAS [BS22]. Bachelard [Ler11, Fab24, Ter10]. Back [Dom13]. Bacon [Ans15, Çim19, Sch17b]. Baconianism [Ver24]. Bailer [Ste13]. Bailer-Jones [Ste13]. Bakker [Zup12]. Ball [Fut14]. Baltas [Bog13]. Baltimore [Büt20]. Banchetti [Eat21]. Banchetti-Robino [Eat21]. Bandwagon [Ant22]. Banks [Lan16]. Bardon [Ryn14]. Baroque [GCM12, Lüt14]. Barrett [Bac13]. Barthes [Ray14]. Bartley [Sep24, Sha20]. Basingstoke [Sin19]. Basis [Sha20]. BC [DeW17a]. Be [Wue13]. Became [Fut14, Bal13]. Beckwith [Mil15a]. Bede [KW10, Lag14]. Beeckman [Art14, vB13]. Been [Kin16, PS23, STP15]. before [Gra13, Mil14a]. beliefs [LM15, Mun16]. Believe [Kle15b, Mis15]. Bella [Hat20b]. Bellis [Bre24]. Benjamin [Mar17, Roq19]. Bensaude [Sim11]. Bensaude-Vincent [Sim11]. Berg [Gam15]. Bergson [Sau17, Can16, Rat24]. Berkel [Art14]. Berlin [Ric15, MP13]. Bernadette [Sim11]. Bernard [Duc17]. Bernays [Ede24]. Bett [Meh14]. Better [BS22]. Between [Gre15, Mar10, McC24, Rot24, Tim19, Coz24, Dam14, FPS11, Gio17, LM15, Mil20, Mun16, Mur24, SB20, Vra20, Mun12]. Beyond [Bok08, San16, Ueb11, Omo21]. Biener [Cal15]. Bigg [Lat11]. Bilder [Ueb16]. Bildpädagogik [Ara19, Gro15]. Biographies [Stu16]. Biological [Ken18]. Biologico [Bre21a]. Biologico-economical [Bre21a]. Biology [AN10, Chi20, Coz24, Gab24, Gam15, GW14, Len14, Man15, Rat24, Ste19a, Zam18a, GLB12, Van13, Zam18b, Hof13, Boy13, Men15]. Birth [Bre24, Dob20a, Fra22]. Bitruji [Çim19]. Black [Tim19]. Black-Body [Tim19]. Bloch [Mon15]. Bloomsbury [Lan19]. Bodies [Nob19]. Body [Pur15, RES^{+}24, Smi11a, Tim19, Gar09]. Boerhaave [Edd15, Pow12]. Böhlandt [Kat12]. Bohnert [Mar24]. Bois [HES^{+}24]. Bokulich [Hep11]. Boner [Ito14, Rot17]. Bonnet [Dun24]. Book [ASF^{+}21, Ada11, And15, AES^{+}24, Ant11, Ant13, Ara19, Are12, Ari17, Art14, Atk24, Bac13, Bal19, Bee24, BKV^{+}22, BM12, Bel13, Ben13, BV11, Bia16, Bla12, Bog13, Bok15, Bon16, Boy13, Bre14, Bre18, Bre24, Bro24, BS13, Bsc19, Büt20, Cal15, Cam20, Cap20, Car16a, Car16b, Car24, CL12, Cat12, Con24a, Con24b, Cow12, Cre20, Cro16, Dam17, DeS12, DeS14, DeW17a, Del19, DP19, Det15, Det16, Dob20a, Dom11, Don24, Dor20, Dow17, Doy24, DL18, Dro11, Duc12, Duc14, Duf11, Dun24, Dun11, Dun14, Ear13, Eat21, Eck18, Edd15, Ekh13, Ela13, Fal17, Fal18, Fal16, Fee18, Feo11, FES^{+}23, Fog12, Fra13, FS12, Fut14, Gab24, Gam15, Gar15a, Gar24, Gar15b, Gaw11, Geo19, Gie12, Gie24, Gig16, Gol11, Gol16, Gol17, Gor11]. Book [GBV^{+}22, Gre13, Gre15, Gre20, Gri13, Gro20, Gui18, Gui14, Haa13, Ham24, Har14, Hat12a, Hat15b, Hat20b, Hen13, Hen24, Hep11, Hic12, Hil12, Hof13, Hoq12, HES^{+}24, Ito14, Iva17, Jes11, Jha12, Jon13, Kas24, Kat12, Ken18, Ken22, Kin18a, Kin16, Kis24, Kjæ11, Kle12, Kob14, Kor18, Kra22b, Kuu13, Lær18, Lag14, Lan19, Lan16, Lat11, Lat14, Lau24, Lav20, Leg24, Ler11, Lol15, Lon23, Loo11, Lov16, Lüt14, Lut24, Mag12, Mar17, McC15, McC24, McD14, Meh14, Mel17, Men15, Mil14a, Mil14b, Mil20, Mil15a, Mir17, Mir11, Mon15, MP18, Mor24, Mos14, Mun12, Mun16, Nas16, Nem12, New14, Nic12, Nor18, Obe14, Pad11, Par11, Pea14, Pea19a, Pea19b, Pen13, Pen17, Pic16, Pod14, Pop16, Pra13, Pur15, Rag19, Rap20, Rat24]. Book [Rei14b, Rei14a, Rem13, Ren18, Rey15, Ric13, Ric15, Rod24, Rog15, Roq19, Ros13, Rot17, Row15, Roz13, Rus19, RES^{+}24, Ryc13, Ryn14, Sal12, San22, Sar17a, Sar17b, Sau17, Sch14a, Sch13, Sch21, Seg14, Ser21, She11, SR18, Sim11, Sin19, Sin22, Slo11a, Slo24, Sme14, Smi11a, Smi24, Sna11, Spi18, ST16, Sta11, Ste19a, Ste19b, Ste13, Str18, Stu11c, Tab11, Teu15, Tub20, Tur24, Tur11, Ueb16, Ueb18, Ueb21a, Van11, Ver20b, Vio24, WG12, Wüt14, Zam24a, Zam18a, Zam24b, Zup12, dP16, vS15]. Boran [Gui18]. Bordogna [Kle12]. Bordoni [Sch21]. Borghero [Bel13]. Born [Str18]. Borrello [Ear13]. Bostanci [Cam20]. Boston [Cap20, Dob20a]. Boundaries [Kle12, Bor08]. Boundary [How22]. Bourdeau [Bre14, Gui14, Pea19a]. Bourdieu [Tur24]. Boyle [Eat21, Jon13, Cor11, Har22]. Brad [San22]. Breaking [Gro20, Mih24]. Brenner [CL12, dP16]. Brentano [Fré20]. Brethren [Ant13]. Brevis [Ado19]. Brief [Plu11, Ryn14, Bar13]. Brill [Gre20, Pal15]. Britain [Mih24]. British [McC15, Mih24, Sga12]. Bronstein [Gol17]. Brown [HES^{+}24, Row15]. Bruno [Ant13]. Brunschvicg [Bre15]. Bschir [Ken22]. Buchdahl [Vat18]. Buchwald [Sme14]. Buridan [Zup12, SBT10]. Business [Dom13, Duc11]. Byrne [Bac13].

C [Car16b, Edd15, Fal17, Lan16, Lev15, Mar17, MP18, Pea19b, Rus19, Sin19, Tes16, Lær18]. C. [McE18, OS17, Sin22]. Cage [Lon23]. Callataÿ [Ant13]. Called [MCP19]. Calvin [Lag14]. Cambridge [Cre20, Meh14, Rem13, Tab11, Ver20b, Bet10, Ger10, HR02]. Can [De 24]. Canales [Sau17]. Canonical [SB20]. Canonizing [Pet18]. can’t [Bru05]. Cantor [Red24]. Cao [Wüt14]. Cardano [Reg21]. Cárdenas [Kas24]. Cardinality [TF24]. Career [Rei14a, EKD^{+}13]. Carl [Sin19, Gar15b]. Carla [Bre24]. Carlo [Bel13]. Carlos [Gre15]. Carnap [Car24, Doy24, Ren18, Smi24, Ant22, Cre20, Dew17b, Kub18, Ueb21b, FA11]. Carnapian [Mar24]. Cartesian [ADJ^{+}15, Det16, DNB13, Gar15a, Ros22]. Cartesianism [Dob20a]. Cartesianisms [Spi18, Sch16]. Cartesians [Bor11]. Cartésiens [Bel13, Bor11]. Carus [Doy24]. Case [DS21, Ela13, ER17, Esa19, Rus19, Wüt14, Cao10, Fan12]. Cases [Cow12, CP21, Gim11]. Cassirer [Bia16, Chi20, Gio24, Hei14]. Castelão [Ler11]. Cat [Cap20]. categories [Vin15, Car16b]. Catherine [Bla12, DL18, Zam24b]. Caught [Dam14]. Causal [Koc13, Vir22, Wes15]. Causality [Fal16]. Causation [Lol15, Ott09, Sch07, Sch14b, Gor11, Slo11a]. Cause [Ado19, De 19]. Causes [Esa19, Sch10]. Cavaillès [Cau20, Cau24]. Cavendish [Gol11, Pet18, Sar10, Sha21]. Cells [Gie24]. Centenary [BS22]. Central [Mil15a, Pal15]. Century [And15, BBK^{+}16, BM16, Bro24, CP21, DeW17a, FS12, Hat12a, Hoq12, LH16, Lat11, Loo11, Lov16, Pen18, Pod14, Roq19, Sch24, Zam24b, Gib15, LNN12, Mai14, ATE11, ABS10, Bor11, FN06, Hol13, Jes14, NL12, Pal15, Rap20, Rhe10, SRK09, Tos09, Vio13, DeS12, Mar17, Sch21, ST16]. Cercle [Nem12, Ler10a]. Certain [AP22]. Certainty [Del17, Dun11, Rot24, Sch17b, Gui09]. Châtelet [Wel23]. Challenge [Sch11a]. Challenges [Agi24]. Cham [Ara19, Bal19, Cap20, Del19, Ste19b]. Change [Fra12, Iva17, Lun10, Stu15a, Nic12]. Changed [Can16, Sau17]. Changing [Dom11, MM09]. Chapter [Buz21, D’A24]. Charge [Kin18b]. Charity [Irv12]. Charles [Atk24, Bro24, Ari17, Dun24, Pea18, Pen18, Sch17a]. Charlotte [Lat11]. Charronian [And15, Mai14]. Chelsea [Fal17]. Chemical [Eat21, Edd15, Pow12]. Chemistry [Edd15, McN16, Wra19, Pow12]. Chemla [Eck18]. Chen [Lüt14]. Chen-Morris [Lüt14]. Chene [Det16]. Cheryl [Pea14]. Chicago [Bsc19, Cam20, Geo19, Rap20, Ste19a]. Chimisso [ASF^{+}21, Dro11]. choice [Bre11, The20]. choix [Bre11]. Chorlay [Eck18]. Christa [Del19]. Christiaan [Mar21]. Christian [Car24, Fee18, Irv12, Mor24]. Christina [GD90]. Christoph [Lut24, She11, DeS14, Ueb16]. Christopher [Kas24, Bro24, Mil15a]. Chromodynamics [Wüt14, Cao10]. Chronos [Fal17]. Chunglin [Sch14a]. Chymical [Eat21]. Circle [Cat12, Haa13, Ueb18, Fer24, Ler10a, Man09, Rom12, Ueb07, vS22]. Civil [Rei14b, Jew12]. civilization [BF12, Sme14]. Claire [Lon23]. Clarke [Ngu23]. classic [RD16, Rus09, Dam17]. Classical [Bok08, Hep11]. Classification [DL18, Ken16]. Clauzade [Bre14]. Clerk [Ach13]. Climate [Wue13]. Cloisters [Mil15a]. Closed [Ela13, Fan12]. Closer [Pec21]. Closing [BBK^{+}16]. cloth [Ara19, Bal19, BKV^{+}22, Büt20, Cam20, Cap20, Cre20, Del19, DP19, Dob20a, Dor20, Dun24, Geo19, Gre20, Gro20, Hat20b, Lav20, Mil20, MP18, Pal15, Pea19a, Pea19b, Rag19, Rap20, Roq19, Rus19, Sin19, Ste19a, Ste19b, Tub20, Ver20b]. Cluster [War24]. Cogito [Sid24]. Cognition [Hat20a]. Cognitive [Ueb11]. Cohen [Bia14, Edg20b, Edg20a]. Cold [Rei14b, Rei14a, Tub20, EKD^{+}13, Jew12]. Collaborative [Agi24]. Collapse [Jes11, Gau10]. Collected [Bac13, BB12, Doy24]. collecting [Tos09]. collezionismo [Feo11, Tos09]. Collision [Gio17]. Color [HES^{+}24]. Combinatorial [Bee24]. Come [Bre21b]. Commentary [BB12, Bac13]. Common [Mur24]. Community [Vio24]. Companion [LH16, Meh14, Roq19, Tab11, Bet10, HR02]. Company [Atk24]. comparative [Ler10a, Ler10b]. comparée [Ler10a, Ler10b]. Compatibility [Bsc15]. Complementarity [Fal16]. Complexity [Len14, Str18]. Computational [MCP19]. Comte [Bre14, Bre18, Gui14, Bou13, Bre21b, Esa19, Pea19a, Pet16]. Conant [Tub20, Wra16]. Concept [Cau20, Fal18, Pad11, Rhe24, Sch14b, Toa13, Rei06, REG08]. Conception [Rom12, TF24]. Concepts [Bal19, Fra12]. Conceptual [De 24, Gav24, Stu15a, Iva17]. Concrete [Hoq12, Rhe10]. Conference [Kis24]. Confirmation [Hun14]. Confronts [CJ21]. Congestions [Rad13]. connection [Sie10]. Consequences [Llo13]. Constitution [McC22]. Constitutional [Llo13]. Constitutivity [Fra12]. Constraints [Gre21]. Construction [Fab24, Jha12, Pod14, LNN12, Nye11]. Constructions [Sch16, Spi18]. Constructive [Dou24]. Constructivism [Sie10]. Contemporaries [Iva20, McE18, Pea19b]. Contemporary [OS17, Sin19, Wra21, Rus09, Car16a]. Contentious [Ear13, Bor10]. Context [Tub18, Tub21]. Contexts [DFK20, Kra22a, IR11, BS13]. Contextualizing [Ngu23]. Contingency [Kin16, STP15]. Contingency/Inevitability [Kin16, STP15]. Continuum [Red24]. contra [Wel23]. Contract [Sch11b]. Contrasting [Rut21]. Contrastive [Hun14]. Contribution [Gor21, Car18, Ren18]. Controversy [Rut21]. convegno [Rig66]. Convention [Jal24]. Conventionalist [Due23]. Conventions [Sch17a]. Converse [Bog13, Bal12]. Convert [Iva13]. Conway [Rus21]. Copernicus [Mos14, God10, Mos14, Par11, Wei08]. Corneanu [Jon13]. Correspondence [Atk24, Har22]. corruption [Zup12]. corruptione [SBT10, Pic16]. Cosci [McC24]. Cosmological [Ito14, Rut21, Bon13]. Cosmology [Bei22, De 22, Fol22, Fra22, GBT17, Har11, Dom13, Rot17]. Cosmopolitics [Ste10, Fra13]. Cosmos [ATE11, DeS12, Mil20]. Cosmotheoros [Mar21]. Could [Kin16, STP15]. Counterfactuals [Gre21]. Countries [Haa13, Sch21, Man09]. Cours [Bre14, CBCD12]. Course [CBCD12]. Court [Atk24]. Craig [BM12, Hat15b]. Craver [Gar15b]. Creation [Coz24, McC24]. Creatively [Fis17, Bsc19]. Creatures [Str18]. Creed [Ham24]. Cristina [ASF^{+}21, Dro11]. critères [Bre11]. criteria [Bre11]. Criterion [Ueb11]. Critias [Gre20]. Critical [Bsc15, Due23, Gam15, Hat15a, How22, Ken22, Kle15a, Kra22b, Neu20, San22, Sep24, Teu15, Tur24, Zup12, Men13, SBT10, Ter10, Van13, dCH11]. Criticism [Dam20]. Criticisms [Fra18]. Critique [De 22, Mil15b, Red24, Sou13, Vat18]. critiques [Ler11, Ter10]. Cross [Fab24]. Cross-Perspectives [Fab24]. Crossing [Ray14]. Crossroads [Cat12, Ueb07]. Crucial [Sch17b]. Cultura [Jon13, Cor11]. Cultural [Pen13, Ste20, MWR12]. Culture [Jes11, Mun12, Ste20, ABS10, GG08, Mar10, Lat11]. Curing [Duc17]. Curiosity [Bal13, Fut14]. Current [Jes14, Wüt14, Cao10]. Cusa [Böh06]. Cusanus [Böh06, Kat12].

D [Rei14a, Sna11]. Damböck [Car24, Fee18, Mor24, Car24]. Daniel [Bre24, Kas24, Jes14, Lan19, Mil14a, Smi11a]. Daniela [Ste13]. dans [Bor11]. Darden [Gar15b]. Dare [Sch24]. Dario [Gre15, Rot17]. Darstellung [Ara19, Gro15]. Darwin [Bal19, Gol16, Par11, HR02, Hon18, Hun14, Pen18, Rus09, Wei08, Pen17, Tab11]. Darwinism [Joh10, Gie12]. Darwinistas [Pod14, LNN12]. Daryn [Str18]. Daston [Dam17, Rei14a, Sal12]. Data [Dom13, Lon13, Har11]. d’Auguste [Bre18, Pet16]. David [Eck18, Bon16, DeS14, DP19, Gol17, Gri13, Iva17, Lat11, Lüt12a, Mon15, Pop16]. Dawkins [AN10, Hof13]. Day [Bei22]. Death [Koo10, WG12]. débat [Ler10b, Nem12]. Debate [Can16, Ede24, Sau17, Ver24, Vra20, Ler10b, Ueb07, Cat12]. Debates [Ekh13, Wra21, Hir11]. Debating [Pen17]. Dedekind [Gan21, Kle18]. Deduction [Car16b, Dik22, Vin15]. Defense [Edg20b, Lut12b, Vir22]. Deficiency [McN22]. Defied [WG12, Koo10]. del [Feo11, Rig66, Tos09]. Deleuze [Duf11, Gaf10]. della [Rut21]. Delphine [Bre24, Mel17]. Demented [Ueb18]. Democracy [Rei14b, Jew12]. Demon [CJ21]. demonstratio} [Ado19]. Demonstrative [Loo11, SRK09]. Denis [Hof13]. Denken [Böh06, Kat12]. Dennis [Det16]. Deontological [Gre13, NZ10]. Dependence [Dun16]. Dependence-Hierarchy [Dun16]. Depth [Rad13]. Descarte [Gre13, Mel17]. Descartes [ADJ^{+}15, Det16, Hof13, dP24, AN10, Ari99, Ari11, Del17, Dik22, Dom11, Gar24, Gre91, Hat09, Hat12b, MM09, Man15, Mar11, NZ10, Roz13, Sch07, Sid24, Slo11a, Van11, BM12]. Descriptive [DFK20, Dam20, Fré20]. Deslauriers [Gab24]. Determination [Cok20]. Determining [Cok20]. Determinismus [FES^{+}23]. deutsche [Mor24, Rot24]. Deutscher [Fee18]. deutschsprachigen [Fee18]. Development [Fal18, Gub19, How22, Teu15, Men13]. Dewey [Bro12, Hen23]. DeWitt [Gaw11]. Dictionary [Det16, ADJ^{+}15]. Did [Bre21b, Iva13, Sep24, Toa21]. Difference [Gab24]. Dika [Gar24]. Dilthey [Dam20, Fré20, Kin18b]. Dimensional [Jal24]. Dingler [Rot24]. Dioscorides [MM23]. Dirk [Ste19b, Gie12]. Disagree [Sep24]. DiSalle [vS15]. Disciple’ [FA11]. discipline [Sep12, Gri13]. Discourse [Pur15]. Discoveries [Gal57, GD90, Gar15b, CD13]. Discovery [Mel19, Nic12, Vae21, Lun10]. Discovery-Justification [Mel19]. Discussing [Kin16, STP15]. Disrupted [Lov16, Gib15]. Dissertation [Bee24]. Distinction [Mel19, Mur24]. Divide [Ray14, SB20]. Divine [Det15, Nob19, Smi11b, Wat13, Duc12]. Dmitri [Lær18]. Do [And13]. Dobre [Gar15a]. Does [GS17, Sga17]. Dogmas [FA11]. Doing [Hud20]. Domain [Stu16]. Domski [BKV^{+}22]. Donata [FES^{+}23]. Done [Wue13]. Dongen [Fog12]. Donohue [Bro24]. Douglas [Bok15, Det16]. Drawing [Gie24]. Dubislav [Mil16]. Duchess [GD90]. Ducheyne [Dom13]. Duhem [Bre15, Buz18, Mur24]. Duncan [Con24b]. Duns [Cro16, BZ14, Cro16]. Dupin [Bre14]. Dupré [Gie24]. during [Gol11, Sar10]. Dutch [Duc14, Spi18, Dob20a, JM12, Sch16]. Dynamic [Mat24, NP22].

e-book [Bal19, Del19, Lav20, Roq19, Sin19, Ste19b]. E. [Lon23]. Early [BF17, Car16a, Dow17, Doy24, Fer22, Gig16, Gor11, Gui18, Hat15b, Hat20b, Irv12, Jon13, Loo11, Ole16, Rot17, RES^{+}24, SB20, Sar17a, Sch24, Sch16, Ueb21a, Cor11, GBT17, Mar14, NS14, Ott09, SRK09, Yeo14, Spi18, Sch10]. Earth [Lat11, Rot17, ABS10, GBT17, McM11]. East [Rag19]. Eastern [Mag12, Nic11]. Ebbinghaus [Dam20]. Ebbs [Cre20]. Ebrey [Pop16]. Eclecticism [Omo21]. Econometrics [Her19]. economical [Bre21a].

Economics [Gon14]. ed [Atk24, Cow12, Det15, Don24, Dow17, DL18, Duf11, Gaw11, GBV^{+}22, Kor18, Kra22b, Lau24, Lol15, Meh14, Pad11, Pop16, Pur15, Rem13, Rey15, Ric13, RES^{+}24, San22, Seg14, Smi24, Sta11, Tur11, dP16]. Edgar [Rom24b]. Edinburgh [Pea19b]. edited [Car24]. Edition [Zup12, SBT10, dCH11]. Editor [Ano22a, Pat17, Pat20, Pat21a, Pat21b, Pat21c, Pat24]. Editors [Ano11b]. Edmond [Buz21]. Edmund [Cau20]. eds [AES^{+}24, Ant13, Bac13, Bia16, Bre14, Bre24, Bro24, BS13, Cal15, Cap20, Car16a, Dam17, Dor20, Doy24, Duc14, Eck18, FS12, Gar15a, Gie24, Gig16, Gre15, Gui18, Haa13, Hat12a, Hat20b, Hic12, Hof13, Kas24, Ken18, Ken22, Kin16, Kis24, Lan19, Lat11, Lav20, Loo11, Lut24, Mar17, McC24, McD14, Mel17, Men15, Mir11, Mor24, Mun16, Nor18, Obe14, Pal15, Pea19a, Pic16, Rag19, Ric15, Roq19, Rot17, Row15, Rus19, Sal12, Ser21, Sim11, Sin19, Sin22, Ste19b, Stu11c, Tab11, Zup12, vS15]. Education [Mos14, God10]. Edward [Mar17]. Effect [Ado19]. Efficient [De 19, Sch14b, Lol15]. Efthymios [Mag12]. Eighteenth [Lov16, Sch24, Gib15, Bor11, Pal15, Tos09]. Eighteenth- [Sch24]. Eighteenth-Century [Lov16, Gib15]. Eino [Nem16]. Einstein [Sau17, Bok15, Can16, Fog12, Gim12, Mee24, Ros13, Sto13, vD10]. Einstein’schen [Gio24]. Eldorado [Toa21]. Elements [Hat20a]. Elizabeth [Sal12]. Elizabethanne [Gui18]. Elliott [MP18, Rus19]. Embedding [Nem16]. Emergence [Gio17, Jes11, Len14, GG08, Eat21]. Emiliano [Kin16]. Emily [Dor20]. Emma [Lat14]. Emmy [Toa21]. Empirical [Ric12, Sha20, Wil20]. Empiricism [AES^{+}24, Bia14, De 14, DeW17a, Dou24, Gio17, Hat15a, Lan16, McC15, MP13, Mor17, Mor24, Nem16, Tub18, Ueb07, Vio13, Vra20, Wes15, Ban14, BS14, Ler10b, Man09, Sga12, Cat12, Cal15, Con24a, Haa13, Lau24, Lut24, Ric15]. Empiricisms [Tub21, DNB13, Gar15a]. Empiricist [KV22]. empirisme [Ler10b]. Empirismus [Fee18, Mor24, Ueb16]. Encounters [Ver20a]. Engaged [Rom24b]. Engineering [De 24]. England [Lær18, Lev15]. English [Sar17a, Yeo14, dCH11]. Enigmatic [DeS14, Lüt12a]. Enlightenment [McD14, Meu10, Pur15]. Enough [dC21]. Entstehung [FES^{+}23]. Environment [Leg24]. Epic [Ueb18]. Epicureanism [Bla12, Wil08]. Epigenesis [Teu15, Men13]. Epistemic [BC17]. Epistemological [Ray14]. épistémologie [Bre15, dP16, Mil15b]. Epistemologist [Due23]. Epistemology [Duc17, Hoq12, McC15, Nem16, Pra13, Rhe10, Rom24a, Sch21, RF10, Sga12, Zam24a]. Epistle [dCH11]. Epistles [Ant13]. Ergo [Sid24]. Eric [Cal15, Men15, Det15, Duc14, Hen24]. Erich [Ser21, Ste20]. Erickson [Mir17, Rei14a]. Erik [Lan16]. Erkenntniskritik [Cas10]. Erklären [Tur11, Fee10]. Erklärungen [Sch10, Hen13]. Ernan [Ano11a]. Ernst [Bia16, Kjæ11, dC21, Bre21a, Buz18, Gor21, Hei14, Ric08]. Eruditum [Duc17]. espace [Car18]. essai [Bre11]. Essay [Dom13, Dun24, Fal21, Jes14, Pal15, Pet18, Sch11a, Tub18, Bre11, RF10, Pra13, Gav24]. Essays [Rot17, GBT17, Hen24, Ken22, Kra22b, San22]. Essence [Har22]. Establish [Rei20]. Ethics [Stu11a]. études [Ler11, Ter10]. Euclid [Mil20]. Euclidean [IW18]. EUR [Ara19, Dob20a]. Europe [BF17, Gui18]. European [Coz24, Pal15]. Everett [Bac13, BB12]. everything [Bal13, Fut14]. Evidence [Ach13, Dom13, Har11, Dun14, SR18]. Evil [CJ21]. Evolution [Ade15, Dil13, FA11, Gon14, Har10b, Hun14, Leg24, Sna11, Har10a, Wil25]. Evolutionary [Bor10, ES17, Gri13, Kin18a, Kjæ11, Mar15, Pea18, Pod14, LNN12, Ric08, Sep12, Ear13]. Evolving [Sim11, BV07]. Exact [Ueb18]. Exceptional [Rus21]. excerpts [GD90]. Existence [BZ14, Cro16]. Experience [DeW17a, Kin18b, HMJ15]. Experiment [Cob12, Har22, Hat20a]. Experimental [BKV^{+}22, Fri15, Sch18]. Experimentalists [PS23]. Experimentation [PS23]. Experimenters [Geo19]. Experiments [BS13, Buz18, Buz21, Feh21, FMB12, Her19, Iva13, Iva20, Pat11, Row15, ST16, SF21, Stu21, IR11]. Explanation [Del17, DP19, Fal16]. Explanations [Sch24, Sch10]. Explicating [Mil15b]. Exploring [ER17, Gim11, Cow12, Rus19]. Extension [Bei22]. Externalism [De 24].

F [Gar15b, GP14, Pen18]. face [Bel13, Bor11]. Facts [Fab24]. Faculties [Dun24]. Failures [Llo13]. Faith [Lag14]. Fall [PS23]. Fancy [Gol11, Sar10]. Fantoli [Ela13]. Fashioned [Duc14, JM12]. Fate [KV22]. Fathers [Mag12, Nic11]. Feels [Pea18]. Feest [Tur11]. Feingold [Gui18, Sme14]. Felicitology [Stu11a]. Feminism [Koe13]. Feyerabend [Ken22, Bsc15, Koe13, Kub18, Sha20, Stu21]. Ficino [Rut21]. Fictions [ATE11, DeS12]. Fifty [Dam17, RD16]. Figure [DeS14, Pal15, Hol13, Lüt12a]. Final [Sch10]. Finalursachen [Hen13, Sch10]. First [Bei22, Ben13, Boy13, Hop24, HA12, How22, Loo11, Ver20a, Bor11, GLB12, SRK09]. Fisch [Bsc19]. fondateurs [Bre15, dP16]. Footed [Wil20]. Force [Duf11, Gaf10, Koc13]. Forces [Zam18a]. Forefront [Rot24]. Forensic [Lon23]. Forging [Tub21]. Form [Dow17]. Formal [Dam14]. Formalism [Ole16]. Formalist [Ueb11]. Formation [Gar24, Toa13]. Forms [Neu20, Van11, Hat09, Hat12b]. Fossil [Gri13, Sep12]. Foucault [Ray14]. Foundation [Omo21]. Foundational [Kin18b]. Foundations [Dun16, Dun17, IW18, Ueb18, Ver20b]. Four [Plu11, Wil20]. Four-Footed [Wil20]. Fourier [Esa19]. Fourth [DeW17a]. Fracastoro [Omo21]. Fraenkel [Gre15]. Framework [Agi24]. Frameworks [CW19]. francaise [dP16]. franccaise [Bre15]. France [Dro11, Chi08, Lon23]. Francesca [Kle12]. Francis [Ans15, Çim19, Sch17b]. Francks [Don24]. Francois [Hop24, Rhe24]. Frank [Mor17, Ueb11]. Frankfurt [Vra20]. Franklin [ST16]. Frappier [Row15]. Frédéric [Bre14]. Frederik [WG12, Koo10]. Frederique [Lav20, DeS12]. Frege [San16]. French [And15, Mai14, Sch21, Spi18, Bor11, Bou13, Bre11, Car18, CBCD12, Ler10a, Ler10b, Sch16, Ter10, Vae21]. French-Speaking [Sch21]. Freud [Par11, Wei08]. Friedel [Par11]. Friedman [Bia16, Doy24, Hat12a, Vat18]. Friedrich [Haa13, Ueb16, dC21]. Frisch [Her19]. Fritz [Mir11, Ryc13, Sie10, Sie10]. Front [Ano20a, Ano20b, Ano21a, Ano21b, Ano22b, Ano22c, Ano23, Ano24a, Ano24b, Ano24c]. Frontiers [Seg14, Kri11]. frühen [Hen13, Sch10]. Functional [De 22, Stu11b]. Functionalism [Hat15a]. Funktionsbegriff [Cas10, Hei14]. Future [Dor20].

G [Fer24, Loo11, Mar24, Mor24, Nas16]. G. [Tes16]. Gaffney [Duf11]. Gal [Lüt14]. Galen [Web24]. Galilei [She11, Sga17]. Galileo [Büt20, Ela13, Mun12, Fan12, Gal57, GD90, Mar10, She11]. Gambarotto [Zam18a]. Game [Eri15, Mir17]. Garber [Bre24, Jes14, Smi11a]. Garey [McD14]. Gary [Lav20, Cre20]. Gassendi [Bre24]. Gasser [Con24a]. Gasser-Wingate [Con24a]. Gaston [Cau24, Ler11, Ter10]. Gaukroger [Dun24, Jes11, Mel17]. Gee [Lat14]. Geisteswissenschaften} [Dew17b]. Gemma [Gie24]. Gender [Ueb21a]. Gene [Cam20]. Genealogy [Gol16]. General [Wil20]. Generality [CCR16, Eck18]. Generalization [CP21]. Generation [Jha12, Nye11, Str18, De 19]. generatione [Pic16, SBT10, Zup12]. Genetics [Cam20, Hop24]. Geoffrey [Mar17]. Geography [Kle12, Bor08]. Geometrical [Bia14, Rom24a]. Geometry [Car16b, Dun16, Dun17, McC22, Vin15]. Georg [Ser21, dC21]. George [Tub20]. Gerd [Vat18]. German [FS12, Mor24, Ste19a, Böh06, Gro15, NL12, Sch24, Sch10, Sie10, Zam18b]. Germany [Del19, JM17, Zam18a]. Gerson [Rem13]. Gestation [Ste19a, Zam18b]. Gibson [Lov16]. Gideon [Dow17, RES^{+}24]. Gijsbert [Gre20]. Gilles [Cau24]. Gilles-Gaston [Cau24]. Gimbel [Cow12, Ros13]. Giovanni [Rut21]. Girolamo [Reg21]. Given [Ueb21b, Sin22]. Glasner [Ant11]. Global [HES^{+}24]. globalization [Nic11, Mag12]. Goblot [Buz21]. God [Böh06, Mil20]. Goddu [Mos14]. Godefroid [Ant13]. Gods [Mun16, LM15]. Goldstein [Chi20]. Gordin [Rei14a]. Gordon [Rag19]. Gorham [Mar17]. Gorlaeus [DeS14, Lüt12a]. Gott [Böh06, Kat12]. Gottfried [Bee24]. Gotthelf [Boy13]. Goy [Men15]. Graham [Mil14a]. Granada [Rot17]. Grand [GD90]. Granger [Cau24]. Gravesande [Dob20a]. Gravity [Dom13, Har11, BKV^{+}22]. Greek [Mag12, Nic11]. Gregor [Nor18, Stu11c]. Gregory [Tab11]. Grete [Fal21]. Grinding [Bog13, Bal12]. Groningen [Kis24]. Groot [DeW17a]. Gross [Nas16, Ara19]. Group [Ear13, Ric15, Bor10, MP13]. Growth [Gri13, Sep12]. Grundfragen [Cas10]. Gruyter [Dob20a]. Guicciardini [Dun11]. Guide [Rey15, Pat14, PK08]. Günther [Mor24].

H [Duc12, Gig16, Gre15, Hic12, Lat11, Seg14, Ser21, Ste19a, Zup12]. H. [Fer24]. Haeckel [Kjæ11, Ric08]. Half [Sch21, Bor11]. Halflants [Ant13]. Han [Bee24]. Hand [Rap20]. Handbook [CCR16, Eck18, Lut24]. Hans [Pen13, Cam20, Hoq12, Pad11, Pra13, Sar17b]. Hans-Joachim [Sar17b]. Hans-Jörg [Pen13, Cam20, Hoq12, Pra13]. Hanson [Nic12, Lun10]. Harmon [Nas16]. Harmonices [Rom24a]. Harmony [Gub19]. Harper [Dom13]. Harrison [Hic12]. Harry [Fal17]. Hart [Sna11]. Hatfield [Kle15a]. Hattab [Van11]. Haytham [IW18]. Health [Web24]. Healthy [De 24]. Heat [Esa19]. Heaven [Rot17, GBT17]. Heavenly [Omo21]. Heavens [Fer22, Lat11, Rut21, ABS10]. Hegel [Fol22, Lin18, Red24, Wes15]. Heidelberger [Nor18, Stu11c]. Hein [Gam15]. Heisenberg [AP22]. Helen [Van11]. Hélène [ASF^{+}21, CJ21]. heliocentrism [God10, Mos14]. Helmholtz [Nor18, McD14, Meu10, Neu18]. Helmholtzian [Gio17]. Helmont [Dem21]. Helmut [Nor18]. Henri [Sin22, Sch17a]. Henrik [Roq19]. Henry [Pur15]. Her [CJ21]. Herbert [Kas24, Mar24]. Herder [Fol22]. heredity [MWR12, Pen13]. Heritage [Rad13]. Herman [Edd15, Pow12]. Hermann [Ryc13, Sie10, Bia14, Edg20b, Edg20a, Fal21, Nor18, Sie10]. Herring [Dor20]. Herschel [Cob12, Hon18, Pen18]. Hesse [Gor21]. hidden [Böh06, Böh06]. Hierarchy [Dun16, Fol19]. Higher [Hat20a]. Hill [Mar17, Roq19]. Himself [Hen23]. Hirai [Ekh13]. Hiro [Ekh13]. histoire [Ler10a, Ler10b]. Historian [ASF^{+}21, CJ21]. Historians [CW19]. Historical [ADJ^{+}15, BS13, Dam14, Det15, Fee10, Gre21, Ken18, Kin18a, Mar15, Mil15b, Mir11, Mur24, Sch18, SR18, IR11, MAKA09, Wat13, Det16, Tur11]. Historicism [DFK20]. Historicizing [Pra13, RF10]. Historico [Ari17]. Historico-Philosophical [Ari17]. Histories [DL11, Gro20, Hoq12, Lær18, Lev15, Rhe10, Sch21, Sal12]. Historiographer [ASF^{+}21]. Historiography [Mel19, Sch21]. History [And13, Bsc19, Buz21, Cap20, Chi08, D’A24, De 24, DeS14, Del19, Dor20, Dro11, Ear13, Fer22, Gaw11, Geo19, Hat15b, Hop24, JM17, Kis24, Lan19, Lau24, MV18, MRR17, Nem16, Par11, Pen13, Plu11, Pre17, Rem13, Rey15, Ric12, Rot17, Ryn14, Sch14a, Sch14b, Seg14, Ste19b, SF21, Tim19, Ueb17, Ueb21a, Vio13, Wra19, Bar13, Bor10, DeW04, Fis17, Ger10, GBT17, Kri11, KM11, Ler10a, Ler10b, Lüt12a, Mar14, MWR12, Pat14, Tos09, Wei08, Lav20, GBV^{+}22, Kor18, Lol15]. Hobbes [Con24b, PG19]. Hodge [Tab11]. Hoffmann [dC21]. Holenstein [Pal15]. Holism [Slo11b]. Holland [Are12]. Holly [Lan19]. Homocentrism [Omo21]. honor [GP14]. Hookway [Kas24]. Hopkins [Büt20]. Horky [Mil14b]. Horrocks [Ben13, HA12, Ben13]. Hot [Reg21]. Hubert [Pal15]. Hugo [Rot24]. Hull [Hon18]. humaines [Bre11, CL12]. Human [DP19, Det15, Har14, Nem16, Bre11, Isa12, Wat13]. Humanism [Bre24, Ekh13, Hir11]. Humanities [MRR17, Ste19b]. Hume [Sch13, DP19, Rus07, Vat18]. Husserl [Cau20, Toa13]. Huygens [Mar21]. Hylozoism [Pec21]. Hypothesis [Fol19]. Hypothetical [HS09, Her19, Sch24, Stu11c].

Ibn [IW18]. Icy [Tub21]. Idea [D’A24, Lin18]. Ideas [Mét24, MRR17, Ste19b]. Ideology [Hof13, AN10]. Ierodiakonou [BS13]. II [Ler10b]. Illustrate [Nas16]. Illustrated [Dam14]. Ilse [Mee24]. im [Böh06, Fee18, Kat12, Ryc13, Sie10]. Image [Gro15, Roc10, BV11]. Imagination [Bre21a, Gre21, Her19, Roc10, BV11]. Immanence [Bog13, Bal12]. Impact [BM16]. Impera} [SB20]. Import [Bsc15]. Ina [Men15]. including [GD90, Nie14]. Inconsistent [Fra12]. Incorporation [Dew17b]. Individuality [Ken18]. Induction [Det18, Ver24]. Inductive [ER17, Rus19]. Inductivism [Cob12]. Inevitability [Kin16, STP15]. Inference [Vio24]. Infinite [TF24]. Infinitesimal [Edg20b, McC22]. Infinitesimals [MK13]. Infinities [Wag14]. Infinity [Gan21]. Influence [Wra16]. Informationalization [Ant22]. Inner [Kin18b]. Innovation [McC24, FPS11, Gre15]. Inquirers [Vio24]. Inquiry [Len11, Cre20]. Inquisition [Reg21]. Insight [Nas16]. Insights [Buz21, Wes15]. Instances [Sch17b]. Instrumentalism [Hen23]. Instrumentalist [Hen23]. Instrumentality [Irv12]. Integrated [Dor20]. Integrating [Ken18]. Integron [Rhe24]. Intelligent [New14, Wax13]. Intention [Rad13]. Interactions [Mun16, LM15]. Interested [Fut14, Bal13]. Internal [Lin18]. International [Tub18]. Interpretation [Bac13, BB12]. interpretatione [Cro16, BZ14]. Interpretations [Dun16, Iva17, Stu15a]. Interpreting [Ken22, Kra22b, San22]. Interpretive [Kor18, Pre17]. Intersection [Ros13, Gim12]. Intersubjectivity [Gub19]. Intolerance [Rot24]. introduced [Dun24]. Introduction [BV22, DRW21, Gaw11, MV24, MP18, Pop14a, Sar13, Sar16, SR21, Stu15b, DeW04, SBT10, Ari17, Zup12]. Intuition [Dun17]. Inutility [Gol16]. Inventing [Pow12, Edd15]. Inventor [Hen23]. irreligion [Rus07, Sch13]. Isaac [Dom13, Duc14, Har14, Ach13, Art14, Dom13, Duc11, Dun11, Gui09, Har11, JM12, vB13]. Isabelle [Fra13]. ISBN [Bal19, BKV^{+}22, Büt20, Cam20, Cap20, Cre20, DP19, Dob20a, Dor20, Dun24, Geo19, Gre20, Gro20, Hat20b, Lav20, Mil20, Pea19a, Rag19, Rap20, Tub20, Ver20b]. Ishiguro [BBK^{+}16]. Isles [McC15, Sga12]. Issue [DRW21, MK13, SR21, dC21]. Italian [Tos09]. Italy [Mar10, Mun12]. Itself [Rei20]. Itzigsohn [HES^{+}24]. IV [Len14, Pop14b]. IV.12 [Gil14]. ix [Dob20a].

J [Atk24, Dam17, Dom11, Ito14, Kjæ11, Loo11, Mil20, Pen17, Rot17, Zup12]. J. [Bia16, Esa19]. Jacob [Agi24, Hop24, Mét24, Rhe24]. Jacques [Dob20b]. James [Kle12, Lan16, Row15, Tub20, Ach13, Ban14, Bor08, Hat15a, Kin18a, Mis15, Pea19b, Wra16]. Jamie [Ken22]. Janiak [GBV^{+}22]. Janssen [Lav20]. Janssen-Lauret [Lav20]. Jean [Nem12, Sin22, Buz21, Cau20, Cau24, Cok20, DeW17a, Iva20, Nem12]. Jean-Philippe [Sin22]. Jeanne [Kis24]. Jed [Sme14]. Jeffrey [Don24, Bac13, Ham24]. Jennifer [Teu15]. Jeremiah [Ben13, Ben13, HA12]. Jeroen [Fog12]. Jesseph [Det16]. Jevons [San16]. Jewett [Rei14b]. Jewish [Ros13, Gim12]. Jill [Loo11]. Jimena [Sau17]. Jo [Jha12]. Joachim [Sar17b]. Jobin [Rag19]. Jocelyn [Are12]. Joel [Har14]. Johann [Are12, D’A24, Hol10, RH10]. Johannes [Zup12]. John [Bal19, Gie24, Pen18, Ada11, Ans11, Blo12, Bro12, Cob12, Cro16, Edd15, Fal16, Hen23, Hil12, Kra22b, Mag15, Mon15, Poo10, Ste19a, SBT10, Zup12]. Johns [Büt20]. Johnson [Gie12]. Jonathan [Tab11]. Jones [Dor20, Ste13]. Jonkers [Gre20]. Jordi [Cap20]. Jörg [Pen13, Cam20, Hoq12, Pra13].

Jorink [Duc14]. Jörn [SR18]. José [And15, HES^{+}24]. Joseph [Nas16, Esa19]. Journal [Don24]. Judgment} [De 22]. Judy [Rei14a]. Jugendbewegung [Mor24]. Juha [Haa13]. Jungnickel [Del19]. Just [Kri12]. Justification [Mel19]. Justin [Gig16, Gre15, Duc12]. Jutta [Geo19].

K. [San22]. Kaila [Nem16, Neu18]. Kammer [Sin22]. Kanjirakkat [Rag19]. Kant [Car16b, Bei22, De 22, Fer22, Fra12, Gam15, Gav24, GW14, How22, Mas08, McN16, McN22, Mee24, Men13, Men15, New14, Pec21, Sou13, Sta11, Teu15, Van13, Vat18, Vin15, Wax13, Wil20, Zam24b]. Kantian [Hat12a, FN06, MK13]. Kantianism [Mee24, Ray14]. Karida [HES^{+}24]. Karim [Ken22]. Karine [Eck18]. Karl [Nie14, Sar17b, Ueb18]. Katerina [BS13]. Kaye [Loo11]. Kekulé [BV11, Roc10]. Kelly [Mir11]. Kemp [Lav20]. Kendall [Lag14]. Kendig [DL18]. Kenneth [Mar17, Vio24]. Kepler [Bon13, Ito14, McM11, Rom24a]. Kevin [Dor20, MP18, Rus19]. Key [RH10, Hol10, Are12]. Kienzler [Doy24]. Kim [Tur24]. Kinêsis} [De 19]. Kindi [Obe14]. Kinds [DL18, Ken16, Mag15, War24]. Kiprijanov [Dor20]. Klaas [Art14]. klassischen [FES^{+}23]. Klein [Rei14a, Mis15]. Know [And13]. Knowing [Pur15, Sch14a, KM11]. Knowledge [Gol17, Gro15, Har14, Kus19, Loo11, Nem16, Pal15, Tes16, Tim13, Bor08, Bro16, Hol13, Isa12, SRK09, Kle12]. Known [Rom24a]. Konstantin [Dor20]. Konstruktivismus [Ryc13, Sie10]. Kooijmans [WG12]. Kopp [BV11, Roc10]. Kreis [Ueb16]. Krieger [Seg14]. Kuhn [Kin18a, Ric13, San22, Tub20, Dam14, Dam17, Isa12, KA12, Mar15, Mel19, Obe14, PK08, RD16, Tim13, Tim19, Wra16, Wra19, Wra21, Har14]. Kwa [Sch14a]. Kyung [Tur24]. Kyung-Man [Tur24].

L [Dom13, Hic12, Hof13, HES^{+}24, Rei14a]. Laboratory [Coz24]. Lagerlund [Roq19]. Lakatos [Gon14]. Laks [Fal18]. Landy [DP19]. Language [Fra18, Koc13, Mar17, Mat24, McE18, Pea19b]. Larry [Sch18]. Last [Mar24, Ari99]. Late [Fal21, Mar21, Mun12, Rem13, Sch24, Ger10, Mar10, Tos09]. Later [Gre13, NZ10, Ueb11]. Latour [Fab24]. Laudan [Sch18]. Laura [Dor20]. Laurence [McD14]. Laurent [Bre14]. Lauret [Lav20]. LaVine [Ueb21a]. Law [Del12, Dik22]. Lawrence [Mun16]. Laws [Gor11, PG19, Ott09, Ans15]. Lazarus [Rei20]. Lea [Don24]. Learn [De 24]. Learning [Gol17, Bro16, Poo10, Ada11]. Lebensreform [Mor24]. Lecons [CBCD12, Bre14]. lecture [Nie14]. Lectures [Mel19, CBCD12]. Legacy [And15, Bro24, Hat12a, FN06, Mai14]. Lehoux [Str18]. Leibniz [Bee24, Duc12, Ham24, Ado19, BBK^{+}16, Don24, Gar09, Nob19, Slo24, Smi11a, Smi11b, Sou13]. Leiden [Gre20, Pal15]. Lemaître [BS22]. Lemov [Rei14a]. L’empirisme [Nem12]. Léna [Kin16]. Lens [Dik22]. Lenses [Bog13, Bal12]. Lenz [Car16a]. Leroux [Nem12]. L’espace [Ren18]. Lessons [WG12, Koo10]. Letitia [Row15]. Letter [GD90]. Letters [GD90]. Levey [Don24]. Levine [Pod14]. Levitin [Lær18]. Lewis [Pic16, Sin22, OS17, Sin19]. Libri [Rom24a]. libros [SBT10, Zup12]. Lichtenberg [Tes16]. Lidgard [Ken18]. Liebscher [FS12]. Life [Bro24, Ekh13, Gar15b, Gie24, Gig16, Kjæ11, Lin18, Mih24, Mor24, Rhe10, Sha21, Stu16, Bru05, CD13, Hir11, NS14, Nie14, Ric08, Smi11b, Bal19, Duc12, Hoq12, Sar17b, Coz24, Hop24]. Light [Gor21]. Like [Sch17a, Sga17]. Lilienau [Lut24, Ueb16]. Limbeck [Lut24, Ueb16]. Limbeck-Lilienau [Lut24, Ueb16]. limitations [Bru05]. Limits [Gil14, Wes15]. Lindley [Gar15b]. Line [HES^{+}24]. Linguistic [D’A24]. Linking [Fer24]. Lisa [Gol11]. Literature [ATE11, DeS12]. Lloyd [Pur15, Rem13]. Locality [Kri12]. Locke [Hil12, Jon13, Ans11, Cor11, Con24b]. Logic [Bro12, Coz24, CP21, Har10b, Hop24, McC15, McC24, Vae21, Zam24a, Bal12, Har10a, Mad14, Sga12, Wil25, Bog13, Rog15, Sna11]. Logical [Fra18, Gio17, Haa13, KV22, Lau24, Lut24, MP13, Mor17, Mor24, Nem16, Ric15, Rog15, Tub18, Tub21, Ver20a, Vio13, Vra20, Ler10b, Mad14, Man09, AES^{+}24]. logique [Ler10b, Nem12]. Logischen [Ueb16]. Logischer [Mor24]. London [DP19, Dor20, Lan19, Rag19]. Look [Sga17]. Lorraine [Dam17, Rei14a, Sal12]. Losee [Fal16]. Lost [Rei14a, EKD^{+}13]. Louis [DS21]. Love [Pea19a]. Lucas [Bal19]. Luft [Bia16]. Lukas [McC24]. Lunbeck [Sal12]. Lund [Nic12]. Lüthy [DeS14]. Lutz [AES^{+}24]. Luuc [WG12]. Lydia [Rey15]. Lynn [Ken18].

M [Det16, Dor20, Hat20b, Lol15, McC24, Rag19, Slo11a, Spi18, Zup12, dP24]. Maas [Duc14]. macchie [Rig66]. Macdonald [Wes24]. Mach [Kra22b, Lan16, Ban14, Bre21a, Buz18, Gor21, Tex21]. Machamer [Dom11]. Machines [Duc12, Smi11b]. Macmillan [Bal19, Lav20, Sin19, Ste19b]. Maddy [Rog15]. Made [Eri15, Toa21, Mir17]. Magic [Ant13, dCH11]. Mahut [Mel17]. Maia [And15]. main [Dom13, Duc11]. Making [Har14, Isa12, vB13, Gro20]. Man [Lut12b, Mel17, Mil20, Tur24]. Manninen [Haa13]. Manning [Dow17, RES^{+}24]. Map [Kle18]. Marc [Mir11, Con24a]. Marco [Gol16, Kat12, McC15]. Marcum [Kin18a]. Margaret [Gol11, Sar10, Wes24]. Marguerite [Gab24]. Maria [Feo11, Fri15, vS15]. Marina [Eat21]. Mark [Mun16, Ear13]. Marr [Mun12]. Marshall [Bon16]. Marsilio [Rut21]. Martin [Bee24, BM12, FS12, Hat15b, Pal15, Car16a]. Martyn [Pur15]. Mary [Pea19a, BKV^{+}22, Gor21, Jha12]. Massimi [Sta11]. Massimo [Bee24]. Masthead [Ano16a, Ano16b, Ano17a, Ano17b, Ano18b, Ano18a, Ano19b, Ano19a]. Materia [MM23]. Material [Len14, Rus21, dP24]. Materialism [Gim21, Ari17, Con24b]. Mathematical [Buz21, Cau24, Çim19, Dob20b, Dun11, Gro20, McN22, Mun12, Pad11, Ser21, Gui09, Mar10, Rei06, REG08]. Mathematicians [LM15, Mun16]. Mathematics [Böh06, CCR16, Dun17, Eck18, IW18, Mil16, Mun16, LM15]. Mathematik [Böh06, Kat12]. Mathematization [Mar17]. Matt [Ueb21a]. Matteo [McC24]. Matter [Ano20a, Ano20b, Ano21a, Ano21b, Ano22b, Ano22c, Ano23, Ano24a, Ano24b, Ano24c, Art14, Ekh13, Rus21, Sha21, dC21, Hir11, vB13, Dow17]. Matthew [Dor20, Nic12]. Mattia [Rod24]. Maxwell [Ach13]. McCartney [Mun16]. McCormmach [Del19]. McDonough [Don24, Ham24]. McElvenny [Pea19b]. McGrew [Mir11]. McGuire [Dom11]. McKaughan [Lan19]. McMullin [Ano11a]. McOuat [Rag19]. Meaning [Pea19b, SR18, Mac82, McE18, Tau09, Kuu13, Nas16]. Measure [Rom24a]. Measurement [De 22, Neu18, San16]. Mechanical [Jes14, vB13]. Mechanics [Bac13, BB12, DeW17a, Fal21, NP22]. Mechanik [FES^{+}23]. Mechanism [Eat21, Ito14, Jes11, Bon13, Gau10]. Mechanisms [Gar15b, CD13, Hat09, Hat12b, Van11]. Mechanization [GR13, Jes14]. Medawar [Nie14]. Medica [MM23]. Medical [Dem21, Hir11, Ekh13]. Medicine [Lon23]. Medicus [Sie10, Ryc13]. Medieval [Ant11, Gla09, Mil15a, GP14]. Meeting [Rig66]. Meike [Mor24]. Mélanie [Row15]. Melting [Tub21]. Memoriam [Pat18]. Men [Tub20]. Menachem [Bsc19]. Mensch [Teu15]. Meritocracy [Llo13]. Mess [Reg21]. Messenger [GD90]. Metaphilosophical [Ole16]. Metaphysical [Man15]. Metaphysics [Gab24, Hen24, Jal24, Mur24, Pea18, Vra20, Zam24a, dP24]. Metasemantic [De 24]. Meteorology [BM12, Gil14, Len14, Pop14b, Mar11]. Method [Boy13, Cow12, Dom13, Fra12, Gar24, Gav24, Geo19, MM23, Pop14b, Ver24, Ach13, Gim11, GLB12, Gui09, Har11, Dun11, Dun14, Edg20b]. Methodical [Gro15]. Methodische [Ara19, Gro15]. Methodological [BS13, Dil13, Ueb21b, IR11]. Methodology [Fra12, Gon14, Kri12, Pen18, Sch18, Duc11, Dom13]. Methods [Cre20]. Metzger [ASF^{+}21, CJ21]. Meulders [McD14]. Meyerson [Mil15b]. Meynell [Row15]. Michael [Doy24, Hic12, Pen17, Rei14a, Hat12a, Jha12, Mil20, Nor18, Nye11, Stu11c, Vat18]. Michel [Bre14, Gui14, McD14, Pea19a]. Michela [Sta11]. Michiel [Zup12]. Microbial [Hop24]. Mid [Pal15]. Mid-Eighteenth [Pal15]. Middle [Dam14]. Miguel [Rot17]. Mihnea [Gar15a]. Milkov [Ric15]. Mill [Esa19, Mag15]. Miller [Bon16]. Mind [Chi08, Dro11, Jon13, Rei14a, RES^{+}24, Cor11, EKD^{+}13, MM09, Wax13, Dom11, New14]. Minds [Tub20]. Mineral [Lov16, Gib15]. Mirabilis} [Feh21]. Miracle [Ham24]. Mirandola [Rut21]. Misak [Pea14]. Mix [Bal19]. Modeling [MCP19, SB20]. Models [Ste13, BJ09]. Modern [BF17, Bre24, Car16a, Dow17, Fra22, Gig16, Gor11, Gui18, Hat15b, Hat20b, Irv12, Jon13, Lon23, Loo11, McC24, Rot17, RES^{+}24, SB20, Sar17a, Sch16, Spi18, Cor11, GBT17, Mar14, NS14, Ott09, Sch10, SRK09, Yeo14]. Modernism [Pea19b, McE18]. Modernist [Fer24]. Modernity [HES^{+}24, Jes11, Wil08, GG08, Gau10, Bla12]. Modesty [Duc17]. moitié [Bor11]. Molecular [Coz24]. Molecules [Gie24]. monad [Gar09, Smi11a]. Monadology [Pec21]. mondo [Feo11, Tos09]. Monism [Tex21, Ban14]. Monod [Mét24]. Montessori [Fri15]. Mordechai [Gui18, Sme14]. Morgan [Ver24]. Morris [Lüt14, Smi24, Ste19b, Ver20b]. Morris-Reich [Ste19b]. Motion [PG19, vB13, Art14]. Motivating [SF21]. Movement [De 19, Rot24, SDR^{+}21, Mor24]. Moving [McM11]. Mud [Str18]. Mugnai [Bee24]. Müller [Cam20, Pen13, Ede24]. Müller-Wille [Pen13]. Multiple [Cok20]. Mundi [Rom24a]. Must [Rei20, Rog15, Mad14]. Mutio [Mun12, Mar10].

N. [Lun10, Nic12]. Naaman [Gre13]. Naaman-Zauderer [Gre13]. Nachtomy [Gig16]. Naess [Rad13]. Narboux [Sin22]. Narratives [Rag19]. Nasim [Rap20]. Natural [Ant11, Böh06, Don24, Ekh13, Fer22, GR13, Gla09, Gol11, HS09, Hil12, Irv12, Jes14, Ken16, Lov16, Mag15, Mar21, Mar17, Mét24, Pop16, PG19, SB20, Sim11, Stu11c, Ans11, BV07, Dom13, Duc11, Ebr15, Gib15, Hir11, Sar10, Tos09, DL18]. Natural-Philosophical [Dom13, Duc11]. naturale [Feo11, Tos09]. Naturalism [DFK20, Dem21, Dil13, Sch13, Rus07]. Naturalistic [Zam24b]. Naturalization [Stu11a]. Nature [AES^{+}24, Cap20, Car16a, Dem21, DP19, Det15, Fal17, GS17, Hic12, Lag14, Mar17, PG19, Rag19, Tes16, Wes24, Wil20, Zam18a, HNS11, Hol10, KW10, Ott09, RH10, Wat13, Are12]. Naturwissen [Böh06, Kat12]. Nebulae [Rap20]. Nebular [Fra22]. Neglected [Buz21]. Neil [Pic16]. nel [Feo11, Tos09]. Neo [Mee24, MK13, Ray14]. Neo-Kantian [MK13]. Neo-Kantianism [Mee24, Ray14]. Neopositivism [Vae21]. Nescire} [Duc17]. Netherlands [Duc14, JM12]. Neto [And15]. Network [SB20]. Networked [Ros22]. Networks [Haa13, Man09]. Neurath [Cap20, Fer24, Gro15, Rad13, Stu11a, Ueb21b]. Neuraths [Ara19, Gro15]. neuroscience [Meu10, McD14]. Neutral [Tex21, Ban14]. Neuzeit [Hen13, Sch10]. Never [PS23]. Newman [Sim11]. Newton [Dom13, Duc14, Dun11, Gui18, Ach13, BKV^{+}22, BS14, BF17, Bor11, BF12, Cal15, Dom13, Duc11, Duc14, Gui09, Har11, Hen24, JM12, Koc13, Ngu23, Sch11a, Slo11b, Sme14, Bel13]. Niccolò [Dun11]. Nicholas [Böh06]. Nicholls [FS12]. Nicolaidis [Mag12]. Nicolaus [Böh06, Kat12]. Niemann [Sar17b]. Nietzsche [Gie12, Joh10, Rod24]. Nieuwentijt [Duc17]. Nikolaus [D’A24]. Nikolay [Ric15]. Nineteenth [CP21, FS12, Hat12a, Lat11, Pen18, Pod14, Rap20, Sch24, Sch21, LNN12, ABS10, FN06, NL12]. Nineteenth-Century [CP21, FS12, Hat12a, Lat11, Pen18, Sch24, ABS10, FN06, NL12]. NJ [Gro20]. Noa [Gre13]. Noble [GP14]. Noether [Toa21]. Noncanonical [SB20]. Nordic [Haa13, Man09]. Nordmann [Hat12a]. Norman [Ryc13]. Norms [Car16a, Len11]. Note [Ano22a, Pat17, Pat20, Pat21a, Pat21b, Pat21c, Pat24]. Notebooks [Yeo14, Sar17a]. Nothing [GS17]. Nouveau [dP21]. Novel [Bia16]. Novoa [Pod14]. Number [Cok20, Böh06]. Numbers [Hic12, Hof13]. NY [Tub20]. Nyden [Gar15a]. Nye [Jha12]. Nyhart [Ken18].

objectif [Bre11]. Objective [Gub19, Bre11]. Objectivity [Cau24, Toa13]. Observation [Ben13, Her19, HA12, Kub18, Nic12, DL11, Lun10, Sal12]. Observatories [Lat11, ABS10]. Observing [Rap20]. Occasions [Ric12]. Oddi [Mun12, Mar10]. Oeuvres [CBCD12, Bre14]. Ofer [Lüt14]. Often [Pec21]. Ogden [Pea19b, McE18]. Ohad [Gig16]. Olen [Sin19]. Omar [Rap20]. Omens [Hic12, HNS11]. Ontology [Hol23, Slo11b]. Open [Atk24]. opinions [Gal57, GD90]. Opposition [Sou13]. Optimality [Ham24]. Opus [Gam15, Van13, McN16]. Order [Det15, Pea19a, Gib15, Wat13, Lov16]. Ordinary [Fra18]. Organicism [Teu15, Men13]. Organism [Leg24]. Organismic [Agi24]. Organisms [Gie24]. Organization [Zam18a]. Origin [Dil13, Fal18, Sme14, BF12]. Origins [Bla12, Edg20a, Jha12, Mil15a, Ros22, Sta16, Wil08, Nye11]. Orthodoxy [Mag12, Nic11]. Ostwald [Del12]. Other [Mar21]. Ott [Gor11]. Otto [Ara19, Gro15, Lat11, Fer24, Gro15]. Our [Can16, Sau17, Rei20]. Overcoming [Vra20]. Oxford [Dun24, Eck18, Hat20b, MP18, Rus19, CCR16].

P [Rem13]. Padua [Omo21]. Paleobiology [Gri13, Sep12]. Palgrave [Bal19, Lav20, Sin19, Ste19b]. Palmerino [Bre24]. Pancritical [Sep24]. Paniel [Kas24]. Pansophia} [Duc17]. Paola [Eat21]. Pap [Stu11b]. paper [Bsc19, Cam20, Lan19, MP18, Rag19, Rus19]. Papers [Kis24]. Paradigms [Tub20]. Paradoxicality [Ade15]. Parmenides [Mil14a, Gra13]. Parsons [Har14, Isa12]. Part [Dun16, Dun17, dCH11]. Particle [ST16]. Parts [Slo11b]. Past [Dor20]. Path [Ain12, Mos14, God10]. Patrick [Rot17, Ito14]. Patton [Rey15]. Paul [Zup12, Mir17, Rei14a, Sch13]. Paved [The20]. Paz [vS15]. Peano [TF24]. Pearce [Leg24]. Peckhaus [Ric15]. Pedagogy [Gro15]. Peeling [Bal12, Bog13]. Peijnenburg [Kis24]. Peirce [Atk24, Pea18, Red24, Vio24]. Penelope [Rog15]. Pennsylvania [Mil20]. Perception [Neu20]. Perinetti [Gre15]. period [Sch10, RES^{+}24]. Periods [Gre21]. Perrin [Cok20, Hud20, Iva13, Iva20]. Personal [Tim13]. Perspective [MCP19]. Perspectives [Car16a, Fab24, Mee24, OS17, Seg14, Sin19, Tur11, Fee10, Kri11, Wat13, dP14, Cap20, Det15, Ken18]. Peter [Bac13, Dom11, Duf11, Dun14, Hic12, Hil12, Koe24, Sin19]. Petit [Bre18]. Petrarch [Sga17]. Phenomena [Çim19]. Phenomenology [Cau20, NP22]. Philadelphia [Mil20]. Philip [Fut14]. Philipp [Mor17, Ueb11]. Philippe [Sin22]. Phillip [Mil14b]. Philosopher [Ant22, Can16, CJ21, Cok20, Sau17, vS15, dP14]. Philosophers [CW19, Hon18, Ric12]. philosophes [Bre15]. Philosophical [Ari17, Dom13, Kas24, Ken18, Kob14, Kra22a, MM23, Rod24, Smi24, Wes24, Duc11, Ric10]. Philosophie [Fee18, Bor11, Bre14, CBCD12, Ler10a, Ler10b, Bre18]. Philosophies [Gro20, Sch21]. Philosophische [Nor18]. Philosophy [And13, And15, AES^{+}24, BM16, Bia16, BC17, Bre21a, Bre21b, Bsc19, Buz21, Cap20, Car16a, Cau20, Chi08, Cob12, CBCD12, Dam14, De 24, DeS14, Del17, Dob20a, Dom13, Dor20, Dro11, Eat21, Ekh13, Fal18, Fal21, Fol22, FMB12, Fri15, Gaf10, Gam15, GR13, Gaw11, Gla09, Gol11, Gor21, Hat15b, How22, Irv12, Iva17, Jes14, KV22, Kin18a, Kis24, Kle12, Koe13, Kor18, Lær18, LH16, Lan19, Lau24, Lev15, Loo11, Lut12b, MCP19, Mar15, Mar17, MAKA09, MV18, MP13, Mil16, MK13, Mos14, Par11, Pat14, Pea19a, Pre17, PG19, Rat24, Rem13, Ric12, Ric15, Rom12, Rom24b, Ros22, Row15, Rus09, Ryn14, SB20, Sch18, Sch24, Sch11a, Seg14, Sid24, Sta11, Ste19a, Ste20, Ste13, SF21, Stu16, Tim19, Tur24, Ueb17, Ver20b]. Philosophy [Vio13, Wes24, Wra19, Wue13, Zam18a, Ans11, ADJ^{+}15, BJ09, Bar13, Bor08, Bor11, DeW04, Duc11, Fis17, Ger10, God10, Hir11, Kri11, Ler10a, Ler10b, Lüt12a, Mai14, Mar14, Mas08, Men13, NS14, Ott09, Sar10, SRK09, Stu15a, Van13, Wei08, Zam18b, vB13, Don24, Hat20b, Mir11, Rey15, Roq19, Ant11, Bre24, Bro24, Det16, Dow17, Duf11, Gig16, Gor11, Ham24, Hil12, Jes14, Leg24, Teu15, Ueb21a, Zam24b]. Phoneme [D’A24]. Physical [AES^{+}24, Çim19, Esa19]. Physicist [Can16, Del19, JM17, Sau17]. Physics [Ant22, Ant11, Coz24, Del19, Dob20b, Fal17, Gla09, Ham24, JM17, Mur24, Ros13, ST16, Gim12, AES^{+}24, Tes16]. Physik [Rot24]. Physiogony [Fer22]. Physiology [Pea18, Ste19a, Zam18b]. Piaget [Buz21]. Pickering [Kin16, Pea19a]. Pico [Rut21]. Picture [Mat24]. Pieces [Dom13]. Pierre [Bre24, Buz18]. Pittsburgh [Pea19a]. Place [Lav20]. Planck [Del12]. Plasticity [De 24]. Plato [Gre20, Hor13, Mil14b]. Platonic [Omo21]. Please [Lon13]. Pluralism [Bok08, Bsc15, Feh21]. Poincaré [BM16, BC17, Bre15, Dun16, Dun17, Hol23, Iva13, Sch17a, Sta16, Stu16, dP14, dP21, vS15]. Point [Ant11, Gla09]. Polanyi [Jha12, Nye11, Tim13]. polarity [BV07, Sim11]. Political [Rom12]. Politically [Rom24b]. Politics [Pea19a, Ros13, Sch17a, Tub20, Gim12, Gab24]. Pomponazzi [BM12, Mar11]. Poole [Ada11]. Popper [Sar17b, Bsc15, Due23, ES17, Nie14, Sep24, Sha20]. Popperian [Sha20]. Populärwissenschaftliche [Nor18]. Positive [Bre21b, Bre14, CBCD12]. Positivism [Bre21b, DS21, Fra18, Mee24, Mil15b, Ver20a]. positivisme} [dP21]. Postempiricism [Gor21]. Posteraro [Rat24]. Posterior [Gol17, Bro16]. Postgenomics [Cam20]. postumum [Van13, Gam15, McN16]. Pot [Tub21]. Potatoes [Bog13, Bal12]. Powers [Edd15, SDR^{+}21]. Pp [Ara19, Bal19, BKV^{+}22, Bsc19, Büt20, Cam20, Cap20, Cre20, Del19, DP19, Dob20a, Dor20, Dun24, Geo19, Gre20, Gro20, Hat20b, Lan19, Lav20, Mil20, MP18, Pal15, Pea19a, Pea19b, Rag19, Rap20, Roq19, Rus19, Sin19, Ste19a, Ste19b, Tub20, Ver20b]. Practice [Cob12, CP21, Ken16, Pal15, Pop16, Ebr15, Hol13, Gro20, DL18]. Pragmatic [Kub18, Kas24]. Pragmatism [OS17, Leg24, Sin19]. Pragmatisms [Tub21]. pragmatists [HMJ15, KK60, Mis13, Pea14]. Praxis [Vra20]. Precization [Rad13]. Precursors [CJ21]. Preestablished [Gub19]. Preformation [Nob19]. Pregnancy [Sid24]. Prehistory [Ser21]. première [Bor11]. Present [Bei22, Dor20, KM11, Sch14a]. Present-Day [Bei22]. Presentism [Vir22]. Presocratic [Fal18]. Press [Bsc19, Büt20, Cam20, Cre20, Dun24, Geo19, Gro20, Hat20b, Mil20, MP18, Pea19a, Pea19b, Rap20, Rus19, Ste19a, Tub20, Ver20b]. Preston [Kor18, Kra22b]. Princeton [Gro20]. Principia [Dom13]. Principle [Ham24, Hun14, Wel23, AP22, Edg20b]. Principles [Boy13, Gio24, Gro15, PG19, GLB12, SRK09, Loo11]. Prinzipien [Ara19, Gro15]. Priori [Sin22, Stu15a, Iva17, Stu11b]. Prize [Gav24]. Probability [Pad11, Ver24, Rei06, REG08]. Problem [Hat20b, Kob14, Mee24, Sha20, Sid24, Wue13, Ric10, STP15, Kin16]. Problems [Plu11, dP14]. Proceedings [Rig66]. Processes [Gie24]. Program [Dam14]. Programmatic [Sch11a]. Progress [Pea19a]. Project [Kin18b, Smi24]. Proliferation [Bsc15]. Proof [Hud20, Lon23]. Proper [Gam15, Van13]. Protocol [Cat12, Ueb07]. Protocol-Sentence [Cat12, Ueb07]. Protoplasm [Pea18]. Psychologism [Kin18b]. Psychology [DFK20, Dam20, Fré20, Fri15, McN22, Ole16, Vae21, Rod24]. Ptolemaic [Cau24]. Public [Stu16]. Publishing [Atk24]. Pulte [Nor18]. Pure [McC22]. Purity [Ant13]. Purposiveness [Lin18]. Putnam [Cre20]. Putting [Dom13]. Pythagoreanism [Hor13, Mil14b].

Quaestiones [SBT10, Zup12]. Qualities [San16]. Quantenphysik [FES^{+}23]. Quantities [San16]. Quantity [Ngu23]. Quantum [Bac13, BB12, Bok08, Bok15, Fal21, Hep11, NP22, Wüt14, vS22, Cao10, Sto13]. Quantum-Classical [Hep11]. Quarter [BBK^{+}16]. Quentin [Sin22]. Quest [Bok15, Kuu13, NP22, Rot24, Sch17b, Ueb18, Mac82, Sto13, Tau09]. question [Fan12, Ela13]. Questions [Cro16, Bru05, BZ14, Gim11, Cow12]. Quine [Cre20, Ade15, FA11, Gub19, Lav20, Ver20b, Smi24].

R [And15, Gar24, Gie12, Hil12, Hof13, Lun10, Nic12, Row15, Sim11]. Rabouin [Eck18]. Race [MRR17, Ste19b, Ueb21a]. Racialized [HES^{+}24]. Radical [Hat15a, Hol23]. Radick [Tab11]. Ragnar [Her19]. Raison [Bre11, CL12]. Ramus [Ans15]. Rand [Mih24]. Raphael [Büt20, Mun12, Mar10]. Rather [Her19]. Rational [Sha21]. Rationalism [Bsc15, Sep24]. Rationalist [Due23]. Rationalists [FPS11, Gre15]. rationality [EKD^{+}13, Rei14a]. Raum [Fee18]. Raz [Lüt14]. Razi [Koe24]. Reader [Cau20, MV18, Pat14, PK08, Lan19, Rey15]. Readers [Fab24]. Reading [BF17, Dom13, Gio24, Mos14, God10, Büt20, Gui18]. readings [Rus09]. Real [Sou13]. Realism [Ain12, BC17, Çim19, DP19, Hat15a, Iva13, Kle15a, Neu20, Wes15, Wra21, Cao10, Wüt14]. Realism/Antirealism [Wra21]. Realistic [Ban14, Lan16]. Reality [BV11, Hud20, Neu20, NP22, HMJ15, Rei06, REG08, Roc10, Pad11]. Really [Hud20, Pal15]. Reason [DP19, Gol11, Gre13, How22, Kas24, NZ10, Rei14a, Wel23, Bre11, EKD^{+}13, Sar10]. Reasoning [Hun14]. Reassessing [Mar17]. Rebecca [Rei14a]. Received [Lut12b]. Reception [Buz21, Iva20, Ver20a, Mel17]. Reck [Ser21]. reconceived [Ban14]. Reconsidered [Cap20]. Reconsidering [Pat11]. Record [Gri13, Sep12]. Reductionism [Bok08]. Reexamining [Bok08, Hep11]. Reflections [Dam17, Iva20, RD16, Tes16]. Reform [Edd15, Mor24, Pow12]. Reformation [Mil20]. Refraction [Dik22]. Rega [Pic16]. Regimens [Cor11, Jon13]. Regius [Dob20a]. Reich [Ste19b]. Reichenbach [Pad11]. Reisch [Tub20]. Rejecting [Ueb21b]. Rejection [Sou13]. related [Nie14].

Relation [Bok08, Coz24, Hep11]. Relationship [SB20, Tim19]. Relativism [DFK20]. Relativitätstheorie [Gio24]. Relativity [Gio24, Sta16, Slo24]. Religion [Hat15b, Kle15b, Bor11, Bre18, GP14, Gim12, Mar14, Ros13]. Religious [Mun16, LM15]. Remarks [Rhe24]. Renaissance [Ekh13, Mar11, Mun12, BM12, Hir11, Mar10, Rut21, Sga17]. Renaud [Eck18]. Renée [Büt20]. Renouvier [Sch11b, Sch17a]. Repercussion [Gon14]. Representation [De 22, Gro15, Pad11, Rei06, REG08]. Representational [Neu18]. Representing [Bon16]. Reproduction [Agi24, Hop24]. Republic [JM12, Duc14]. Rereading [Sep12, Gri13]. Response [Vat18]. Restraints [Ear13, Bor10]. Rethinking [Irv12]. Review [ASF^{+}21, Ada11, And15, AES^{+}24, Ant11, Ant13, Ara19, Are12, Ari17, Art14, Atk24, Bac13, Bal19, Bee24, BKV^{+}22, BM12, Bel13, Ben13, BV11, Bia16, Bla12, Bog13, Bok15, Bon16, Boy13, Bre14, Bre18, Bre24, Bro24, BS13, Bsc19, Büt20, Cal15, Cam20, Cap20, Car16a, Car16b, Car24, CL12, Cat12, Con24a, Con24b, Cow12, Cre20, Cro16, Dam17, DeS12, DeS14, DeW17a, Del19, DP19, Det15, Det16, Dob20a, Dom11, Dom13, Don24, Dor20, Dow17, Doy24, DL18, Dro11, Duc12, Duc14, Duf11, Dun24, Dun11, Dun14, Ear13, Eat21, Eck18, Edd15, Ekh13, Ela13, Fal17, Fal18, Fal16, Fal21, Fee18, Feo11, FES^{+}23, Fog12, Fra13, FS12, Fut14, Gab24, Gam15, Gar15a, Gar24, Gar15b, Gaw11, Geo19, Gie12, Gie24, Gig16, Gol11, Gol16]. Review [Gol17, Gor11, GBV^{+}22, Gre13, Gre15, Gre20, Gri13, Gro20, Gui18, Gui14, Haa13, Ham24, Har14, Hat12a, Hat15b, Hat20b, Hen13, Hen24, Hep11, Hic12, Hil12, Hof13, Hoq12, HES^{+}24, Ito14, Iva17, Jes14, Jha12, Jon13, Kas24, Kat12, Ken18, Ken22, Kin18a, Kin16, Kis24, Kjæ11, Kle12, Kob14, Kor18, Kra22b, Kuu13, Lær18, Lag14, Lan19, Lan16, Lat11, Lat14, Lau24, Lav20, Leg24, Ler11, Lol15, Lon23, Loo11, Lov16, Lüt14, Lut24, Mag12, Mar17, McC15, McC24, McD14, Meh14, Mel17, Men15, Mil14a, Mil14b, Mil20, Mil15a, Mir17, Mir11, Mon15, MP18, Mor24, Mos14, Mun12, Mun16, Nas16, New14, Nic12, Nor18, Obe14, Pad11, Pal15, Par11, Pea14, Pea19a, Pea19b, Pen13, Pen17, Pet18, Pic16, Pod14, Pop16, Pra13, Pur15]. Review [Rag19, Rap20, Rat24, Rei14b, Rei14a, Rem13, Ren18, Rey15, Ric13, Ric15, Rod24, Rog15, Roq19, Ros13, Rot17, Row15, Roz13, Rus19, RES^{+}24, Ryc13, Ryn14, Sal12, San22, Sar17a, Sar17b, Sau17, Sch14a, Sch13, Sch21, Seg14, Ser21, She11, SR18, Sim11, Sin19, Sin22, Slo11a, Slo24, Sme14, Smi11a, Smi24, Sna11, Spi18, ST16, Sta11, Ste19a, Ste19b, Ste13, Str18, Stu11c, Tab11, Teu15, Tub18, Tub20, Tur24, Tur11, Ueb16, Ueb18, Ueb21a, Van11, Ver20b, Vio24, WG12, Wüt14, Zam24a, Zam18a, Zam24b, Zup12, dP16, vS15]. Reviews [Jes11, Nem12]. Revision [IW18]. Revisited [Dam20, Obe14, KA12]. revolution [Sar10, Bon16, Gol11]. Revolutions [Cau24, Dam17, Kin18a, Kuh62, Kuh70, Kuh96, KH12, Mar15, Par11, KA12, Wei08, Obe14, Ric13, RD16, Wra16, PK08]. Rey [Bre21b]. Reyes [Kas24]. Rheinberger [Cam20, Hoq12, Pen13, Pra13]. Riccardi [Rod24]. Riccardo [Zam24a]. Richard [Don24, Don24, Gaw11, Kob14, Meh14, Sar17a, Slo24, Vio24]. Richards [Dam17, Kjæ11, Kob14, Pen17, Rus19]. Richardson [Doy24]. Riddle [Sch13, Rus07]. Riehl [Bia14, Mee24]. Riemannian [Gio17]. Rise [Jes11, Jes14, KV22, McC15, PS23, Zam18a, Gau10, Sga12]. Risk [ER17, Rus19]. Rita [Bre24]. Ritter [Are12, Hol10, RH10]. Road [Kle18, The20]. Robert [Eat21, vS15, Dam17, Kas24, Kjæ11, Pen17]. Robino [Eat21]. Robinson [Atk24]. Rocke [BV11]. Roger [Don24, Det16, Roz13]. Rogers [Loo11]. Rohault [Dob20b]. Roi [Gro20]. Role [Bre21a, Cok20, Pea18]. Rome [GP14]. Ronald [Hic12, Hof13]. Roots [Mur24]. Rose [Mih24]. Rothacker [Ste20]. Rougier [DS21]. Routledge [BKV^{+}22, DP19, Dor20, Lut24, Rag19, Roq19, LH16]. Roux [BS13, Jes14]. Roy [Gim21, dP21]. Royce [Gan21]. Rudolf [Car24, Doy24, Car24, Dew17b, Ren18]. Rule [BKV^{+}22, Dik22]. Ruler [Bee24]. Rupnow [Ste19b]. Ruse [Pen17, Hon18]. Rüsen [SR18]. Russell [Del19, Sch13, Ban14, Lan16]. Ruth [Ant11]. Ruysch [Koo10, WG12].

S [Atk24, Dor20, Esa19, Rat24, Ric13, Tub20]. Sachs [Sin19]. Salisbury [Mon15, Blo12]. Samuel [Don24]. Sander [Kis24]. Sanders [Pea18]. Sandner [Mor24]. sanitate [Web24]. Santayana [Neu20]. Santayana/Strong [Neu20]. Sarasohn [Gol11]. Sarukkai [Rag19]. Sauter [Mil20]. Savant [Pal15, Hol13]. Saving [Çim19]. Scale [BS22]. scepticism [Bet10, Meh14]. Scheiner [She11]. Schelling [Fol22, Ste19a, Zam18b]. Schemata [Wil20]. Schickore [Geo19]. Schiemann [Nor18, Stu11c]. Schiemer [Ser21]. Schliesser [Cal15, Hen24]. Schlotter [Doy24]. Schmaltz [Det16, Hat20b, Lol15, Slo11a, Spi18, dP24]. Schmaus [Pea19a]. Schmid [Hen13]. Schneider [Mee24]. Scholars [Hol13, Pal15]. Scholarship [Jes14]. scholastics [Gre91, Ari99, Ari11, Roz13]. School [Omo21, Vra20]. Schriften [Nor18]. Schroeder [Don24]. Science [ATE11, And13, Are12, BM16, BC17, Bre21a, Bre21b, Bre24, Bro12, Cap20, Chi08, CJ21, Cob12, CP21, Dam14, Dam17, De 24, DeS12, Del17, DP19, Dob20a, Dow17, Dro11, Duc12, Duf11, ER17, Fol19, FMB12, Fut14, Gaf10, Gam15, Gor21, Gra13, Gui14, Iva17, Jes11, Jes14, Jew12, KV22, Kin18a, Kle12, Kle15b, Kri11, Lær18, Lan19, Lat11, Lev15, Lon23, Lov16, Lut12b, Mac82, Mar15, MV18, Mel19, Mil16, Mil15a, Mir11, MK13, Nic11, Pal15, Pea19a, Pen18, Rei20, Rey15, Ric12, Rom12, Rom24b, Ros13, Ros22, Row15, Sch14a, Sch18, Sch24, Sch11b, Sch17a, Sch21, STP15, Sta11, Stu21, Stu16, Tau09, Tim19, Tur24, Ueb17, Wel23, Wes24, Wra19, Zam24a, ABS10, BJ09, Bal13, Blo12, Bor08, Bor11, Bou13, Bre18, Bru05]. science [Car18, DeW04, Ebr15, FN06, GCM12, GG08, Gau10, Gib15, Gim12, HNS11, Hol10, KM11, Ler10a, Ler10b, Lüt12a, Mar14, Mas08, MAKA09, Nye11, Pat14, RD16, RH10, Stu15a, Ter10, Van13, Wei08, Yeo14, dP14, MCP19, Dor20, Kin16, Kuu13, Mag12, Mil14a, MP18, Nas16, Rag19, Rei14b, Rus19, Seg14, DeS14, Gar24, Gaw11, Hat12a, Hat15b, Hic12, Jha12, Lüt14, Mon15, Par11, Pop16, Sar17a, Seg14, Ste13, Ueb18, vS15, Ler11, Ren18]. Sciences [AES^{+}24, Böh06, Bro24, Büt20, CCR16, Gig16, Har14, HS09, CD13, Isa12, Ler10a, Ler10b, NS14, Smi11b, ASF^{+}21, Eck18, Gar15b, Stu11c]. Scientia [SRK09, Loo11]. Scientific [BJ09, BV11, Bon16, Boy13, Bsc19, Cow12, CP21, Dam17, DP19, Dom13, DL18, Fab24, Feh21, Fra12, Gol11, Iva13, Jes11, Ken16, Ken18, Kuh62, Kuh70, Kuh96, KH12, Nic12, Obe14, Pop14b, Rei20, RD16, Ric13, Rom12, Sal12, Tes16, Toa13, Vae21, Ver24, Vio13, Wra16, Ach13, Bru05, DL11, Fis17, GG08, Gim11, GLB12, Har11, KA12, Lun10, PK08, Roc10, Sar10, Bre11, Ste13]. Scientifically [Geo19]. scientifique [Bre11, CL12]. Scientism [Vra20]. Scientists [Nas16]. Scott [Ade15, Ken18]. Scotus [Cro16, BZ14]. Scribal [Büt20]. Sean [Smi24, Ver20b]. Search [Gar15b, CD13]. Sebastian [Bia16, AES^{+}24]. Second [Del19, JM17, Pec21, Sch21, Vat18, Del12, Mih24]. secondo [Feo11, Tos09]. Secrets [Sar17b, Nie14]. Section [BV22, MV24]. Seen [Ben13, HA12]. Selected [Kis24, KK60]. Selection [Mét24, Bor10, Ear13]. Self [Gan21]. Sellars [Mat24, Dou24, Gim21, Hat15a, Neu20, Ole16]. Sellers [Dor20]. Semantic [Det18]. Sense [Gro20, Kjæ11, Ric08, Mur24]. Senses [SDR^{+}21]. Sensibility [Jes11, Pur15, Gau10]. Sensitive [Sha21]. Sentence [Cat12, Ueb07]. Sepkoski [Gri13]. Set [Ver20b]. Settecento [Tos09, Feo11]. Seventeenth [And15, DeS12, Loo11, Mar17, Mai14, ATE11, Jes14, SRK09]. Seventeenth-Century [And15, Loo11, Mai14, Jes14, SRK09]. Sexless [Hop24]. Sexual [Gab24]. Sgarbi [McC15]. Shank [Hic12]. Shape [Dik22, Wil20]. Shaped [Hop24]. Shaping [Jes11, GG08, Gau10]. Shared [Gub19]. Shaw [Ken22]. Shifting [ES17, ST16]. Shifts [Ole16]. Sibum [Lat11]. Sidney [Mil14b]. siècle [Bor11]. Sieroka [Ryc13]. Sierpiński [The20]. Sight [Nas16]. Sigmund [Ueb18]. Significance [HS09, Stu11c, Ueb11, Fal18]. Single [TF24]. Sir [Pen18]. Sitter [BS22]. Sixteenth [LH16, Roq19]. Skepticism [And15, Sch13, Mai14, Rus07]. Sketching [Rap20]. Slime [Str18]. Slipher [Fra22]. Slopes [Tub21]. Slowik [Mar17]. Smith [Duc12, Gig16, Gre15]. Snake [Geo19]. Sneedified [Dam14]. Snezana [Mun16]. Soames [Ade15]. Social [Jha12, Sch11b, Vra20, Nye11]. société [Bou13, Gui14]. society [Bou13]. Sociology [HES^{+}24, Kus19, Tur24, Vae21]. Socrates [Mil14a, Gra13]. solari [Rig66]. Soler [Kin16]. Solinas [Gol16]. Solipsism [Ueb21b]. Some [Wes15]. Sophie [BS13, Jes14]. Sophisticated [Sch21]. Sorana [Jon13]. Sorell [Loo11]. Soul [Dem21, Reg21, Bon13, Hir11, Dun24, Ekh13, Ito14]. Souls [Nob19, Bal19]. Sources [Cap20, Ler10a, Nem12]. Space [Bon16, Mee24, Slo11b, Slo24, Vin15, Car18, Car16b, GBV^{+}22]. Spatial [Mil20, Slo11b]. Spatiotemporal [Bei22]. Speaking [Sch21]. Special [BV22, DRW21, MV24, SR21, Sta16]. Species [Dil13, Kob14, Ric10]. Specificity [Vio13]. Spectroscopy [Fra22]. Speculations [vS22]. Speculative [KV22]. Spinoza [Bog13, Bal12, Har22, dP24]. Spirits [Rus21]. Spontaneous [Str18]. Springer [Ara19, Cap20, Del19]. Spur [FES^{+}23]. Stadler [Haa13, Ueb16]. Staffan [Cam20, Pen13]. Stahl [Ste19a, Pec21, Zam18b, dC21]. Stalled [Red24]. Stamp [Gol16]. Standards [ST16]. Starry [GD90]. State [Kra22a]. States [RES^{+}24]. Status [Rei20]. Steady [Kra22a]. Steady-State [Kra22a]. Stefano [Hat20b, Sch21]. Steffen [Dom13]. Steinke [Pal15]. Steinthal [Rei20]. Stengers [Fra13]. Stephan [Hen13]. Stephen [Dun24, Jes11, Mel17, Jes11]. Stetson [Atk24]. Steven [Cow12, Ros13]. Stewart [Con24b]. Stimuli [Gub19]. Stone [Bok15]. storia [Feo11, Tos09]. Stories [Cok20, Stu21]. Story [Mih24]. Strange [Rei14a, EKD^{+}13]. strategies [Ach13]. Straw [Lut12b]. Strazzoni [Dob20a]. Streijger [Zup12]. Strength [Vio24]. Strobino [Zam24a]. Strong [Neu20]. Structural [Ain12, BC17, Wüt14, Cao10]. Structuralism [Ser21]. Structure [Dam17, Koc13, Kuh62, Kuh70, Kuh96, KH12, NP22, Obe14, Ric13, Tim13, KA12, PK08, Tim19, RD16, Wra16]. Struggle [Çim19, Kjæ11, Tub20, Ric08]. Stuart [Mag15]. Stuber [Pal15]. Studien [Fee18]. Studies [Cap20, ER17, Rus19, GP14, Ter10, SR18]. study [Sch10]. Stump [Iva17]. Sturm [Rei14a]. Styles [KM11, Sch14a]. Suárez [dP24]. Subject [Gar24]. Substance [Smi11a, Gar09]. Substantial [DP19]. Substanzbegriff [Cas10, Hei14]. Subverting [Mar14, Hat15b]. Succeed [Pal15]. Successful [Tim13]. Sufficient [Wel23]. sulle [Rig66]. Sumus [Sid24]. Sun [Ben13, HA12]. Sundar [Rag19]. sunspots [GRVS10, Rig66, She11, GD90]. SUNY [Tub20]. super [SBT10, Zup12]. Surroundings [Sie10]. Susannah [Lov16]. Sven [Doy24]. Swabian [Sto13, Bok15]. Syllogism [McC24]. Symbolic [De 22, Sie10]. Symbolischer [Ryc13, Sie10]. Syncategoremania [BBK^{+}16]. Synthesis [Ito14, Bon13]. Synthetic [Fra12]. System [Koc13, Wil20]. système [Bre18, Pet16].

T [Ari17, Bro24, Don24, Gol11, Slo24]. Tad [Det16, Hat20b, dP24, Lol15, Slo11a, Spi18]. Tagebücher [Car24]. Talisse [Kas24]. Tamas [AES^{+}24, Cap20, Lau24]. Tammy [Gar15a]. Tano [Rat24]. Tapestry [MP18]. Tarek [Gar24]. Tauber [Kuu13]. Taxonomy [Mag15]. Technologies [Büt20]. Ted [Rus19]. teleological [Sch10]. teleologischer [Hen13, Sch10]. Teleology [De 22, Gil14, Gol16, GLB12, Zam18a, Boy13]. Telling [Stu21]. Tendency [Rat24]. Teresa [Ler11]. Tessicini [Rot17]. Testimonies [RES^{+}24]. Tetens [D’A24]. Texte [Ueb16]. textes [Bre15, dP16]. Texts [Are12, Hol10, Nie14, RH10]. Textual [Gre20]. Thalheimer [Mel19]. Their [Llo13, Mun16, LM15]. Theo [Det16]. Theodore [Obe14]. Theology [Mar21]. Theorem [Kle18]. Theoretical [Bsc15, Del19, JM17]. théorie [Car18, Ren18]. Theories [Tes16, Wes24]. Theorist [War24]. Theorists [Eri15, Mir17]. Theory [Bsc15, Ebr15, ES17, Esa19, Gio24, GW14, Kri12, Kub18, Men15, Neu18, SR18, Stu11b, Zam24a, Car18, Tim19, Pop16, Ver20b]. There [Wue13]. Thermal [Ant22]. Thermodynamics [Del12]. Thesis [Ade15]. Thijssen [Zup12]. Thing [MCP19]. Things [Lag14, KW10]. Think [Toa21]. Thinkers [Loo11, SRK09]. Thinking [Ade15, Gub19, HMJ15, NL12, Sch24, Ueb18, Böh06, FS12]. Third [Ain12, BKV^{+}22]. Tho [Don24]. Thomas [Rei14a, Tub20, Car16b, Cat12, GP14, Kin18a, Lut24, Mar15, Mel19, Ric13]. Thought [Buz18, Buz21, D’A24, Feh21, FMB12, Her19, Pod14, SF21, Stu21, LNN12, NL12, Ric08, BS13, IR11, Row15, Car16a, FS12, Kjæ11]. Thoughts [Pec21]. Thrasybulus [Web24]. Three [Stu16, Nie14]. Thrice [Sha21]. Tian [Wüt14]. Tilburg [Kis24]. Tilburg-Groningen [Kis24]. Timaeus [Gre20]. Time [BZ14, Can16, Cro16, Mee24, Slo24, Bar13, Fal17, Rat24, Ryn14, Sau17]. Times [Lag14, Sch14a, Ueb18, KM11, KW10]. Timothy [Mir11]. today [Mas08, Sta11]. Together [Dom13]. Tom [Loo11]. Topic [MCP19]. Topic-Modeling [MCP19]. Toscano [Feo11]. Touati [DeS12]. Touch [SDR^{+}21]. Tradition [Gre15, Gre20, McC15, McC24, Mos14, Sha20, Cor11, FPS11, Gee13, God10, Sga12, Jon13, Kas24, Lat14]. Traditions [Gio17]. Tragic [Kjæ11, Ric08]. Trans [Cam20, Bee24, Don24]. Transcendental [Car16b, Vin15]. Transformation [Hen13, Sch10]. Transformations [Haa13, Man09]. Transit [Ben13, HA12]. Transition [Gre21, OS17, Sin19]. Translated [Dun24]. translation [dCH11]. Treatise [Mel17, Sch13, Rus07]. Trevor [Leg24]. Trizio [Kin16]. Truth [Pec21, Bru05]. Trying [Pal15]. Tuboly [AES^{+}24, Cap20, Lau24]. tuenda [Web24]. Turn [Dam14, Gre13, NZ10, PS23, Toa13, Zam24b]. Turning [Ant11, Dom13, Gla09, Har11]. Turns [Sch18]. Twentieth [BM16, Bro24, Hoq12, ST16, Rhe10, Vio13]. Twentieth-Century [BM16, Hoq12, Rhe10, Vio13]. Two [Büt20, FA11, Neu20, Sar17b, Nie14]. Tyler [Bia16]. Typology [Sch18]. Tzuchien [Don24].

Uebel [Cat12, Lut24]. Uljana [Tur11]. ultimate [Bru05]. Umgebungen [Ryc13, Sie10]. Uncertainty [Rot24, AP22]. Unconscious [FS12, NL12]. Undecided [Bsc19, Fis17]. Underdetermination [Ade15, Bei22, Har22]. Understanding [Agi24, Can16, Lau24, Sau17]. Unexpected [Wes15]. unification [vD10, Fog12]. Unifying [GBT17, Rot17]. Unity [Dun17]. Universal [BKV^{+}22, Fer22]. Universals [Hat20b]. Universe [Bei22, Hop24, Kra22a]. University [Bsc19, Büt20, Cam20, Cre20, Dun24, Geo19, Gro20, Hat20b, Mil20, MP18, Pea19a, Pea19b, Rap20, Rei14b, Rus19, Ste19a, Ver20b, Jew12]. Untersuchung [Hen13, Sch10]. Untersuchungen [Cas10]. Use [Hun14]. Utopianism [Fer24].

V [Rom24a]. Vain [GS17]. valeurs [Bre11, CL12]. Valiant [Bok15, Sto13]. Validity [Gub19, Vio24]. Values [ER17, MP18, Rus19, Bre11]. VandeWall [Lan19]. Vasso [Obe14]. Vegetable [Bal19, Lov16, Gib15]. Venom [Geo19]. Venus [Ben13, Ben13, HA12]. Verbeek [Det16]. Verborgene [Böh06, Kat12]. Verborgener [Kat12, Böh06]. Verburgt [McC24]. Verhaegh [Kis24]. Verstehen [Fee10, Tur11]. versus [BBK^{+}16, Mét24]. Vesto [Fra22]. Vienna [Cat12, Haa13, Ueb18, Fer24, Ler10a, Man09, Rom12, Ueb07, vS22]. Vienne [Ler10a, Nem12]. Viennese [Rad13]. View [Bre21a, Fer22, Lut12b]. Views [Rut21]. viii [Sin19]. Vincent [Sim11]. Vinci [Car16b]. Virtual [Duf11, Gaf10]. Virtuality [Rat24]. Virtue [Due23, Gre13, NZ10]. Virtuosi [Sar17a, Yeo14]. Vital [Zam18a]. Vitalism [Rus21, Bro24]. Vol [Car24, Doy24]. Volker [Ric15]. Völkerpsychologie} [DFK20, Kus19, Edg20a, Rei20, Ste20]. vols [Pal15]. Volume [Ler10a, Ler10b, Ler10a, Ler10b].

W [Don24, Doy24, HES^{+}24, Pen18, Rap20, Slo24]. W. [Sna11]. Wagerklärungen [Sch24]. Wagner [Gro20, Sin22]. Waldow [Car16a]. Wallis [Lag14]. Walter [Gor11, Mil16]. War [Mih24, Rei14b, Rei14a, Tub20, EKD^{+}13, Jew12, Rei14b]. Warren [Pea19a]. Warriors [Mil15a]. Was [Duc14, Hen23, Hud20, Pec21, Tim13, JM12]. Water [Kri12]. Waters [Mar17]. Watkins [Det15, Men15]. Waxman [New14]. Way [Rot24]. Wayne [New14]. Weinert [Par11]. Wells [Fer24]. Werner [Mor24]. West [Rag19]. Weyl [Ryc13, Sie10, McC22, Sie10, Toa13, Toa21]. Whewell [Det18, Hon18, Ver24, War24]. While [Pal15]. Whitehead [HMJ15]. Wiener [Ueb16]. Wilfrid [Dou24, Ole16]. Wilhelm [Are12, Bee24, Hol10, Kin18b]. Will [Gre13, Kle15b, Mis15, NZ10]. Wille [Cam20, Pen13]. William [Dom13, Sim11, Ada11, Bor08, Det18, Kle12, Seg14, War24]. Wilson [Bla12, Zam24b, Bee24]. Wingate [Con24a]. Wisdom [Lær18, Lev15]. Wise [Sha21]. Wissen [Ara19, Gro15]. wissenschaftlichen [FES^{+}23]. Wittgenstein [Bog13, Rog15, Bal12, Mad14]. Wolfe [Ari17, Bro24]. Wolff [Wel23]. Wolfgang [Doy24]. Women [And13, Wue13, Kis24]. Wonder [Str18]. Wood [Gim21, Pic16]. Woolgar [Fab24]. Woolhouse [Don24]. Work [Ric12]. Working [Isa12, Har14]. Works [Bac13, BB12, Doy24, Ole16, CBCD12]. World [Eri15, How22, Irv12, Koc13, McC22, Mih24, Mir17, Rom12, dP24, GP14, Tos09, Mil15a]. Worldviews [DeW04, Gaw11]. Wray [San22]. Wrestling [HNS11, Hic12]. Writing [Chi08, Geo19, Dro11]. Writings [Mar21, NZ10, KK60, Doy24, Gre13, Web24]. Wronski [Wag14]. Wundt [Hat20a].

X [GP14, Ver20b]. Xavier [NP22]. xi [BKV^{+}22, Cre20, DP19, Rag19]. xii [Ara19, Cap20, Dor20, Lav20]. xiii [Ste19b]. xiv [Pea19a, Rus19]. xlv [Tub20]. xviii [Gre20]. XVIIIe [Bor11]. xxiii [Lan19]. xxx [Pal15]. xxxi [Del19]. xxxiv [Dun24].

Yeo [Sar17a]. Yes [Wue13]. York [BKV^{+}22, Lav20, Roq19, Rus19]. Young [Bog13, Bal12]. Youth [Mor24]. Yu [Wüt14].

Z [Sme14]. Zahl [Böh06, Kat12]. Zahl-Verborgener [Kat12]. Zammito [Ste19a]. Zauderer [Gre13]. Zilsel [Rom24b]. Zubiri [NP22]. zum [Ueb16]. zur [Fee18, FES^{+}23, Gio24]. Zvi [Cal15].

References

Aubin:2010:HEO

[ABS10] David Aubin, Charlotte Bigg, and Heinz Otto Sibum, editors. The heavens on earth: observatories and astronomy in Nineteenth-Century science and culture. Science and cultural theory. Duke University Press, Durham, NC, USA, 2010. ISBN 0-8223-4628-1 (hardcover), 0-8223-4640-0 (paperback). xii + 384 pp. LCCN QB32 .H43 2010.

Achinstein:2013:EMS

[Ach13] Peter Achinstein. Evidence and method: scientific strategies of Isaac Newton and James Clerk Maxwell. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-992185-7 (hardcover), 0-19-933289-4 (e-book). xv + 177 pp. LCCN Q174.8 .A238 2013.

Adams:2011:BRW

[Ada11] Marcus P. Adams. Book review: William Poole, John Aubrey and the Advancement of Learning. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):375–377, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/660742.

Adeel:2015:EQT

[Ade15] M. Ashraf Adeel. Evolution of Quine’s thinking on the thesis of underdetermination and Scott Soames’s accusation of paradoxicality. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):56–69, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/680369.

Ariew:2015:HDD

[ADJ^{+}15] Roger Ariew, Dennis Des Chene, Douglas Michael Jesseph, Tad M. Schmaltz, and Theo Verbeek. Historical dictionary of Descartes and Cartesian philosophy. Historical dictionaries of religions, philosophies, and movements. Rowman and Littlefield, Lanham, MD, USA, second edition, 2015. ISBN 1-4422-4768-1 (hardcover), 1-4422-4769-X (e-book). xx + 388 pp. LCCN B1831 .H57 2015.

Adomaitis:2019:CEL

[Ado19] Laurynas Adomaitis. Cause and effect in Leibniz’s Brevis demonstratio. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):120–134, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anta:2024:BRS

[AES^{+}24] Javier Anta, Mehmet Elgin, Elliott Sober, Thomas Uebel, Daniel Schwartz, Pierre-Olivier Méthot, Daniele Cozzoli, and Hans-Jörg Rheinberger. Book review: Sebastian Lutz and Adam Tamas Tuboly, eds., Logical Empiricism and the Physical Sciences: From Philosophy of Nature to Philosophy of Physics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):560–563, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Agiriano:2024:JUR

[Agi24] Arantza Etxeberria Agiriano. Jacob’s understanding of reproduction: Challenges from an organismic collaborative framework. HO-POS: Journal of the International Society for the History of Philosophy of Science, 13(2):535–553, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ainsworth:2012:TPS

[Ain12] Peter M. Ainsworth. The third path to structural realism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):307–320, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/664775.

Alexander:2010:BID

[AN10] Denis Alexander and Ronald L. Numbers, editors. Biology and ideology from Descartes to Dawkins. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-60840-9 (hardcover), 0-226-60841-7 (paperback), 0-226-60842-5 (e-book). 453 pp. LCCN QH331 .B477 2010.

Andersen:2013:WHP

[And13] Hanne Andersen. Women in the history of philosophy of science: What we do and do not know. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):136–139, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/669954.

Anderson:2015:BRO

[And15] Joseph Anderson. Book review: José R. Maia Neto, Academic Skepticism in Seventeenth-Century French Philosophy: The Charronian Legacy, 1601–1662. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):371–374, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/682378.

Anonymous:2011:EM

[Ano11a] Anonymous. Ernan McMullin, 1924–2011. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1 (1):iv, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-

6240 (electronic). URL http://www.jstor.org/stable/10.1086/
660690.

Anonymous:2011:E

[Ano11b] Anonymous. From the Editors. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):1–2, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/659251.

Anonymous:2016:Ma

[Ano16a] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):??, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2016:Mb

[Ano16b] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):??, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2017:Ma

[Ano17a] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):??, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2017:Mb

[Ano17b] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):??, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2018:Ma

[Ano18a] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):??, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2018:M

[Ano18b] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):??, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2019:Ma

[Ano19a] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):??, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2019:M

[Ano19b] Anonymous. Masthead. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):??, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2020:FMa

[Ano20a] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):??, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2020:FMb

[Ano20b] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):??, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2021:FMa

[Ano21a] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):??, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2021:FMb

[Ano21b] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):??, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2022:EN

[Ano22a] Anonymous. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):??, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2022:FMa

[Ano22b] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):??, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2022:FMb

[Ano22c] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2):??, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2023:FM

[Ano23] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):??, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2024:FMa

[Ano24a] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):??, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2024:FMb

[Ano24b] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):??, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anonymous:2024:FMc

[Ano24c] Anonymous. Front matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):??, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Anstey:2011:JLN

[Ans11] Peter R. Anstey. John Locke and natural philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-958977-1. x + 252 pp. LCCN B1298.

Anstey:2015:FBL

[Ans15] Peter R. Anstey. Francis Bacon and the Laws of Ramus. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):1–23, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/677037.

Anthony:2011:BRR

[Ant11] Sean W. Anthony. Book review: Ruth Glasner, Averroes’ Physics: A Turning Point in Medieval Natural Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):175–178, Spring 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/656813.

Anthony:2013:BRB

[Ant13] Sean W. Anthony. Book review: Godefroid de Callataÿ and Bruno Halflants, eds., Epistles of the Brethren of Purity: On Magic I. HO-POS: Journal of the International Society for the History of Philosophy of Science, 3(2):384–387, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/671108.

Anta:2022:PAB

[Ant22] Javier Anta. A philosopher against the bandwagon: Carnap and the informationalization of thermal physics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):43–67, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Atkinson:2022:HCH

[AP22] David Atkinson and Jeanne Peijnenburg. How certain is Heisenberg’s Uncertainty Principle? HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):1–21, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Aray:2019:BRA

[Ara19] Basak Aray. Book review: Angélique Groß, Die Bildpädagogik Otto Neuraths: Methodische Prinzipien der Darstellung von Wissen. Cham: Springer, 2015. Pp. xii + 288. EUR 89.99 (cloth). HO-POS: Journal of the International Society for the History of Philosophy of Science, 9(1):198–201, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Arens:2012:BRJ

[Are12] Katherine Arens. Book review: Jocelyn Holland, Key Texts of Johann Wilhelm Ritter (1776–1810) on the Science and Art of Nature. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):169–172, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/661974.

Ariew:1999:DLS

[Ari99] Roger Ariew. Descartes and the last Scholastics. Cornell University Press, Ithaca, NY, USA, 1999. ISBN 0-8014-3603-6. xi + 230 pp. LCCN B1875 .A65 1999.

Ariew:2011:DAS

[Ari11] Roger Ariew. Descartes among the Scholastics, volume 20(1) of History of science and medicine library; Scientific and learned cultures and their institutions. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-20724-4 (hardcover). x + 358 pp. LCCN B1875 .A65 2011.

Arikha:2017:BRC

[Ari17] Noga Arikha. Book review: Charles T. Wolfe, Materialism: a Historico-Philosophical Introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2): 386–391, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Arthur:2014:BRB

[Art14] Richard T. W. Arthur. Book review: Klaas van Berkel, Isaac Beeckman on Matter and Motion. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):192–196, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675432.

Abadia:2021:BRC

[ASF^{+}21] Oscar Moro Abadía, Michael T. Stuart, Yiftach Fehige, Lydia Patton, and K. Brad Wray. Book review: Cristina Chimisso, Hélène Metzger, Historian and Historiographer of the Sciences. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):327–330, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ait-Touati:2011:FCS

[ATE11] Frédérique Aït-Touati and Susan Emanuel. Fictions of the Cosmos: Science and Literature in the Seventeenth Century. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 1-283-28157-0, 0-226-01122-4, 0-226-01124-0 (e-book). ix + 261 pp. LCCN PN56.C68 A5813 2011.

Atkins:2024:BRS

[Atk24] Richard Kenneth Atkins. Book review: Stetson J. Robinson, ed., The Correspondence of Charles S. Peirce and the Open Court Publishing Company, 1890–1913. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):219–222, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bacciagaluppi:2013:BRJ

[Bac13] Guido Bacciagaluppi. Book review: Jeffrey A. Barrett and Peter Byrne, eds., The Everett Interpretation of Quantum Mechanics: Collected Works 1955–1980 with Commentary. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):348–352, Fall 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/671743.

Baltas:2012:PPG

[Bal12] Aristeidēs Baltas. Peeling potatoes or grinding lenses: Spinoza and young Wittgenstein converse on immanence and its logic. University of Pittsburgh Press, Pittsburgh, PA, USA, 2012. ISBN 0-8229-4416-2, 0-8229-7790-7 (e-book). xvii + 290 pp. LCCN B3974 .B35 2012.

Ball:2013:CHS

[Bal13] Philip Ball. Curiosity: how science became interested in everything. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-04579-X (hardcover), 0-226-04582-X (e-book). viii + 465 pp. LCCN Q125 .B297 2013.

Baldassarri:2019:BRL

[Bal19] Fabrizio Baldassarri. Book review: Lucas John Mix, Life Concepts from Aristotle to Darwin: On Vegetable Souls. Cham: Palgrave Macmillan, 2018. Pp. 272. $85.00 (cloth); $65.00 (e-book). ISBN 978-3-319-96046-3 (cloth); ISBN 978-3-319-96047-0 (e-book).

HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):423–426, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Banks:2014:REM

[Ban14] Erik C. Banks. Realistic empiricism of Mach, James, and Russell: neutral monism reconceived. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-07386-3 (hardcover). viii + 217 pp. LCCN B816 .B36 2014.

Bardon:2013:BHP

[Bar13] Adrian Bardon. A brief history of the philosophy of time. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-997645-7, 0-19-930108-5 (paperback), 0-19-997773-9 (updf), 0-19-997774-7 (epub), 1-299-67459-3 (e-book). ix + 185 pp. LCCN BD638 .B335 2013.

Barrett:2012:EIQ

[BB12] Jeffrey A. Barrett and Peter Byrne. The Everett Interpretation of Quantum Mechanics: Collected Works 1955–1980 with Commentary. EBL-Schweitzer. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14507-5 (hardcover), 1-280-49440-9, 1-4008-4274-3 (e-book). xii + 389 pp. LCCN QC174.12.

Bascelli:2016:LVI

[BBK^{+}16] Tiziana Bascelli, Piotr Błaszczyk, Vladimir Kanovei, Karin U. Katz, Mikhail G. Katz, David M. Schaps, and David Sherry. Leibniz versus Ishiguro: Closing a quarter century of syncategoremania. HO-POS: Journal of the International Society for the History of Philosophy of Science, 6(1):117–147, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Brading:2017:ESR

[BC17] Katherine Brading and Elise Crull. Epistemic structural realism and Poincaré’s philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1): 108–129, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Beeley:2024:BRG

[Bee24] Philip Beeley. Book review: Gottfried Wilhelm Leibniz, Dissertation on Combinatorial Art, trans. Massimo Mugnai, Han van Ruler, and Martin Wilson. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):620–623, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Beisbart:2022:WSE

[Bei22] Claus Beisbart. What is the spatiotemporal extension of the universe? underdetermination according to Kant’s first antinomy and in present-day cosmology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):286–307, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bellis:2013:BRB

[Bel13] Delphine Bellis. Book review: Carlo Borghero, Les Cartésiens face à Newton. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):364–367, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/671467.

Bennett:2013:BRJ

[Ben13] Jim Bennett. Book review: Jeremiah Horrocks, Venus Seen on the Sun: The First Observation of a Transit of Venus by Jeremiah Horrocks. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):375–376, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/670134.

Bett:2010:CCA

[Bet10] Richard Bett, editor. The Cambridge companion to ancient scepticism. Cambridge companions to philosophy. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-87476-9 (hardcover), 0-521-69754-9 (paperback). xii + 380 pp. LCCN B525 .C36 2010.

URL http://catalogue.bnf.fr/ark:/12148/cb42183097h.

Buchwald:2012:NOC

[BF12] Jed Z. Buchwald and Mordechai Feingold. Newton and the origin of civilization. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15478-3, 1-283-85127-X (MyiLibrary). ix + 529 pp. LCCN QC16.N7 B93 2012.

Boran:2017:RNE

[BF17] Elizabethanne Boran and Mordechai Feingold, editors. Reading Newton in Early Modern Europe, volume 19 of Scientific and learned cultures and their institutions. Brill, Leiden, The Netherlands, 2017.

ISBN 90-04-33664-8 (hardcover), 90-04-33665-6 (e-book). ISSN 2352-1325. x + 358 pp. LCCN QA803 .R43 2017.

Biagioli:2014:HCA

[Bia14] Francesca Biagioli. Hermann Cohen and Alois Riehl on geometrical empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):83–105, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/675400.

Biagioli:2016:BRJ

[Bia16] Francesca Biagioli. Book review: J. Tyler Friedman and Sebastian Luft, eds., The Philosophy of Ernst Cassirer: a Novel Assessment. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):164–167, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bailer-Jones:2009:SMP

[BJ09] Daniela Bailer-Jones. Scientific models in philosophy of science. University of Pittsburgh Press, Pittsburgh, PA, USA, 2009. ISBN 0-8229-4376-X (hardcover), 0-8229-7123-2 (e-book). x + 235 pp. LCCN Q175 .B163 2009.

Belkind:2022:BRM

[BKV^{+}22] Ori Belkind, Joel Katzav, Krist Vaesen, Marij van Strien, and Ori Belkind. Book review: Mary Domski, Newton’s Third Rule and the Experimental Argument for Universal Gravity, New York: Routledge, 2022. Pp. xi + 116. $47.96 (cloth). ISBN 978-1-032-02036-5. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2):510–515, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Blank:2012:BRC

[Bla12] Andreas Blank. Book review: Catherine Wilson, Epicureanism at the Origins of Modernity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):200–203, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/663836.

Bloch:2012:JSA

[Blo12] David Bloch. John of Salisbury on Aristotelian science, volume 8 of Disputatio. Brepols, Turnhout, Belgium, 2012. ISBN 2-503-54099-6. xv + 1 + 243 pp. LCCN B765.J44 B56 2012.

Bellis:2012:BRC

[BM12] Delphine Bellis and Gideon Manning. Book review: Craig Martin, Renaissance Meteorology: Pomponazzi to Descartes. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):394–398, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/666475.

Ben-Menahem:2016:PIT

[BM16] Yemima Ben-Menahem. Poincaré’s impact on twentieth-century philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):257–273, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bogen:2013:BRA

[Bog13] Jim Bogen. Book review: Aristides Baltas, Peeling Potatoes or Grinding Lenses: Spinoza and Young Wittgenstein Converse on Immanence and Its Logic. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):352–356, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/671199.

Bohlandt:2006:VZV

[Böh06] Marco Böhlandt. Verborgene Zahl — verborgener Gott: Mathematik und Naturwissen im Denken des Nicolaus Cusanus (1401– 1464). (German) [Hidden number — hidden God and Mathematics: Natural Sciences in the thinking of Nicholas of Cusa (1401– 1464)], volume 58 of Sudhoffs Archiv: […], Beiheft; Wissenschaftsgeschichte. Steiner, Stuttgart, Germany, 2006. ISBN 3-515-09289-7. 358 pp. LCCN BX4705.N58 B64 2009. URL http://d-nb.info/

995315957/04.

Bokulich:2008:RQC

[Bok08] Alisa Bokulich. Reexamining the Quantum–Classical Relation: Beyond Reductionism and Pluralism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85720-1. x + 195 pp. LCCN QA805

.B685 2008. URL http://www.loc.gov/catdir/enhancements/
fy0834/2008018902-b.html;   http://www.loc.gov/catdir/
enhancements/fy0834/2008018902-d.html;   http://www.
loc.gov/catdir/enhancements/fy0834/2008018902-t.html.

Bokulich:2015:BRD

[Bok15] Alisa Bokulich. Book review: A. Douglas Stone, Einstein and the Quantum: The Quest of the Valiant Swabian. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):177–179, Spring 2015. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/678188.

Boner:2013:KCS

[Bon13] Patrick Boner. Kepler’s cosmological synthesis: astrology, mechanism and the soul, volume 39 of History of science and medicine library, Medieval and early modern science; volume 20. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-24608-8 (hardcover), 90-04-24609-6 (e-book). ISSN 1872-0684. x + 187 pp. LCCN QB36.K4 .B638 2013.

Boner:2016:BRD

[Bon16] Patrick J. Boner. Book review: David Marshall Miller, Representing Space in the Scientific Revolution. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):172– 173, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bordogna:2008:WJB

[Bor08] Francesca Bordogna. William James at the boundaries: philosophy, science and the geography of knowledge. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-06652-5. x + 382 pp. LCCN B945.J24 B47 2008.

Borrello:2010:ERC

[Bor10] Mark E. Borrello. Evolutionary restraints: the contentious history of group selection. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-06701-7 (hardcover). xi + 215 pp. LCCN QH376 .B67 2010.

Borghero:2011:CFN

[Bor11] Carlo Borghero. Les Cartésiens face à Newton: philosophie, science et religion dans la première moitié du XVIIIe siècle. (French) [The Cartesians against Newton: philosophy, science, and religion in the first half of the Eighteenth Century]. Brepols, Turnhout, Belgium, 2011. ISBN 2-503-54177-1 (paperback). 156 + 4 pp. LCCN B802 .B674 2011. Translated from the Italian by Tomaso Berni Canani.

Bourdeau:2013:ACS

[Bou13] Michel Bourdeau. Auguste Comte: science et société. (French) [Auguste Comte: science and society]. Philosophie en cours: séries générales. SCÉRÉN-CNDP-CRDP, Futuroscope, ????, 2013. ISBN 2-240-03423-8 (paperback). ISSN 2109-8786. 85 pp. LCCN ????

URL http://catalogue.bnf.fr/ark:/12148/cb43751843v.

Boylan:2013:BRA

[Boy13] Michael Boylan. Book review: Allan Gotthelf, Teleology, First Principles, and Scientific Method in Aristotle’s Biology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):387–390, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/671107.

Brenner:2011:RSV

[Bre11] Anastasios Brenner. Raison scientifique et valeurs humaines: essai sur les critères du choix objectif. (French) [Scientific reason and human values: an essay on the criteria of objective choice]. Science, histoire et société. Essais. Presses universitaires de France, Paris, France, 2011. ISBN 2-13-059055-1. viii + 125 pp. LCCN Q175.32.V35 B75 2011.

Brenner:2014:BRA

[Bre14] Anastasios Brenner. Book review: Auguste Comte, Michel Bourdeau, Laurent Clauzade, and Frédéric Dupin, eds., Oeuvres: Cours de philosophie positive, leçons 46-51. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1): 176–178, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-

6240 (electronic). URL http://www.jstor.org/stable/10.1086/
672153.

Brenner:2015:TFE

[Bre15] Anastasios Brenner. Les textes fondateurs de l’épistémologie françcaise: Duhem, Poincaré, Brunschvicg et autres philosophes. Collection “Hermann Philosophie”. Hermann, Paris, France, 2015. ISBN 2-7056-9090-5 (paperback). 293 pp. LCCN BD162 .T49 2015.

Brenner:2018:BRA

[Bre18] Anastasios Brenner. Book review: Annie Petit, Le système d’Auguste Comte: De la science à la religion par la philosophie.

HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):226–229, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Brecevic:2021:RIE

[Bre21a] Char Brecevic. The role of imagination in Ernst Mach’s philosophy of science: a biologico-economical view. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11 (1):241–261, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Brenner:2021:HDP

[Bre21b] Anastasios Brenner. How did philosophy of science come about?: From Comte’s positive philosophy to Abel Rey’s absolute positivism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):428–445, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bresci:2024:BRD

[Bre24] Simone Bresci. Book review: Delphine Bellis, Daniel Garber, and Carla Rita Palmerino, eds., Pierre Gassendi: Humanism, Science, and the Birth of Modern Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1): 256–259, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Brown:2012:JDL

[Bro12] Matthew J. Brown. John Dewey’s logic of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):258–306, Fall 2012. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/666843.

Bronstein:2016:AKL

[Bro16] David Bronstein. Aristotle on knowledge and learning: the posterior analytics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-872490-X (hardcover), 0-19-179242-X (e-book), 0-19-103791-5 (e-book). xii + 272 pp. LCCN B441 .B76 2016.

Brooks:2024:BRC

[Bro24] Daniel S. Brooks. Book review: Christopher Donohue and Charles T. Wolfe, eds., Vitalism and Its Legacy in Twentieth Century Life Sciences and Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):634–637, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Brush:2005:LST

[Bru05] Nigel Brush. The limitations of scientific truth: why science can’t answer life’s ultimate questions. Kregel Publications, Grand Rapids, MI, USA, 2005. ISBN 0-8254-2253-1. 300 pp. LCCN BL240.3

.B78 2005. URL http://www.loc.gov/catdir/toc/ecip0517/
2005022166.html.

Brown:2013:BRK

[BS13] James Robert Brown and Michael T. Stuart. Book review: Katerina Ierodiakonou and Sophie Roux, eds., Thought Experiments in Methodological and Historical Contexts. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3 (1):154–157, Spring 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/667764.

Biener:2014:NE

[BS14] Zvi Biener and Eric Schliesser, editors. Newton and empiricism. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-937399-X (e-book). LCCN QC16.N7.

Baerdemaeker:2022:BAS

[BS22] Siska De Baerdemaeker and Mike D. Schneider. Better appreciating the scale of it: Lemaître and de Sitter at the BAAS centenary. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):170–188, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bschir:2015:FPT

[Bsc15] Karim Bschir. Feyerabend and Popper on theory proliferation and anomaly import: On the compatibility of theoretical pluralism and critical rationalism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):24–55, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/680368.

Bschir:2019:BRM

[Bsc19] Karim Bschir. Book review: Menachem Fisch, Creatively Undecided: Toward a History and Philosophy of Scientific Agency. Chicago:

University of Chicago Press, 2017. Pp. 304. $27.92 (paper). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):189–192, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Buttner:2020:BRR

[Büt20] Jochen Büttner. Book review: Renée Raphael, Reading Galileo: Scribal Technologies and the “Two New Sciences”. Baltimore: Johns Hopkins University Press, 2017. Pp. 280. $54.95 (cloth). ISBN 978-1-4214-2178-0. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):350–353, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Buzzoni:2018:PDE

[Buz18] Marco Buzzoni. Pierre Duhem and Ernst Mach on thought experiments. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):1–27, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Buzzoni:2021:NCH

[Buz21] Marco Buzzoni. A neglected chapter in the history of philosophy of mathematical thought experiments: Insights from Jean Piaget’s reception of Edmond Goblot. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):282–304, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bensaude-Vincent:2007:ANE

[BV07] Bernadette Bensaude-Vincent. The artificial and the natural: an evolving polarity. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-02620-1. vi + 331 pp. LCCN Q127.E8 A78 2007.

Bensaude-Vincent:2011:BRA

[BV11] Bernadette Bensaude-Vincent. Book review: Alan Rocke, Image and Reality: Kekulé, Kopp, and the Scientific Imagination. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):329–330, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/661198.

Bianchi:2022:SSI

[BV22] Silvia De Bianchi and Federico Viglione. Special section introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):122–128, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Buckner:2014:DST

[BZ14] Edward Buckner and Jack Zupko. Duns Scotus on Time and Existence: the questions on Aristotle’s “De interpretatione”. Catholic University of America Press, Washington, DC, USA, 2014. ISBN 0-8132-2603-1 (hardcover). xi + 386 pp. LCCN B439 .D8613 2014.

URL http://muse.jhu.edu/books/9780813226040/;   http://
www.jstor.org/stable/10.2307/j.ctt9qdqj3.

Callergaard:2015:BRZ

[Cal15] Robert Callergård. Book review: Zvi Biener and Eric Schliesser, eds., Newton and Empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):194–197, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/680376.

Campos:2020:BRH

[Cam20] Luis Campos. Book review: Hans-Jörg Rheinberger and Staffan Müller–Wille, The Gene: From Genetics to Postgenomics. Trans. Adam Bostanci. Chicago: University of Chicago Press, 2017. Pp.

147. $25.00 (paper); $75.00 (cloth). ISBN 978-0-226-47478-6. HO-POS: Journal of the International Society for the History of Philosophy of Science, 10(1):344–347, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Canales:2016:PPE

[Can16] Jimena Canales. Physicist and the Philosopher: Einstein, Bergson, and the Debate That Changed Our Understanding of Time. Princeton University Press, Princeton, NJ, USA, 2016. ISBN 0-691-16534-3 (hardcover), 0-691-17317-6, 1-4008-9722-X. viii + 479 pp. LCCN BD638 .C326 2015. URL http://public.eblib.com/

choice/publicfullrecord.aspx?p=1936462.

Cao:2010:CAQ

[Cao10] Tian Yu Cao. From current algebra to quantum chromodynamics: a case for structural realism. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-88933-2, 0-511-78175-X (e-book). x + 308 pp. LCCN QC793.3.Q35 C36 2010.

Capps:2020:BRJ

[Cap20] John Capps. Book review: Jordi Cat and Adam Tamas Tuboly, eds., Neurath Reconsidered: New Sources and Perspectives. Boston Studies in the Philosophy and History of Science 336. Cham: Springer Nature, 2019. Pp. xii + 706. $149.00 (cloth). ISBN 978-3-030-02128-3. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):353–357, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Carlin:2016:BRM

[Car16a] Laurence Carlin. Book review: Martin Lenz and Anik Waldow, eds., Contemporary Perspectives on Early Modern Philosophy: Nature and Norms in Thought. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):178–180, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Carson:2016:BRT

[Car16b] Emily Carson. Book review: Thomas C. Vinci, Space, Geometry, and Kant’s Transcendental Deduction of the Categories. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):341–344, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Carnap:2018:LCT

[Car18] Rudolf Carnap. L’espace: une contribution à la théorie de la science. (French) [Space: a contribution to the theory of science]. Gallimard, Paris, France, 2018. ISBN 2-07-012808-3. 188 pp. LCCN ????

Carus:2024:BRR

[Car24] A. W. Carus. Book review: Rudolf Carnap, Tagebücher, Vol. 1, 1908–1919, edited by Christian Damböck; Rudolf Carnap, Tagebücher, Vol. 2, 1920–1935, edited by Christian Damböck. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(2):648–654, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Cassirer:1910:SFU

[Cas10] Ernst Cassirer. Substanzbegriff und Funktionsbegriff: Untersuchungen über die Grundfragen der Erkenntniskritik. B. Cassirer, Berlin, Germany, 1910. xv + 459 pp. LCCN BD221 .C2840 1910.

Cat:2012:BRT

[Cat12] Jordi Cat. Book review: Thomas Uebel, Empiricism at the Crossroads: The Vienna Circle’s Protocol-Sentence Debate. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):354–360, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/666513.

Cauvin:2020:PPC

[Cau20] Jean-Paul Cauvin. From phenomenology to the philosophy of the concept: Jean Cavaillès as a reader of Edmund Husserl. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):24–47, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Cauvin:2024:PRM

[Cau24] Jean-Paul Cauvin. Ptolemaic revolutions: Mathematical objectivity in Jean Cavaillès and Gilles-Gaston Granger. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):397–434, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Comte:2012:O

[CBCD12] Auguste Comte, Michel Bourdeau, Laurent Clauzade, and Frédéric Dupin. Oeuvres: Cours de philosophie positive. Leçons 46–51. (French) [Works: Course of Positive Philosophy. Lectures 46–51]. Collection Hermann Philosophie. Hermann, Paris, France, 2012. ISBN 2-7056-8163-9 (hardcover: v. 1). 427 pp. LCCN B2204 .B68 2012.

Chemla:2016:OHG

[CCR16] Karine Chemla, Renaud Chorlay, and David Rabouin, editors. The Oxford Handbook of Generality in Mathematics and the Sciences. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-877726-4. xii + 507 pp. LCCN QA8.4 .O93 2016.

Craver:2013:SMD

[CD13] Carl F. Craver and Lindley Darden. In search of mechanisms: discoveries across the life sciences. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-03965-X (hardcover), 0-226-03979-X (paperback), 0-226-03982-X (e-book). xxii + 228 pp. LCCN QH331 .C898 2013.

Chimisso:2008:WHM

[Chi08] Cristina Chimisso. Writing the History of the Mind: Philosophy and Science in France, 1900 to 1960s. Science, Technology and Culture. Ashgate., Farnham, Surrey, England, 2008. ISBN 0-7546-5705-1, 0-7546-9025-3 (e-book). ISSN 1700-1945. vi + 209 pp. LCCN BD418.3.

Chirimuuta:2020:CGA

[Chi20] M. Chirimuuta. Cassirer and Goldstein on abstraction and the autonomy of biology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):471–503, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Cimen:2019:SAP

[Çim19] Ünsal Çimen. On saving the astronomical phenomena: Physical realism in struggle with mathematical realism in Francis Bacon, al-Bitruji, and Averroës. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):135–151, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Chimisso:2021:HMP

[CJ21] Cristina Chimisso and Nicholas Jardine. Hélène Metzger on precursors: a historian and philosopher of science confronts her evil demon. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):331–353, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Castelao-Lawless:2012:BRA

[CL12] Teresa Castelão-Lawless. Book review: Anastasios Brenner, Raison scientifique et valeurs humaines. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):182–185, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/663732.

Cobb:2012:IPE

[Cob12] Aaron D. Cobb. Inductivism in practice: Experiment in John Herschel’s philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):21–54, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/663302.

Coko:2020:JPP

[Cok20] Klodian Coko. Jean Perrin and the philosophers’ stories: The role of multiple determination in determining Avogadro’s number. HO-POS: Journal of the International Society for the History of Philosophy of Science, 10(1):143–193, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Connell:2024:BRM

[Con24a] Sophia Connell. Book review: Marc Gasser-Wingate, Aristotle’s Empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):617–620, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Connolly:2024:BRS

[Con24b] Patrick J. Connolly. Book review: Stewart Duncan, Materialism from Hobbes to Locke. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):610–613, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Corneanu:2011:RMB

[Cor11] Sorana Corneanu. Regimens of the mind: Boyle, Locke, and the early modern cultura animi tradition. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-11639-5 (hardcover), 1-283-35515-9, 0-226-11641-7 (e-book). ix + 308 pp. LCCN B1131 .C67 2011.

Cowles:2012:BRS

[Cow12] Henry M. Cowles. Book review: Steven Gimbel, ed., Exploring the Scientific Method: Cases and Questions. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2 (1):154–157, Spring 2012. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/663303.

Cozzoli:2024:LLC

[Coz24] Daniele Cozzoli. The Logic of Life, the creation of the European Molecular Biology Laboratory, and the relation between molecular biology and physics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):463–482, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Cristalli:2021:AGL

[CP21] Claudia Cristalli and Ahti-Veikko Pietarinen. Abstraction and generalization in the logic of science: Cases from nineteenth-century scientific practice. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):93–121, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Creath:2020:BRG

[Cre20] Richard Creath. Book review: Gary Ebbs, Carnap, Quine, and Putnam on Methods of Inquiry. Cambridge: Cambridge University Press, 2017. Pp. xi + 278. $105.00 (cloth). ISBN 978-1-107-17815-1. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):337–340, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Cross:2016:BRJ

[Cro16] Richard Cross. Book review: John Duns Scotus, Duns Scotus on Time and Existence: The Questions on Aristotle’s “De interpretatione”. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):352–353, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Currie:2019:FHP

[CW19] Adrian Currie and Kirsten Walsh. Frameworks for historians and philosophers. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):1–34, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

DAgostino:2024:JNT

[D’A24] Pierluigi D’Agostino. Johann Nikolaus Tetens (1736–1807) and the idea of phoneme: a chapter in the history of linguistic thought. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(1):185–209, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dambock:2014:CMP

[Dam14] Christian Damböck. Caught in the middle: Philosophy of science between the historical turn and formal philosophy as illustrated by the program of “Kuhn Sneedified”. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4 (1):62–82, Spring 2014. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/673297.

Dambock:2017:BRR

[Dam17] Christian Damböck. Book review: Robert J. Richards and Lorraine Daston, eds., Kuhn’s Structure of Scientific Revolutions at Fifty: Reflections on a Science Classic. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):154– 156, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dambock:2020:WDP

[Dam20] Christian Damböck. What is descriptive psychology? Ebbinghaus’s 1896 criticism of Dilthey revisited. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):274– 289, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

deCeglia:2021:MEG

[dC21] Francesco Paolo de Ceglia. Matter is not enough: Georg Ernst Stahl, Friedrich Hoffmann, and the issue of animism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):502–527, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

deCallatay:2011:MPA

[dCH11] Godefroid de Callataÿ and Bruno Halflants, editors. On magic. Part I: an Arabic critical edition and English translation of Epistle 52a. Epistles of the Brethren of Purity. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-963895-0 (hardcover). xxv + 1 + 198 + 110 pp. LCCN B746.A4 E52 2011.

DeGroot:2014:AE

[De 14] Jean De Groot. Aristotle’s Empiricism. Parmenides Publishing, Las Vegas, NV, USA, 2014. ISBN 1-930972-83-0 (paperback), 1-322-20576-0 (MyiLibrary), 1-930972-84-9 (e-book). xxv + 442 pp.

LCCN Q124.95 .D43 2014. URL http://www.myilibrary.com?
id=651856.

DeRibera-Martin:2019:MKE

[De 19] Ignacio De Ribera-Martin. Movement (Kinêsis) as efficient cause in Aristotle’s Generation of Animals. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2): 296–326, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Bianchi:2022:KFC

[De 22] Silvia De Bianchi. Kant’s functional cosmology: Teleology, measurement, and symbolic representation in the Critique of Judgment. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):209–224, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

DeBenedetto:2024:WCE

[De 24] Matteo De Benedetto. What conceptual engineering can learn from the history of philosophy of science: Healthy externalism and metasemantic plasticity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):1–24, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Deltete:2012:POS

[Del12] Robert J. Deltete. Planck, Ostwald, and the Second Law of Thermodynamics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):121–146, Spring 2012. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/663835.

Dellsen:2017:CED

[Del17] Finnur Dellsén. Certainty and explanation in Descartes’s philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):302–327, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Deltete:2019:BRC

[Del19] Robert J. Deltete. Book review: Christa Jungnickel and Russell McCormmach, The Second Physicist: On the History of Theoretical Physics in Germany. Cham: Springer, 2017. Pp. xxxi + 460.

$180.00 (cloth); $140.00 (e-book). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):209– 211, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Demarest:2021:SAN

[Dem21] Boris Demarest. Soul, archeus, and nature in van Helmont’s medical naturalism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):564–584, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

DeSantis:2012:BRF

[DeS12] Anthony J. DeSantis. Book review: Frédérique Aït-Touati, Fictions of the Cosmos: Science and Literature in the Seventeenth Century. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):387–390, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/666968.

DeSantis:2014:BRC

[DeS14] Anthony J. DeSantis. Book review: Christoph Lüthy, David Gorlaeus (1591–1612): An Enigmatic Figure in the History of Philosophy and Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):196–199, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/675436.

Detlefsen:2015:BRE

[Det15] Karen Detlefsen. Book review: Eric Watkins, ed., The Divine Order, the Human Order, and the Order of Nature: Historical Perspectives. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):187–190, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/678185.

Detlefsen:2016:BRR

[Det16] Karen Detlefsen. Book review: Roger Ariew, Dennis Des Chene, Douglas M. Jesseph, Tad M. Schmaltz, and Theo Verbeek, Historical Dictionary of Descartes and Cartesian Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):345–348, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dethier:2018:WWS

[Det18] Corey Dethier. William Whewell’s semantic account of induction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):141–156, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

DeWitt:2004:WIH

[DeW04] Richard DeWitt. Worldviews: an introduction to the history and philosophy of science. Blackwell, Oxford, UK, 2004. ISBN 1-4051-1619-6, 1-4051-1620-X (paperback). xii + 326 pp. LCCN Q125 .D38 2008.

DeWitt:2017:BRJ

[DeW17a] Richard DeWitt. Book review: Jean De Groot, Aristotle’s Empiricism: Experience and Mechanics in the Fourth Century BC. HO-POS: Journal of the International Society for the History of Philosophy of Science, 7(1):176–179, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dewulf:2017:RCI

[Dew17b] Fons Dewulf. Rudolf Carnap’s incorporation of the Geisteswissenschaften in the Aufbau. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):199–225, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dambock:2020:DPV

[DFK20] Christian Damböck, Uljana Feest, and Martin Kusch. Descriptive psychology and Völkerpsychologie — in the contexts of historicism, relativism, and naturalism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):226–233, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dika:2022:DDL

[Dik22] Tarek R. Dika. Descartes’s deduction of the law of refraction and the shape of the anaclastic lens in Rule 8. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2):395–446, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dilley:2013:EMN

[Dil13] Stephen Dilley. The evolution of methodological naturalism in the origin of species. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):20–58, Spring 2013. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/667897.

Daston:2011:HSO

[DL11] Lorraine J. Daston and Elizabeth Lunbeck, editors. Histories of scientific observation. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-13677-9 (hardcover), 0-226-13678-7 (paperback). 460 + 1 pp. LCCN Q174.8 .H57 2011.

Dresow:2018:BRC

[DL18] Max Dresow and Alan C. Love. Book review: Catherine Kendig, ed., Natural Kinds and Classification in Scientific Practice. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):217–222, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dobre:2013:CE

[DNB13] Mihnea Dobre and Tammy Nyden-Bullock. Cartesian empiricisms. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-7689-6, 94-007-7690-X (e-book). xiii + 326 pp. LCCN B1875 .C3

2013. URL http://www.loc.gov/catdir/enhancements/fy1405/
2013956125-d.html;   http://www.loc.gov/catdir/
enhancements/fy1405/2013956125-t.html.

Dobre:2020:BRA

[Dob20a] Mihnea Dobre. Book review: Andrea Strazzoni, Dutch Cartesianism and the Birth of Philosophy of Science: From Regius to ’s Gravesande. Boston: de Gruyter, 2019. Pp. ix + 245. EUR 99.95 (cloth). ISBN 978-3-110-56782-3. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):609– 612, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dobre:2020:JRM

[Dob20b] Mihnea Dobre. Jacques Rohault’s mathematical physics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):414–439, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Domski:2011:BRP

[Dom11] Mary Domski. Book review: Peter Machamer and J. E. McGuire, Descartes’s Changing Mind. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):162–165, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/657628.

Domski:2013:ERP

[Dom13] Mary Domski. Essay review: Putting the pieces back together again: Reading Newton’s Principia through Newton’s method: Steffen Ducheyne, “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology. William L. Harper. Isaac Newton’s Scientific Method: Turning Data into Evidence about Gravity and Cosmology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):318–333, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/670269.

Dong:2024:BRR

[Don24] Hao Dong. Book review: Richard T. W. Arthur, ed., Leibniz: Journal Articles on Natural Philosophy, trans. Jeffrey K. McDonough, Lea Aurelia Schroeder, Samuel Levey, Richard T. W. Arthur, Richard Francks (with Roger Woolhouse), and Tzuchien Tho. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(2):631–634, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dorsch:2020:BRE

[Dor20] Kate Dorsch. Book review: Emily Herring, Kevin Matthew Jones, Konstantin S. Kiprijanov, and Laura M. Sellers, eds., The Past, Present, and Future of Integrated History and Philosophy of Science. London: Routledge, 2019. Pp. xii + 255. $155.00 (cloth). ISBN 978-0-815-37985-0. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):598–601, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dougherty:2024:WSC

[Dou24] John Dougherty. Wilfrid Sellars and constructive empiricism. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(2):435–478, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Downing:2017:BRG

[Dow17] Lisa Downing. Book review: Gideon Manning, ed., Matter and Form in Early Modern Science and Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):381–383, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Doyle:2024:BRW

[Doy24] Emerson P. Doyle. Book review: A. W. Carus, Michael Friedman, Wolfgang Kienzler, Alan Richardson, and Sven Schlotter, eds., The Collected Works of Rudolf Carnap. Vol. 1, Rudolf Carnap: Early Writings. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):210–215, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

dePaz:2014:PPS

[dP14] Maria de Paz. Poincaré, philosopher of science: problems and perspectives. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 94-017-8779-4, 94-017-8780-8 (e-book). xiv + 191 pp. LCCN B2430.P564 P64

2014. URL http://www.loc.gov/catdir/enhancements/fy1407/
2014936171-d.html;   http://www.loc.gov/catdir/
enhancements/fy1407/2014936171-t.html.

dePaz:2016:BRA

[dP16] María de Paz. Book review: Anastasios Brenner, ed., Les textes fondateurs de l’épistémologie française. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2): 325–329, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Demeter:2019:BRD

[DP19] Tamás Demeter and Krisztián Pete. Book review: David Landy, Hume’s Science of Human Nature: Scientific Realism, Reason, and Substantial Explanation. London: Routledge, 2018. Pp. xi + 266. £120.00 (cloth). ISBN 978-1-138-50313-7. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2): 415–419, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

dePaz:2021:PRN

[dP21] María de Paz. Poincaré, Le Roy, and the Nouveau positivisme. HO-POS: Journal of the International Society for the History of Philosophy of Science, 11(2):446–460, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

diPoppa:2024:TMS

[dP24] Francesca di Poppa. Tad m. schmaltz, The Metaphysics of the Material World: Suárez, Descartes, Spinoza. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14 (1):222–225, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Drozdova:2011:BRC

[Dro11] Daria Drozdova. Book review: Cristina Chimisso, Writing the History of the Mind: Philosophy and Science in France, 1900 to 1960s. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):348–351, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/661747.

Demarest:2021:SII

[DRW21] Boris Demarest, Jonathan Regier, and Charles Wolfe. Special issue introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):494–501, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dewulf:2021:PAC

[DS21] Fons Dewulf and Massimiliano Simons. Positivism in action: The case of Louis Rougier. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):461–487, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ducheyne:2011:MBN

[Duc11] Steffen Ducheyne. The main business of natural philosophy: Isaac Newton’s natural-philosophical methodology, volume 5644 of Archimedes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-2125-0 (hardcover), 94-007-2126-9 (e-book). xxv + 352 pp. LCCN QC6 .D83 2012.

Duchesneau:2012:BRJ

[Duc12] François Duchesneau. Book review: Justin E. H. Smith, Divine Machines: Leibniz and the Science of Life. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):186–189, Spring 2012. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/664492.

Ducheyne:2014:BRE

[Duc14] Steffen Ducheyne. Book review: Eric Jorink and Ad Maas, eds., Newton and the Netherlands: How Isaac Newton Was Fashioned in the Dutch Republic. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):189–192, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/674979.

Ducheyne:2017:CPT

[Duc17] Steffen Ducheyne. Curing Pansophia through Eruditum Nescire: Bernard Nieuwentijt’s (1654–1718) epistemology of modesty. HO-POS: Journal of the International Society for the History of Philosophy of Science, 7(2):272–301, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Duerr:2023:PCR

[Due23] Patrick M. Duerr. Popper: Critical rationalist, conventionalist, and virtue epistemologist. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):54–90, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Duffy:2011:BRP

[Duf11] Simon Duffy. Book review: Peter Gaffney, ed., The Force of the Virtual: Deleuze, Science, and Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):340–343, Fall 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/660122.

Dunlop:2011:BRN

[Dun11] Katherine Dunlop. Book review: Niccolò Guicciardini, Isaac Newton on Mathematical Certainty and Method. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):359–364, Fall 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/659400.

Dunlop:2014:BRB

[Dun14] Katherine Dunlop. Book review: Peter Achinstein, Evidence and Method. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):361–365, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/677538.

Dunlop:2016:PFA

[Dun16] Katherine Dunlop. Poincaré on the foundations of arithmetic and geometry. Part 1: Against “dependence-hierarchy” interpretations. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):274–308, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dunlop:2017:PFA

[Dun17] Katherine Dunlop. Poincaré on the foundations of arithmetic and geometry. Part 2: Intuition and unity in mathematics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):88–107, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Dunham:2024:BRC

[Dun24] Jeremy Dunham. Book review: Charles Bonnet, Analytical Essay on the Faculties of the Soul. Translated and introduced by Stephen Gaukroger. Oxford: Oxford University Press, 2022. Pp. xxxiv + 234. $85.00 (cloth). ISBN 978-0-19-284677-8. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):554–557, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Earnshaw:2013:BRM

[Ear13] Eugene Earnshaw. Book review: Mark Borrello, Evolutionary Restraints: The Contentious History of Group Selection. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):356–360, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/670132.

Eaton:2021:BRM

[Eat21] William Eaton. Book review: Marina Paola Banchetti-Robino, The Chemical Philosophy of Robert Boyle: Mechanism, Chymical Atoms, and Emergence. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):415–419, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ebrey:2015:TPA

[Ebr15] David Ebrey, editor. Theory and practice in Aristotle’s natural science. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-05513-X. viii + 261 pp. LCCN B491.N43 T44 2015.

Eckes:2018:BRK

[Eck18] Christophe Eckes. Book review: Karine Chemla, Renaud Chorlay, and David Rabouin, eds., The Oxford Handbook of Generality in Mathematics and the Sciences. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):214– 217, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Eddy:2015:BRJ

[Edd15] Matthew Daniel Eddy. Book review: John C. Powers, Inventing Chemistry: Herman Boerhaave and the Reform of the Chemical Arts. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):385–387, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/682380.

Eder:2024:BMD

[Ede24] Günther Eder. The Bernays–Müller debate. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):317–361, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Edgar:2020:VOH

[Edg20a] Scott Edgar. Völkerpsychologie and the origins of Hermann Cohen’s antipsychologism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):254–273, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Edgar:2020:HCP

[Edg20b] Scott Edgar. Hermann Cohen’s Principle of the Infinitesimal Method: A defense. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):440–470, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Erickson:2013:HRA

[EKD^{+}13] Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca M. (Rebecca Maura) Lemov, Thomas Sturm, and Michael D. Gordin. How reason almost lost its mind: the strange career of Cold War rationality. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-04663-X (hardcover), 0-226-04677-X (e-book). viii + 259 pp. LCCN D843 .E69 2013.

Ekholm:2013:BRH

[Ekh13] Karin Ekholm. Book review: Hiro Hirai, Medical Humanism and Natural Philosophy: Renaissance Debates on Matter, Life and the Soul. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):367–371, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/670133.

Elazar:2013:BRA

[Ela13] Michael Elazar. Book review: Annibale Fantoli, The Case of Galileo: A Closed Question? HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):380–384, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/671115.

Elliott:2017:EIR

[ER17] Kevin Christopher Elliott and Ted Richards. Exploring Inductive Risk: Case Studies of Values in Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-046771-1 (hardcover), 0-19-046772-X (paperback), 0-19-046773-8 (e-book), 0-19-046775-4. xvi + 290 pp. LCCN Q175.32.I54.

Erickson:2015:WGT

[Eri15] Paul Erickson. The World the Game Theorists Made. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-09703-X (hardcover), 0-226-09717-X (paperback), 0-226-09720-X (e-book). 390 pp. LCCN HB144 .E75 2015.

Elgin:2017:PSA

[ES17] Mehmet Elgin and Elliott Sober. Popper’s shifting appraisal of evolutionary theory. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):31–55, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Esanu:2019:ACJ

[Esa19] Andreea Esanu. Auguste Comte and J. S. Mill on physical causes: The case of Joseph Fourier’s analytical theory of heat. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):275–295, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Frost-Arnold:2011:QEC

[FA11] Greg Frost-Arnold. Quine’s evolution from ‘Carnap’s Disciple’ to the author of “Two Dogmas”. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):291–316, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/660011.

Fabry:2024:CPC

[Fab24] Lucie Fabry. Cross-perspectives on the construction of scientific facts: Latour and Woolgar as readers of Bachelard. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):52–77, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Falkenburg:2016:BRJ

[Fal16] Brigitte Falkenburg. Book review: John Losee, Complementarity, Causality, and Explanation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):162– 164, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Falcon:2017:BRC

[Fal17] Andrea Falcon. Book review: Chelsea C. Harry, Chronos in Aristotle’s Physics: On the Nature of Time. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2): 395–397, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Falcon:2018:BRA

[Fal18] Andrea Falcon. Book review: André Laks, The Concept of Presocratic Philosophy: Its Origin, Development, and Significance. HO-POS: Journal of the International Society for the History of Philosophy of Science, 8(2):486–489, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Falkenburg:2021:ERG

[Fal21] Brigitte Falkenburg. Essay review: Grete Hermann’s philosophy of quantum mechanics: a late appraisal. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11 (1):201–210, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Fantoli:2012:CGC

[Fan12] Annibale Fantoli. The case of Galileo: a closed question? University of Notre Dame, Notre Dame, IN, USA, 2012. ISBN 0-268-02891-5 (paperback), 0-268-07972-2 (e-book). xii + 271 pp. LCCN QB36.G2

F2613 2010. URL http://muse.jhu.edu/books/9780268079727/.

Feest:2010:HPE

[Fee10] Uljana Feest, editor. Historical perspectives on Erklären and Verstehen, volume 21 of Archimedes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3539-7, 90-481-3540-0 (e-book). vi + 320 pp. LCCN BD237 .H57

2010. URL http://www.loc.gov/catdir/enhancements/fy1008/
2009942131-d.html;   http://www.loc.gov/catdir/
enhancements/fy1008/2009942131-t.html.

Feest:2018:BRC

[Fee18] Uljana Feest. Book review: Christian Damböck, <Deutscher Empirismus>: Studien zur Philosophie im deutschsprachigen Raum 1830–1930. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):480–485, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Fehige:2021:AMT

[Feh21] Yiftach Fehige. The Annus Mirabilis of 1986: Thought experiments and scientific pluralism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):222–240, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Feola:2011:BRM

[Feo11] Vittoria Feola. Book review: Maria Toscano, Gli archivi del mondo: Antiquaria, storia naturale e collezionismo nel secondo Settecento. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):178–181, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/658490.

Ferrini:2022:SPH

[Fer22] Cinzia Ferrini. A “physiogony” of the heavens: Kant’s early view of universal natural history. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):261–285, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

FerreiradaCunha:2024:ONM

[Fer24] Ivan Ferreira da Cunha. Otto Neurath’s modernist utopianism: Linking the Vienna Circle and H. G. Wells. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):25–51, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ferrari:2023:BRD

[FES^{+}23] Massimo Ferrari, Mehmet Elgin, Elliott Sober, Thomas Uebel, and Donata Romizi. Book review: Donata Romizi, Dem wissenschaftlichen Determinismus auf der Spur: Von der klassischen Mechanik zur Entstehung der Quantenphysik. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):202–208, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Fisch:2017:CUT

[Fis17] Menachem Fisch. Creatively undecided toward a history and philosophy of scientific agency. University of Chicago Press, Chicago, IL, USA, 2017. ISBN 0-226-51448-X (hardcover), 0-226-51451-X (paperback), 0-226-51465-X (e-book). vii + 295 pp. LCCN Q175 .F5452 2017.

Frappier:2012:TES

[FMB12] Melanie Frappier, Letitia Meynell, and James Robert Brown. Thought Experiments in Science, Philosophy, and the Arts. Routledge Studies in the Philosophy of Science. Taylor and Francis, Boca Raton, FL, USA, 2012. ISBN 1-136-28600-4, 0-415-88544-2. 283 (est.) pp. LCCN BD265 .T465 2012. URL http://public.eblib.

com/choice/publicfullrecord.aspx?p=1024628;
http://www.123library.org/book\_details/?id=68629.

Friedman:2006:KLN

[FN06] Michael Friedman and Alfred Nordmann, editors. The Kantian legacy in Nineteenth-Century science. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-06254-2. vi + 370 pp.

LCCN Q174.8 .K36 2006. URL http://www.loc.gov/catdir/toc/
fy0703/2005058035.html.

Fogel:2012:BRJ

[Fog12] Brandon Fogel. Book review: Jeroen van Dongen, Einstein’s Unification. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):372–376, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/666957.

Folina:2019:SHH

[Fol19] Janet Folina. Science, hypothesis, and hierarchy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):388–406, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Follesa:2022:CAP

[Fol22] Laura Follesa. Cosmology, astronomy, and philosophy around 1800: Schelling, Hegel, Herder. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):242–260, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Fraenkel:2011:RBT

[FPS11] Carlos Fraenkel, Dario Andres Perinetti, and Justin E. H. Smith, editors. The Rationalists: between tradition and innovation, volume 65 of The new synthese historical library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 90-481-9384-2. vii + 224 pp. LCCN BL2785 .R38

2011. URL http://www.loc.gov/catdir/enhancements/fy1114/
2010936524-b.html;   http://www.loc.gov/catdir/
enhancements/fy1114/2010936524-d.html;   http://www.
loc.gov/catdir/enhancements/fy1114/2010936524-t.html.

Franco:2012:KCM

[Fra12] Paul L. Franco. Are Kant’s concepts and methodology inconsistent with scientific change? constitutivity and the synthetic method in Kant. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):321–353, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/664819.

Frappier:2013:BRI

[Fra13] Mélanie Frappier. Book review: Isabelle Stengers, Cosmopolitics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):341–344, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/671744.

Franco:2018:OLC

[Fra18] Paul L. Franco. Ordinary language criticisms of logical positivism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):157–190, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Fraser:2022:VSN

[Fra22] Craig Fraser. Vesto Slipher, nebular spectroscopy, and the birth of modern cosmology, 1912–22. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):146– 169, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Frechette:2020:DPB

[Fré20] Guillaume Fréchette. Descriptive psychology: Brentano and Dilthey. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):290–307, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Frierson:2015:MMP

[Fri15] Patrick R. Frierson. Maria Montessori’s philosophy of experimental psychology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):240–268, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/682395.

Fulkerson-Smith:2012:BRA

[FS12] Brett A. Fulkerson-Smith. Book review: Angus Nicholls and Martin Liebscher, eds., Thinking the Unconscious: Nineteenth-Century German Thought. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):166–169, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/664186.

Futch:2014:BRB

[Fut14] Michael Futch. Book review: Philip Ball, Curiosity: How Science Became Interested in Everything. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):186–189, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675431.

Gabbe:2024:BRM

[Gab24] Myrna Gabbe. Book review: Marguerite Deslauriers, Aristotle on Sexual Difference: Metaphysics, Biology, Politics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):267–270, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gaffney:2010:FVD

[Gaf10] Peter Gaffney. Force of the Virtual: Deleuze, Science, and Philosophy. University of Minnesota Press, Minneapolis, MN, USA, 2010. ISBN 0-8166-6597-4 (hardcover), 0-8166-6598-2 (paperback), 0-8166-7356-X (e-book). ix + 393 pp. LCCN Q175 .F7126 2010.

URL http://muse.jhu.edu/books/9780816673568/;   http://
public.eblib.com/choice/publicfullrecord.aspx?p=661346.

Galilei:1957:DOG

[Gal57] Galileo Galilei. Discoveries and opinions of Galileo. Doubleday anchor books, A94. Doubleday, Garden City, NY, USA, 1957. viii + 302 pp. LCCN QB36.G2 G313.

Gambarotto:2015:BRH

[Gam15] Andrea Gambarotto. Book review: Hein van den Berg, Kant on Proper Science: Biology in the Critical Philosophy and the Opus postumum. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):364–367, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/682377.

Gandon:2021:ISR

[Gan21] Sébastien Gandon. Infinity and the self: Royce on Dedekind. HO-POS: Journal of the International Society for the History of Philosophy of Science, 11(2):354–382, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Garber:2009:LBS

[Gar09] Daniel Garber. Leibniz: body, substance, monad. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956664-X (hardcover), 0-19-157062-1 (e-book). xxi + 428 pp. LCCN B2599.M7 .G37 2009.

Garber:2015:BRM

[Gar15a] Daniel Garber. Book review: Mihnea Dobre and Tammy Nyden, eds., Cartesian Empiricisms. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):374–377, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/682722.

Garson:2015:BRC

[Gar15b] Justin Garson. Book review: Carl F. Craver and Lindley Darden, In Search of Mechanisms: Discoveries across the Life Sciences. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):180–183, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/678187.

Garnick:2024:BRT

[Gar24] Alex Garnick. Book review: Tarek R. Dika, Descartes’s Method: The Formation of the Subject of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14 (2):603–606, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gaukroger:2010:CMR

[Gau10] Stephen Gaukroger. The collapse of mechanism and the rise of sensibility: science and the shaping of modernity, 1680–1760. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-959493-7. ix + 505 pp. LCCN Q127.E8.

Gava:2024:CAA

[Gav24] Gabriele Gava. Conceptual analysis and the analytic method in Kant’s Prize Essay. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):164–184, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gawne:2011:BRR

[Gaw11] Richard Gawne. Book review: Richard DeWitt, ed., Worldviews: An Introduction to the History and Philosophy of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):331–333, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/660808.

Granada:2017:UHE

[GBT17] Miguel Ángel Granada, Patrick Boner, and Dario Tessicini. Unifying Heaven and Earth: essays in the history of early modern cosmology. Biblioteca universitària. Edicions de la Universitat de Barcelona, Barcelona, Spain, 2017. ISBN 84-475-3960-1. LCCN Q124.6.

Gorham:2022:BRA

[GBV^{+}22] Geoffrey Gorham, Silvia De Bianchi, Federico Viglione, Helge Kragh, Siska De Baerdemaeker, and Mike D. Schneider. Book review: Andrew Janiak, ed., Space: a History. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):322–325, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gal:2012:BS

[GCM12] Ofer Gal and Raz D. Chen-Morris. Baroque science. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-92398-3 (hardcover), 0-226-92399-1 (paperback), 0-226-92399-1 (e-book). xiv + 333 pp. LCCN Q127.E85 G35 2013.

Galilei:1990:DOG

[GD90] Galileo Galilei and Stillman Drake. Discoveries and opinions of Galileo: including The Starry Messenger (1610), Letter to the Grand Duchess Christina (1615), and excerpts from Letters on Sunspots (1613), The assayer (1623). Anchor Books, New York, NY, USA, 1990. ISBN 0-385-09239-3. viii + 301 pp. LCCN QB41

.G123 1990. URL http://www.loc.gov/catdir/description/
random043/90030691.html.

Gee:2013:AAT

[Gee13] Emma Gee. Aratus and the astronomical tradition. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-978168-0 (hardcover), 0-19-978178-8 (e-book). xi + 298 pp. LCCN PA3873.A7 G44 2013.

Georgescu:2019:BRJ

[Geo19] Laura Georgescu. Book review: Jutta Schickore, About Method: Experimenters, Snake Venom, and the History of Writing Scientifically. Chicago: University of Chicago Press, 2017. Pp. 316. $50.00 (cloth). ISBN 978-0-226-44998-2. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):410– 415, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gerson:2010:CHP

[Ger10] Lloyd P. Gerson. The Cambridge history of philosophy in late antiquity, volume 1 of Cambridge histories online. Cambridge University Press, Cambridge, UK, 1. publ. edition, 2010. ISBN 0-521-76440-8 (vol. 1), 0-521-19484-9 (vol. 2), 0-521-87642-7 (set), 1-139-09545-5. xiii + 581 pp. LCCN B111 .C37 2010.

Gaukroger:2008:ESC

[GG08] Stephen Gaukroger and Stephen Gaukroger. The emergence of a scientific culture: science and the shaping of modernity 1210–1685. Clarendon Press, New York, NY, USA, 2008. ISBN 0-19-955001-8 (paperback), 0-19-929644-8. ix + 563 pp. LCCN BL240.3 .I58 v. 199.

Gibson:2015:AVM

[Gib15] Susannah Gibson. Animal, vegetable, mineral?: how eighteenthcentury science disrupted the natural order. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-870513-1 (hardcover). xv + 215 pp. LCCN QH83 .G53 2015.

Giere:2012:BRD

[Gie12] Ronald N. Giere. Book review: Dirk R. Johnson, Nietzsche’s Anti-Darwinism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):380–382, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/666830.

Giesen:2024:BRG

[Gie24] Alexander Theo Giesen. Book review: Gemma Anderson and John Dupré, eds., Drawing Processes of Life: Molecules, Cells, Organisms. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):637–640, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Giglioni:2016:BRO

[Gig16] Guido Giglioni. Book review: Ohad Nachtomy and Justin E. H. Smith, eds., The Life Sciences in Early Modern Philosophy. HO-POS: Journal of the International Society for the History of Philosophy of Science, 6(2):348–352, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gill:2014:LTA

[Gil14] Mary Louise Gill. The limits of teleology in Aristotle’s meteorology IV.12. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):335–350, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/677567.

Gimbel:2011:ESM

[Gim11] Steven Gimbel, editor. Exploring the scientific method: cases and questions. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-29481-1 (hardcover), 0-226-29483-8 (paperback). xvii + 406 pp. LCCN Q175.3 .E97 2011.

Gimbel:2012:EJS

[Gim12] Steven Gimbel. Einstein’s Jewish science: physics at the intersection of politics and religion. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0554-7 (hardcover), 1-4214-0575-X (e-book). viii + 245 pp. LCCN QC173.55 .G55 2012.

Gimes:2021:MRW

[Gim21] Balázs Gimes. The materialism of Roy Wood Sellars. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):166–182, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Giovanelli:2017:TCE

[Gio17] Marco Giovanelli. Traditions in collision: The emergence of logical empiricism between the Riemannian and Helmholtzian traditions. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):328–380, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Giovanelli:2024:RTT

[Gio24] Marco Giovanelli. Relativity theory as a theory of principles: A reading of Cassirer’s Zur Einstein’schen Relativitätstheorie. HO-POS: Journal of the International Society for the History of Philosophy of Science, 13(2):261–296, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Glasner:2009:APT

[Gla09] Ruth Glasner. Averroes’ Physics: A Turning Point in Medieval Natural Philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956773-5 (hardcover). ix + 229 pp. LCCN QC7 .G65 2009.

Gotthelf:2012:TFP

[GLB12] Allan Gotthelf, Mariska Leunissen, and Pieter Beullens. Teleology, first principles, and scientific method in Aristotle’s biology. Oxford Aristotle studies. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-928795-3. xvi + 440 pp. LCCN QL41; QH331 .G668

2012. URL http://catdir.loc.gov/catdir/enhancements/
fy1215/2011278272-b.html;   http://catdir.loc.gov/
catdir/enhancements/fy1215/2011278272-d.html;   http://
catdir.loc.gov/catdir/enhancements/fy1215/2011278272-t.
html;   http://www.oxfordscholarship.com/view/10.1093/
acprof:oso/9780199287956.001.0001/acprof-9780199287956.

Goddu:2010:CAT

[God10] André Goddu. Copernicus and the Aristotelian tradition: education, reading, and philosophy in Copernicus’s path to heliocentrism, volume 15(12) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-18107-5 (hardcover). ISSN 1872-0684. xxvii + 545 pp. LCCN QB36.C8 G635 2010.

Goldberg:2011:BRL

[Gol11] Benjamin Goldberg. Book review: Lisa T. Sarasohn, The Natural Philosophy of Margaret Cavendish: Reason and Fancy during the Scientific Revolution. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):169–172, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/657932.

Goldberg:2016:BRM

[Gol16] Benjamin Goldberg. Book review: Marco Solinas, From Aristotle’s Teleology to Darwin’s Genealogy: The Stamp of Inutility. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):174–177, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Goldin:2017:BRD

[Gol17] Owen Goldin. Book review: David Bronstein, Aristotle on Knowledge and Learning: The Posterior Analytics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):173–176, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gonzalez:2014:ELR

[Gon14] Wenceslao J. Gonzalez. The evolution of Lakatos’s repercussion on the methodology of economics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):1–25, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675401.

Gorham:2011:BRW

[Gor11] Geoffrey Gorham. Book review: Walter Ott, Causation & Laws in Early Modern Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):371–375, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/659759.

Gori:2021:EMC

[Gor21] Pietro Gori. Ernst Mach’s contribution to the philosophy of science in light of Mary B. Hesse’s postempiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):383–411, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Garver:2014:RRM

[GP14] Valerie L. (Valerie Louise) Garver and Owen Michael Phelan, editors. Rome and religion in the medieval world: studies in honor of Thomas F. X. Noble. Church, faith and culture in the medieval West. Ashgate, Farnham, Surrey, England, 2014. ISBN 1-4724-2112-4 (hardcover), 1-4724-2113-2 (e-book), 1-4724-2114-0 (epub). xxviii + 349 pp. LCCN DG811 .R625 2014.

Garber:2013:MNP

[GR13] Daniel Garber and Sophie Roux, editors. The Mechanization of Natural Philosophy, volume 300 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-4344-0 (hardcover), 94-007-4345-9 (e-book). ISSN 0068-0346. LCCN QC6 .M3616 2013; QC6

.M43 2013. URL https://link.springer.com/book/10.1007/
978-94-007-4345-8.

Graham:2013:SBS

[Gra13] Daniel Graham. Science before Socrates. Parmenides, Anaxagoras, and the new astronomy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-995978-1 (hardcover), 0-19-995979-X. xiii + 287 pp. LCCN QB21 .G73 2013.

Grene:1991:DAS

[Gre91] Marjorie Grene. Descartes among the scholastics, volume 1991 of Aquinas lecture. Marquette University Press, Milwaukee, WI, USA, 1991. ISBN 0-87462-158-5. 50 pp. LCCN B1873 .G74 1991.

Greenberg:2013:BRN

[Gre13] Sean Greenberg. Book review: Noa Naaman-Zauderer, Descartes’ Deontological Turn: Reason, Will, and Virtue in the Later Writings. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):182–186, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/668083.

Greenberg:2015:BRC

[Gre15] Sean Greenberg. Book review: Carlos Fraenkel, Dario Perinetti, and Justin E. H. Smith, eds., The Rationalists: Between Tradition and Innovation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):378–381, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/682713.

Gregory:2020:BRG

[Gre20] Andrew Gregory. Book review: Gijsbert Jonkers, The Textual Tradition of Plato’s “Timaeus” and “Critias”. Leiden: Brill, 2017. Pp. xviii + 548. $202.00 (cloth). ISBN 978-90-04-32591-3. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):361–364, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Greene:2021:HCT

[Gre21] Catherine Greene. Historical counterfactuals, transition periods, and the constraints on imagination. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1): 305–323, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Griesemer:2013:BRD

[Gri13] James Griesemer. Book review: David Sepkoski, Rereading the Fossil Record: The Growth of Paleobiology as an Evolutionary Discipline. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):360–364, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/670326.

Gross:2015:BON

[Gro15] Angélique Groß. Die Bildpädagogik Otto Neuraths: Methodische Prinzipien der Darstellung von Wissen. (German) [Otto Neurath’s Image Pedagogy: Methodical Principles of the Representation of Knowledge], volume 21 of Veröffentlichungen des Instituts Wiener Kreis. Springer International Publishing AG, Cham, Switzerland, 2015. ISBN 3-319-16315-9 (hardcover), 3-319-16316-7 (e-book). ISSN 2363-5118. xii + 288 pp. LCCN PM8999 . G76 2015.

Grosholz:2020:BRR

[Gro20] Emily Grosholz. Book review: Roi Wagner, Making and Breaking Mathematical Sense: Histories and Philosophies of Mathematical Practice. Princeton, NJ: Princeton University Press, 2017. Pp. 256. $45.00 (cloth). ISBN 978-0-691-17171-5. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10 (1):332–337, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Galilei:2010:S

[GRVS10] Galileo Galilei, Eileen Adair Reeves, Albert Van Helden, and Christoph Scheiner. On sunspots. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-70715-6 (hardcover), 0-226-70716-4 (paperback), 1-282-90203-2 (e-book). xi + 418 pp. LCCN QB525 .G22 2010.

Gottlieb:2017:AND

[GS17] Paula Gottlieb and Elliott Sober. Aristotle on “nature does nothing in vain”. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):246–271, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Gubelmann:2019:SSP

[Gub19] Reto Gubelmann. From shared stimuli to preestablished harmony: The development of Quine’s thinking on intersubjectivity and objective validity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):343–370, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Guicciardini:2009:INM

[Gui09] Niccolò Guicciardini. Isaac Newton on mathematical certainty and method. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-01317-7. xxiii + 422 pp. LCCN QA29.N4 G85 2009.

Guillin:2014:BRM

[Gui14] Vincent Guillin. Book review: Michel Bourdeau, Auguste Comte: Science et société. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):371–375, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/677398.

Guicciardini:2018:BRE

[Gui18] Niccolò Guicciardini. Book review: Elizabethanne Boran and Mordechai Feingold, eds., Reading Newton in Early Modern Europe. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):205–209, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Goy:2014:KTB

[GW14] Ina Goy and Eric Watkins, editors. Kant’s Theory of Biology. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2014. ISBN 3-11-022578-6 (hardcover), 3-11-022579-4 (e). x + 321 pp. LCCN QH331 .K273 2014.

Horrocks:2012:VSS

[HA12] Jeremiah Horrocks and Wilbur Applebaum. Venus Seen on the Sun: the First Observation of a Transit of Venus by Jeremiah Horrocks, volume 29(18) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2012. ISBN 1-280-12669-8, 90-04-22193-X, 90-04-22194-8 (e-book). xxiv + 82 pp. LCCN QB509.5.

Haaparanta:2013:BRJ

[Haa13] Leila Haaparanta. Book review: Juha Manninen and Friedrich Stadler, eds., The Vienna Circle in the Nordic Countries: Networks and Transformations of Logical Empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):337–341, Fall 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/670129.

Hamid:2024:BRJ

[Ham24] Nabeel Hamid. Book review: Jeffrey K. McDonough, A Miracle Creed: The Principle of Optimality in Leibniz’s Physics and Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):624–627, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hart:2010:EL

[Har10a] W. D. (Wilbur Dyre) Hart. The evolution of logic. The evolution of modern philosophy. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76681-8 (hardcover), 0-521-74772-4 (paperback). xi + 299 pp. LCCN BC15 .H37 2010. URL http://assets.

cambridge.org/97805217/47721/cover/9780521747721.jpg.

Hart:2006:EL

[Har10b] Wilbur Dyre Hart. The Evolution of Logic. The evolution of modern philosophy. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76681-8 (hardcover), 0-521-74772-4 (paperback), 0-511-78620-4 (e-book). xi + 299 pp. LCCN BC15 .H37 2010.

Harper:2011:INS

[Har11] William L. Harper. Isaac Newton’s scientific method: turning data into evidence about gravity and cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-957040-X. xvii + 424 pp. LCCN QC16.N7 H37 2011.

Hardcastle:2014:BRJ

[Har14] Gary L. Hardcastle. Book review: Joel Isaac, Working Knowledge: Making the Human Sciences from Parsons to Kuhn. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):154–157, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/675454.

Harrop:2022:EEU

[Har22] Stephen Harrop. Essence, experiment, and underdetermination in the Spinoza–Boyle correspondence. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2): 447–484, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hattab:2009:DFM

[Hat09] Helen Hattab. Descartes on forms and mechanisms. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-51892-X (hardcover). x + 236 pp. LCCN B1878.M43 H38 2009.

Hatfield:2012:BRM

[Hat12a] Gary Hatfield. Book review: Michael Friedman and Alfred Nordmann, eds., The Kantian Legacy in Nineteenth-Century Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):172–177, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/662734.

Hattab:2012:DFM

[Hat12b] Helen Hattab. Descartes on forms and mechanisms. Cambridge University Press, Cambridge, UK, 2012. ISBN 0-521-51892-X (hardcover), 1-107-40515-7 (paperback). x + 236 pp. LCCN B1878.M43 H38 2009.

Hatfield:2015:REC

[Hat15a] Gary Hatfield. Radical empiricism, critical realism, and American functionalism: James and Sellars. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):129–153, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/680373.

Hattab:2015:BRC

[Hat15b] Helen Hattab. Book review: Craig Martin, Subverting Aristotle: Religion, History and Philosophy in Early Modern Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):381–385, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/682379.

Hatfield:2020:WHC

[Hat20a] Gary Hatfield. Wundt and “higher cognition”: Elements, association, apperception, and experiment. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1): 48–75, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hattab:2020:BRS

[Hat20b] Helen Hattab. Book review: Stefano di Bella and Tad M. Schmaltz, eds., The Problem of Universals in Early Modern Philosophy. Oxford: Oxford University Press, 2017. Pp. 352. $105.00 (cloth). ISBN 978-0-19-060804-0. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):329–332, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Heis:2014:ECB

[Hei14] Jeremy Heis. Ernst Cassirer’s Substanzbegriff und Funktionsbegriff. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):241–270, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/676959.

Hennig:2013:BRS

[Hen13] Boris Hennig. Book review: Stephan Schmid, Finalursachen in der frühen Neuzeit: Eine Untersuchung der Transformation teleologischer Erklärungen. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):179–182, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/669214.

Henne:2023:JDW

[Hen23] Céline Henne. John Dewey: Was the inventor of instrumentalism himself an instrumentalist? HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):120–150, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Henry:2024:BRE

[Hen24] John Henry. Book review: Eric Schliesser, Newton’s Metaphysics: Essays. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):246–250, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hepburn:2011:BRA

[Hep11] Brian Hepburn. Book review: Alisa Bokulich, Reexamining the Quantum-Classical Relation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):142–146, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/657933.

Herfeld:2019:IRT

[Her19] Catherine Herfeld. Imagination rather than observation in econometrics: Ragnar Frisch’s hypothetical experiments as thought experiments. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):35–74, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Huebner:2024:BRJ

[HES^{+}24] Daniel R. Huebner, Mehmet Elgin, Elliott Sober, Thomas Uebel, and Daniel Schwartz. Book review: José Itzigsohn and Karida L. Brown, The Sociology of W. E. B. Du Bois: Racialized Modernity and the Global Color Line. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):270–274, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hicks:2012:BRP

[Hic12] Marie Hicks. Book review: Peter Harrison, Ronald L. Numbers, and Michael H. Shank, eds., Wrestling with Nature: From Omens to Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):361–364, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/665253.

Hill:2012:BRP

[Hil12] Benjamin Hill. Book review: Peter R. Anstey, John Locke and Natural Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):382–387, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/666523.

Hirai:2011:MHN

[Hir11] Hiro Hirai. Medical humanism and natural philosophy: Renaissance debates on matter, life, and the soul, volume 17 of Medieval and early modern science; History of science and medicine library, v. 26.

Brill, Leiden, The Netherlands, 2011. ISBN 90-04-21871-8 (hardcover), 90-04-21872-6 (e-book). ISSN 1872-0684. xiii + 227 pp. LCCN B778 .H57 2011.

Henning:2015:TWA

[HMJ15] Brian G. Henning, William T. Myers, and Joseph David John, editors. Thinking with Whitehead and the American pragmatists: experience and reality. Lexington Books, Lexington, MA, USA, 2015. ISBN 0-7391-9031-8 (hardcover), 1-4985-1570-3 (paperback), 0-7391-9032-6 (electronic). xviii + 253 pp. LCCN B1674.W354 T48 2015.

Harrison:2011:WNO

[HNS11] Peter Harrison, Ronald L. Numbers, and Michael H. Shank, editors. Wrestling with nature: from omens to science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-31781-1, 0-226-31783-8, 1-283-36258-9 (e-book). x + 416 pp. LCCN Q125 .W86 2011.

Hofmann:2013:BRD

[Hof13] James R. Hofmann. Book review: Denis R. Alexander and Ronald L. Numbers, eds., Biology and Ideology from Descartes to Dawkins. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):175–179, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/667770.

Holland:2010:JWR

[Hol10] Jocelyn Holland. Johann Wilhelm Ritter (1776–1810): key texts on the science and art of nature, volume 16 of Medieval and early modern science. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-18367-1 (hardcover). xiv + 713 pp. LCCN Q113.

Holenstein:2013:SAP

[Hol13] André Holenstein. Scholars in action: the practice of knowledge and the figure of the savant in the 18th Century, volume 1; 34(1); 9(1) of History of science and medicine library; Scientific and learned cultures and their institutions. Brill, Leiden, The Netherlands, 2013. ISBN 90-04-24390-9 (set), 90-04-24947-8 (vol.1), 90-04-24948-6 (vol.2), 90-04-24391-7 (e-book). ISSN 1872-0684. xxx + 501 pp. LCCN AZ604 .S36 2013.

Holder:2023:PRO

[Hol23] Justin P. Holder. Poincaré’s radical ontology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):151–179, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Honenberger:2018:DAP

[Hon18] Phillip Honenberger. Darwin among the philosophers: Hull and Ruse on Darwin, Herschel, and Whewell. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2): 278–309, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hopwood:2024:SUH

[Hop24] Nick Hopwood. A sexless universe: How microbial genetics shaped the first history of reproduction, François Jacob’s The Logic of Life. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):511–534, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hoquet:2012:BRH

[Hoq12] Thierry Hoquet. Book review: Hans-Jörg Rheinberger, An Epistemology of the Concrete: Twentieth-Century Histories of Life. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):157–161, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/664461.

Horky:2013:PP

[Hor13] Phillip Sidney Horky. Plato and Pythagoreanism. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-989822-7, 0-19-989823-5 (e-book). xxi + 305 pp. LCCN QA22 .H67 2013; B243 .H67 2013.

Howard:2022:BWB

[How22] Stephen Howard. From the boundary of the world to the boundary of reason: The first antinomy and the development of Kant’s critical philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):225–241, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hodge:2002:CCD

[HR02] Michael Jonathan Sessions Hodge and Gregory Radick. The Cambridge companion to Darwin. Cambridge companions to philosophy. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-77197-8 (hardcover), 0-521-77730-5 (paperback). xiii + 486 pp. LCCN QH31.D2 C185 2002. URL http://catdir.loc.gov/

catdir/description/cam0210/2002031070.html;
http://catdir.loc.gov/catdir/toc/cam031/2002031070.html.

Heidelberger:2009:SHN

[HS09] Michael Heidelberger and Gregor Schiemann, editors. The Significance of the Hypothetical in the Natural Sciences. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2009. ISBN 1-282-71449-X. 376 pp. LCCN Q175.32.R45.

Hudson:2020:WWP

[Hud20] Robert Hudson. What was Perrin really doing in his proof of the reality of atoms? HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):194–218, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Hunter:2014:DPU

[Hun14] Cornelius Hunter. Darwin’s principle: The use of contrastive reasoning in the confirmation of evolution. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4 (1):106–149, Spring 2014. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/675437.

Ierodiakonou:2011:TEM

[IR11] Katerina Ierodiakonou and Sophie Roux, editors. Thought experiments in methodological and historical contexts, volume 19(15) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-20176-9 (hardcover). ISSN 1872-0684. LCCN BD265 .T46 2011. URL

http://catalogue.bnf.fr/ark:/12148/cb42470329f.

Irving:2012:RIN

[Irv12] Sarah Irving. Rethinking instrumentality: Natural philosophy and Christian charity in the early modern Atlantic world. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):55–76, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/662014.

Isaac:2012:WKM

[Isa12] Joel Isaac. Working knowledge: making the human sciences from Parsons to Kuhn. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06522-0. vi + 314 pp. LCCN H62 .I77 2012.

Itokazu:2014:BRP

[Ito14] Anastasia Guidi Itokazu. Book review: Patrick J. Boner, Kepler’s Cosmological Synthesis: Astrology, Mechanism and the Soul. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):203–206, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/675438.

Ivanova:2013:DPE

[Iva13] Milena Ivanova. Did Perrin’s experiments convert Poincaré to scientific realism? HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):1–19, Spring 2013. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/669714.

Ivanova:2017:BRD

[Iva17] Milena Ivanova. Book review: David Stump, Conceptual Change and the Philosophy of Science: Alternative Interpretations of the A Priori. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):151–153, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ivanova:2020:RRJ

[Iva20] Milena Ivanova. Reflections on the reception of Jean Perrin’s experiments by his contemporaries. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):219– 224, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ighbariah:2018:IAH

[IW18] Ahmad Ighbariah and Roy Wagner. Ibn al-Haytham’s revision of the Euclidean foundations of mathematics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1): 62–86, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Jalloh:2024:MCD

[Jal24] Mahmoud Jalloh. Metaphysics and convention in dimensional analysis, 1914–1917. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):275–322, Fall 2024. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Jesseph:2011:BRS

[Jes11] Douglas M. Jesseph. Book reviews: Stephen Gaukroger, The Emergence of a Scientific Culture: Science and the Shaping of Modernity, 1210–1685, Stephen Gaukroger, The Collapse of Mechanism and the Rise of Sensibility: Science and the Shaping of Modernity, 1680–1760. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):317–328, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/661292.

Jesseph:2014:BRB

[Jes14] Doug Jesseph. Essay review: The rise of the “mechanical philosophy”: Current scholarship on Seventeenth-Century science: Daniel Garber and Sophie Roux, The Mechanization of Natural Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):351–357, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/676672.

Jewett:2012:SDA

[Jew12] Andrew Jewett. Science, democracy, and the American university: from the Civil War to the Cold War. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-02726-8 (hardcover), 1-139-57115-X, 1-139-22571-5 (e-book). xii + 402 pp. LCCN Q127.U6 J49 2012.

Jha:2012:BRM

[Jha12] Stefania Ruzsits Jha. Book review: Mary Jo Nye, Michael Polanyi and His Generation: Origins of the Social Construction of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):368–372, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/665595.

Jorink:2012:NNH

[JM12] Eric Jorink and Ad Maas, editors. Newton and The Netherlands: how Isaac Newton was fashioned in the Dutch Republic. Leiden University Press, Amsterdam, The Netherlands, 2012. ISBN 90-8728-137-4. 256 pp. LCCN QC16.N7 N46

2012. URL http://www.loc.gov/catdir/enhancements/fy1407/
2013445090-b.html;   http://www.loc.gov/catdir/
enhancements/fy1407/2013445090-d.html.

Jungnickel:2017:SPH

[JM17] Christa Jungnickel and Russell McCormmach. The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-49564-X (hardcover), 3-319-49565-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxxi + 460 pp. LCCN QC9.G3 J86 2017.

Johnson:2010:NAD

[Joh10] Dirk R. Johnson. Nietzsche’s Anti-Darwinism. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19678-7 (hardcover), 1-107-62152-6 (paperback), 0-511-76221-6 (e-book). x + 240 pp.

LCCN B3317 .J64 201. URL http://www.survey.bris.ac.uk/
standrews/cambridgebooksonlinetrial.

Jones:2013:BRB

[Jon13] Jan-Erik Jones. Book review: Sorana Corneanu, Regimens of the Mind: Boyle, Locke and the Early Modern Cultura Animi Tradition. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):371–374, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/671742.

Kindi:2012:KSS

[KA12] Vasso Kindi and Theodore Arabatzis, editors. Kuhn’s The structure of scientific revolutions revisited, volume 12 of Routledge studies in the philosophy of science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2012. ISBN 0-415-80855-3 (hardcover), 0-203-10315-7 (e-book). LCCN Q175.K953 K84 2012.

Kasser:2024:BRR

[Kas24] Jeff Kasser. Book review: Robert Talisse, Paniel Reyes Cárdenas, and Daniel Herbert, eds., Pragmatic Reason: Christopher Hookway and the American Philosophical Tradition. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14 (1):225–229, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kather:2012:BRM

[Kat12] Regine Kather. Book review: Marco Böhlandt, Verborgene Zahl-Verborgener Gott: Mathematik und Naturwissen im Denken des Nicolaus Cusanus (1401–1464). HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):189–192, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/662662.

Kendig:2016:NKC

[Ken16] Catherine Kendig, editor. Natural Kinds and Classification in Scientific Practice. History and philosophy of biology. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2016. ISBN 1-84893-540-4 (hardcover), 1-317-21541-9 (e-book), 1-317-21543-5 (ebook), 1-78144-732-2 (e-book), 1-78144-733-0 (e-book). xx + 247 pp. LCCN Q177 .N38 2016.

Kendig:2018:BRS

[Ken18] Catherine Kendig. Book review: Scott Lidgard and Lynn K. Nyhart, eds., Biological Individuality: Integrating Scientific, Philosophical, and Historical Perspectives. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):475–480, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kent:2022:BRK

[Ken22] Rory Kent. Book review: Karim Bschir and Jamie Shaw, eds., Interpreting Feyerabend: Critical Essays. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12 (1):308–311, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kuhn:2012:SSR

[KH12] Thomas S. Kuhn and Ian Hacking. The Structure of Scientific Revolutions. University of Chicago Press, Chicago, IL, USA, fourth edition, 2012. ISBN 0-226-45811-3 (hardcover), 0-226-45812-1 (paperback), 0-226-45814-8 (e-book). xlvi + 217 pp. LCCN Q175 .K95 2012. With an introductory essay by Ian Hacking.

Kinzel:2016:BRL

[Kin16] Katherina Kinzel. Book review: Léna Soler, Emiliano Trizio, and Andrew Pickering, eds., Science as It Could Have Been: Discussing the Contingency/Inevitability Problem. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2): 319–323, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kindi:2018:BRJ

[Kin18a] Vasso Kindi. Book review: James A. Marcum, Thomas Kuhn’s Revolutions: a Historical and an Evolutionary Philosophy of Science? HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):233–236, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kinzel:2018:IEA

[Kin18b] Katherina Kinzel. Inner experience and articulation: Wilhelm Dilthey’s foundational project and the charge of psychologism. HO-POS: Journal of the International Society for the History of Philosophy of Science, 8(2):347–375, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kissel:2024:BRJ

[Kis24] Teresa Kouri Kissel. Book review: Jeanne Peijnenburg and Sander Verhaegh, eds., Women in the History of Analytic Philosophy: Selected Papers of the Tilburg-Groningen Conference, 2019. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):215–219, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kjaergaard:2011:BRR

[Kjæ11] Peter C. Kjærgaard. Book review: Robert J. Richards, The Tragic Sense of Life: Ernst Haeckel and the Struggle over Evolutionary Thought. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):149–152, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/659037.

Konvitz:1960:APS

[KK60] Milton R. (Milton Ridvas) Konvitz and Gail Kennedy, editors. The American pragmatists: selected writings. Meridian books, M105. Meridian Books, New York, NY, USA, 1960. 413 pp. LCCN B832 .K6.

Klein:2012:BRF

[Kle12] Alexander Klein. Book review: Francesca Bordogna, William James at the Boundaries: Philosophy, Science, and the Geography of Knowledge. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):161–166, Spring 2012. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/661794.

Klein:2015:HAC

[Kle15a] Alexander Klein. Hatfield on American critical realism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):154–166, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/680374.

Klein:2015:SRW

[Kle15b] Alexander Klein. Science, religion, and “the will to believe”. HO-POS: Journal of the International Society for the History of Philosophy of Science, 5(1):72–117, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/680371.

Klev:2018:RMD

[Kle18] Ansten Klev. A road map of Dedekind’s theorem 66. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):241–277, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kwa:2011:SKN

[KM11] Chunglin Kwa and David McKay. Styles of knowing: a new history of science from ancient times to the present. University of Pittsburgh Press, Pittsburgh, PA, USA, 2011. ISBN 0-8229-6151-2 (paperback), 0-8229-7774-5 (e-book). viii + 366 pp. LCCN Q125 .K83 2011.

Kober:2014:BRR

[Kob14] Gal Kober. Book review: Richard A. Richards, The Species Problem: A Philosophical Analysis. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):169–172, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675455.

Kochiras:2013:CLS

[Koc13] Hylarie Kochiras. Causal language and the structure of force in Newton’s system of the world. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):210–235, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/670327.

Koertge:2013:FFP

[Koe13] Noretta Koertge. Feyerabend, feminism, and philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):139–141, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/669955.

Koetschet:2024:PAA

[Koe24] Pauline Koetschet. Peter Adamson, Al-Razi. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):236–239, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kooijmans:2010:DDA

[Koo10] L. Kooijmans. Death defied: the anatomy lessons of Frederik Ruysch, volume 18 of History of science and medicine library. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-18784-7 (hardcover), 90-04-18856-8 (e-book). ISSN 1872-0684. xvi + 470 pp. LCCN QM16.R89 K6613 2010.

Korhonen:2018:BRA

[Kor18] Anssi Korhonen. Book review: Aaron Preston, ed., Analytic Philosophy: an Interpretive History. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):222– 226, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kragh:2022:PCS

[Kra22a] Helge Kragh. Philosophical contexts of the steady-state universe. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):129–145, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Krauss:2022:BRJ

[Kra22b] Chiara Russo Krauss. Book review: John Preston, ed., Interpreting Mach: Critical Essays. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):319–322, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Krieger:2011:SFP

[Kri11] William Krieger. Science at the frontiers: perspectives on the history and philosophy of science. Lexington Books, Lexington, MA, USA, 2011. ISBN 0-7391-5014-6 (hardcover), 0-7391-5015-4 (paperback). xiv + 231 pp. LCCN Q175.3. URL http://covers.

rowmanlittlefield.com/L/07/391/0739150146.jpg.

Krieger:2012:TLM

[Kri12] William H. Krieger. Theory, locality, and methodology in archaeology: Just add water? HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):243–257, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/666956.

Kuby:2018:CFP

[Kub18] Daniel Kuby. Carnap, Feyerabend, and the pragmatic theory of observation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):432–470, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kuhn:1962:SSR

[Kuh62] Thomas S. Kuhn. The Structure of Scientific Revolutions. University of Chicago Press, Chicago, IL, USA, 1962. xv + 172 pp. LCCN Q121 .I5 vol. 2, no. 2; Q175 .K95.

Kuhn:1970:SSR

[Kuh70] Thomas S. Kuhn. The Structure of Scientific Revolutions. International encyclopedia of unified science. Foundations of the unity of science, v. 2, no. 2. University of Chicago Press, Chicago, IL, USA, second enlarged edition, 1970. ISBN 0-226-45803-2 (hardcover), 0-226-45804-0 (paperback), 0-226-45807-5. xii + 210 pp. LCCN Q175 .K95 1970.

Kuhn:1996:SSR

[Kuh96] Thomas S. Kuhn. The Structure of Scientific Revolutions. University of Chicago Press, Chicago, IL, USA, third edition, 1996. ISBN 0-226-45807-5 (hardcover), 0-226-45808-3 (paperback). xiv + 212

pp. LCCN Q175 .K95 1996. URL http://www.loc.gov/catdir/
description/uchi051/96013195.html;
http://www.loc.gov/catdir/toc/uchi051/96013195.html.

Kusch:2019:VSK

[Kus19] Martin Kusch. From Völkerpsychologie to the sociology of knowledge. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):250–274, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kuukkanen:2013:BRA

[Kuu13] Jouni-Matti Kuukkanen. Book review: Alfred I. Tauber, Science and the Quest for Meaning. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):157–160, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/668912.

Katzav:2022:RLE

[KV22] Joel Katzav and Krist Vaesen. The rise of logical empiricist philosophy of science and the fate of speculative philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2):327–358, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Kendall:2010:BBN

[KW10] Calvin B. Kendall and Faith Wallis. Bede: On the nature of things and On times, volume 56 of Translated texts for historians. Liverpool University Press, Liverpool, UK, 2010. ISBN 1-84631-495-X (hardcover), 1-84631-496-8 (paperback). xiv + 222 pp. LCCN

PA8260 .A6 2010. URL http://www.gbv.de/dms/bowker/toc/
9781846314957.pdf.

Laerke:2018:BRD

[Lær18] Mogens Lærke. Book review: Dmitri Levitin, Ancient Wisdom in the Age of the New Science: Histories of Philosophy in England, c. 1640–1700. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):209–213, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lagerlund:2014:BRC

[Lag14] Henrik Lagerlund. Book review: Calvin B. Kendall and Faith Wallis, Bede: “On the Nature of Things” and “On Times”. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):206–208, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/674980.

Landini:2016:BRE

[Lan16] Gregory Landini. Book review: Erik C. Banks, The Realistic Empiricism of Mach, James and Russell. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2): 329–333, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lancaster:2019:BRD

[Lan19] James A. T. Lancaster. Book review: Daniel McKaughan and Holly VandeWall, eds., The History and Philosophy of Science: a Reader. London: Bloomsbury Academic, 2018. Pp. xxiii + 1073. $49.95 (paper). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):211–214, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lattis:2011:BRD

[Lat11] James Lattis. Book review: David Aubin, Charlotte Bigg, and H. Otto Sibum, eds., The Heavens on Earth: Observatories and Astronomy in Nineteenth-Century Science and Culture. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):146–149, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/659004.

Lattis:2014:BRE

[Lat14] James Lattis. Book review: Emma Gee, Aratus and the Astronomical Tradition. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):387–390, Fall 2014. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/677343.

Lauer:2024:BRA

[Lau24] Richard Lauer. Book review: Adam Tamas Tuboly, ed., The History of Understanding in Analytic Philosophy: Around Logical Empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):253–256, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lavers:2020:BRF

[Lav20] Gregory Lavers. Book review: Frederique Janssen-Lauret and Gary Kemp, eds., Quine and His Place in History. New York: Palgrave Macmillan, 2016. Pp. xii + 224. $110.00 (cloth); $85.00 (e-book).

ISBN 978-1-137-47250-2. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):357–361, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Legg:2024:BRT

[Leg24] Catherine Legg. Book review: Trevor Pearce, Pragmatism’s Evolution: Organism and Environment in American Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):557–560, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lennox:2011:ANI

[Len11] James G. Lennox. Aristotle on norms of inquiry. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):23–46, Spring 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/658482.

Lennox:2014:AEM

[Len14] James G. Lennox. Aristotle on the emergence of material complexity: Meteorology IV and Aristotle’s biology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):272–305, Fall 2014. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/677568.

Leroux:2010:HCPa

[Ler10a] Jean Leroux. Une histoire comparée de la philosophie des sciences. Volume I, Aux sources du Cercle de Vienne. (French) [A comparative history of the philosophy of science. Volume I, The sources of the Vienna Circle]. Logique de la science. Presses de l’Université Laval, Laval, Québec, Canada, 2010. ISBN 2-7637-8788-6 (paperback). x + 190 pp. LCCN B824.6 L47 2010; Q174.8 .L47 2010 v. 1.

URL http://catalogue.bnf.fr/ark:/12148/cb424331540.

Leroux:2010:HCPb

[Ler10b] Jean Leroux. Une histoire comparée de la philosophie des sciences. Volume II, L’empirisme logique en débat. (French) [A comparative history of the philosophy of science. Volume II. The logical empiricism debate]. Logique de la science. Presses de l’Université Laval, Laval, Québec, Canada, 2010. ISBN 2-7637-8957-9 (paperback). xi + 187 pp. LCCN B824.6 L47 2010; Q174.8 .L47 2010 v. 2. URL

http://catalogue.bnf.fr/ark:/12148/cb42252827v.

Leroux:2011:BRT

[Ler11] Jean Leroux. Book review: Teresa Castelão, Gaston Bachelard et les études critiques de la science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):139–142, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/659771.

Levitin:2015:AWA

[Lev15] Dmitri Levitin. Ancient Wisdom in the Age of the New Science: Histories of Philosophy in England, C. 1640–1700. Ideas in context. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-10588-9. xii + 670 pp. LCCN B1131 .L48

2015. URL http://www.loc.gov/catdir/enhancements/fy1508/
2015008921-b.html;   http://www.loc.gov/catdir/
enhancements/fy1508/2015008921-d.html;   http://www.
loc.gov/catdir/enhancements/fy1508/2015008921-t.html.

Lagerlund:2016:RCS

[LH16] Henrik Lagerlund and Benjamin Hill. Routledge Companion to Sixteenth Century Philosophy. Routledge Philosophy Companions. Routledge, New York, NY, USA, 2016. ISBN 0-415-65860-8, 1-317-67261-5. xiv + 645 pp. LCCN B770.R68 2016. URL http://

public.eblib.com/choice/publicfullrecord.aspx?p=4779398.

Lindquist:2018:HIL

[Lin18] Daniel Lindquist. Hegel’s “idea of life” and internal purposiveness. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):376–408, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lloyd:2013:CFM

[Llo13] Elisabeth A. Lloyd. Constitutional failures of meritocracy and their consequences. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):142–144, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/669956.

Lawrence:2015:MTG

[LM15] Snezana Lawrence and Mark McCartney, editors. Mathematicians and their Gods: interactions between mathematics and religious beliefs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-870305-8, 0-19-100754-4 (PDF with Adobe DRM). vi + 298 pp. LCCN BL265.M3 M38

2015. URL http://www.loc.gov/catdir/enhancements/fy1604/
2014956601-b.html;   http://www.loc.gov/catdir/
enhancements/fy1604/2014956601-d.html;   http://www.
loc.gov/catdir/enhancements/fy1604/2014956601-t.html.

Levine:2012:DCE

[LNN12] Alex Levine, Adriana Novoa, and Adriana Novoa, From man to ape. ‘!Darwinistas!: the construction of evolutionary thought in nineteenth century Argentina, volume 27(5) of History of science and medicine library; Scientific and learned cultures and their institutions. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-22136-0 (hardcover). ISSN 1872-0684. xvii + 279 pp. LCCN QH361 .L48 2012.

Lolordo:2015:BRT

[Lol15] Antonia Lolordo. Book review: Tad M. Schmaltz, ed., Efficient Causation: A History. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):356–360, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/682376.

Longino:2013:DP

[Lon13] Helen E. Longino. Data, please. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):144–146, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/669957.

Long:2023:BRC

[Lon23] Brandon Long. Book review: E. Claire Cage, The Science of Proof: Forensic Medicine in Modern France. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1): 199–202, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Look:2011:BRT

[Loo11] Brandon C. Look. Book review: Tom Sorell, G. A. J. Rogers, and Jill Kaye, eds., Scientia in Early Modern Philosophy: Seventeenth-Century Thinkers on Demonstrative Knowledge from First Principles. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):367–371, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/661233.

Love:2016:BRS

[Lov16] Alan C. Love. Book review: Susannah Gibson, Animal, Vegetable, Mineral? How Eighteenth-Century Science Disrupted the Natural Order. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):337–340, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Lund:2010:RHO

[Lun10] Matthew D. Lund. N. R. Hanson: observation, discovery, and scientific change. Humanity Books, Amherst, NY, USA, 2010. ISBN 1-59102-772-1 (paperback). 253 pp. LCCN Q175 .L846 2010.

Luthy:2012:DGE

[Lüt12a] Christoph Herbert Lüthy. David Gorlæus (1591–1612): an enigmatic figure in the history of philosophy and science, volume 13 of History of science and scholarship in the Netherlands. Amsterdam University Press, Amsterdam, The Netherlands, 2012. ISBN 90-8964-438-5 (hardcover), 90-485-1680-3 (eISBN: PDF), 90-485-1681-1 (eISBN: ePub). ISSN 1569-3481. 225 pp. LCCN QC16.G635 L87 2012.

Lutz:2012:SMP

[Lut12b] Sebastian Lutz. On a straw man in the philosophy of science: A defense of the received view. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):77–120, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/664460.

Luthy:2014:BRO

[Lüt14] Christoph Lüthy. Book review: Ofer Gal and Raz Chen-Morris, Baroque Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):379–382, Fall 2014. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/677320.

Lutz:2024:BRT

[Lut24] Sebastian Lutz. Book review: Thomas Uebel and Christoph Limbeck-Lilienau, eds., The Routledge Handbook of Logical Empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):641–644, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

MacKay:1982:SQM

[Mac82] Donald MacCrimmon MacKay. Science and the quest for meaning. Eerdmans, Grand Rapids, MI, USA, 1982. ISBN 0-8028-1914-1. xi + 75 pp. LCCN BL241 .M18 1982.

Maddy:2014:LMW

[Mad14] Penelope Maddy. The logical must: Wittgenstein on logic. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-939175-0 (hardcover). x + 135 pp. LCCN B3376.W564 M155 2014.

Magdalino:2012:BRE

[Mag12] Paul Magdalino. Book review: Efthymios Nicolaidis, Science and Eastern Orthodoxy: From the Greek Fathers to the Age of Globalization. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):376–380, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/666864.

Magnus:2015:JSM

[Mag15] P. D. Magnus. John Stuart Mill on taxonomy and natural kinds. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):269–280, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/682373.

MaiaNeto:2014:ASS

[Mai14] Jose R. Maia Neto. Academic skepticism in seventeenthcentury French philosophy: the Charronian legacy, 1601–1662. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-07358-3 (hardcover), 3-319-07359-1 (e-book). xii + 165 pp. LCCN B1818.S3 M35

2014. URL http://www.loc.gov/catdir/enhancements/fy1411/
2014943514-b.html;   http://www.loc.gov/catdir/
enhancements/fy1411/2014943514-d.html;   http://www.
loc.gov/catdir/enhancements/fy1411/2014943514-t.html.

McGrew:2009:PSH

[MAKA09] Timothy J. McGrew, Marc Alspector-Kelly, and Fritz Allhoff. Philosophy of science: an historical anthology, volume 30 of Blackwell Philosophy Anthologies. Wiley-Blackwell, Chichester, UK, 2009. ISBN 1-4051-7543-5 (hardcover), 1-4051-7542-7 (paperback). xix + 1 + 660 pp. LCCN Q175 .P51235 2009.

Manninen:2009:VCN

[Man09] Juha Manninen. The Vienna Circle in the Nordic countries: networks and transformations of logical empiricism, volume 14 of Vienna Circle Institute Yearbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-3682-2, 90-481-3683-0 (e-book). 326 pp. LCCN B824.6 V54 2010.

Manning:2015:DMB

[Man15] Gideon Manning. Descartes’s metaphysical biology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):209–239, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/682372.

Marr:2010:BRG

[Mar10] Alexander Marr. Between Raphael and Galileo: Mutio Oddi and the mathematical culture of late Renaissance Italy. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-50628-2 (hardcover). xiii + 359 pp. LCCN QA27.I8.

Martin:2011:RMP

[Mar11] Craig Martin. Renaissance meteorology: Pomponazzi to Descartes. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0187-8. viii + 213 pp. LCCN QC855 .M37 2011.

Martin:2014:SAR

[Mar14] Craig Martin. Subverting Aristotle: religion, history, and philosophy in early modern science. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1316-7 (hardcover), 1-4214-1317-5 (e-book). 262 pp. LCCN B485 .M267 2014. URL

http://muse.jhu.edu/books/9781421413174/.

Marcum:2015:TKR

[Mar15] James A. Marcum. Thomas Kuhn’s Revolutions: a Historical and an Evolutionary Philosophy of Science? Bloomsbury Academic, an imprint of Bloomsbury Publishing Plc, London, UK, 2015. ISBN 1-4725-3049-7 (paperback), 1-4725-2208-7 (epub), 1-4725-3040-3 (epdf). viii + 287 pp. LCCN Q175.K953 M37 2015.

Marshall:2017:BRG

[Mar17] Douglas Bertrand Marshall. Book review: Geoffrey Gorham, Benjamin Hill, Edward Slowik, and C. Kenneth Waters, eds., The Language of Nature: Reassessing the Mathematization of Natural Philosophy in the Seventeenth Century. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2): 383–386, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Marinucci:2021:CHN

[Mar21] Ludovica Marinucci. Christiaan Huygens’s natural theology in his Cosmotheoros and other late writings. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2): 642–659, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Marschall:2024:HGB

[Mar24] Benjamin Marschall. Herbert G. Bohnert: The last Carnapian. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(2):361–396, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Massimi:2008:KPS

[Mas08] Michela Massimi. Kant and philosophy of science today, volume 63 of Royal Institute of Philosophy supplement. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-73544-0 (paperback), 0-521-74851-8. vii + 204 pp. LCCN B2798 .K36 2008.

Matsui:2024:SAD

[Mat24] Takaaki Matsui. Sellars, analyticity, and a dynamic picture of language. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):78–102, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

McCaskey:2015:BRM

[McC15] John P. McCaskey. Book review: Marco Sgarbi, The Aristotelian Tradition and the Rise of British Empiricism: Logic and Epistemology in the British Isles, 1570–1689. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5 (1):204–207, Spring 2015. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/680379.

McCoy:2022:CWP

[McC22] C. D. McCoy. The constitution of Weyl’s pure infinitesimal world geometry. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):189–208, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

McConaughey:2024:BRL

[McC24] Zoe McConaughey. Book review: Lukas M. Verburgt and Matteo Cosci, eds., Aristotle’s Syllogism and the Creation of Modern Logic: Between Tradition and Innovation, 1820s–1930s. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):232–236, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

McDonald:2014:BRM

[McD14] Patrick J. McDonald. Book review: Michel Meulders and Laurence Garey, eds., Helmholtz: From Enlightenment to Neuroscience. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):182–186, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/672154.

McElvenny:2018:LMA

[McE18] James McElvenny. Language and meaning in the age of modernism: C. K. Ogden and his contemporaries. Edinburgh University Press, Edinburgh, Scotland, 2018. ISBN 1-4744-2503-8. 188 (est.) pp. LCCN P107 .M374 2018.

McMullin:2011:KME

[McM11] Ernan McMullin. Kepler: Moving the Earth. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):3–22, Spring 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/658917.

McNulty:2016:CKO

[McN16] Michael Bennett McNulty. Chemistry in Kant’s Opus Postumum. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):64–95, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

McNulty:2022:KMD

[McN22] Michael Bennett McNulty. Kant on the mathematical deficiency of psychology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2):485–509, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Malaterre:2019:WTC

[MCP19] Christophe Malaterre, Jean-François Chartier, and Davide Pulizzotto. What is this thing called Philosophy of Science? A computational topic-modeling perspective, 1934–2015. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):215–249, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Meer:2024:ISA

[Mee24] Rudolf Meer. Ilse Schneider (and Alois Riehl) on the space–time problem in Kant and Einstein: New perspectives on neo-Kantianism and positivism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):508–526, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Mehdi:2014:BRR

[Meh14] Mohamed Mehdi. Book review: Richard Bett, ed., The Cambridge Companion to Ancient Scepticism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):208–212, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675446.

Meli:2017:BRD

[Mel17] Domenico Bertoloni Meli. Book review: Delphine Antoine-Mahut and Stephen Gaukroger, eds., Descartes’ “Treatise on Man” and Its Reception. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):391–395, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Melogno:2019:DJD

[Mel19] Pablo Melogno. The discovery-justification distinction and the new historiography of science: On Thomas Kuhn’s Thalheimer lectures. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):152–178, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Mensch:2013:KOE

[Men13] Jennifer Mensch. Kant’s organicism: epigenesis and the development of critical philosophy. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-02198-X (hardcover), 0-226-02203-X (e-book). xi + 246 pp. LCCN B2799.N37 M46 2013.

Mensch:2015:BRI

[Men15] Jennifer Mensch. Book review: Ina Goy and Eric Watkins, eds., Kant’s Theory of Biology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):367–370, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/682712.

Methot:2024:JVM

[Mét24] Pierre-Olivier Méthot. Jacob versus Monod on the natural selection of ideas. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):492–510, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Meulders:2010:HEN

[Meu10] Michel Meulders, editor. Helmholtz: from enlightenment to neuroscience. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01448-3 (hardcover). xvii + 235 pp. LCCN Q143.H5 M4813 2010.

Mihaljevic:2024:BBA

[Mih24] Katarina Mihaljević. Breaking into British academic life in Second World War Britain: The story of Rose Rand. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):297–316, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Miller:2014:BRD

[Mil14a] Joseph G. Miller. Book review: Daniel Graham, Science before Socrates: Parmenides, Anaxagoras, and the New Astronomy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):212–215, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/673862.

Miller:2014:BRP

[Mil14b] Joseph G. Miller. Book review: Phillip Sidney Horky, Plato and Pythagoreanism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):391–393, Fall 2014. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/676673.

Mills:2015:BRC

[Mil15a] Ethan Mills. Book review: Christopher I. Beckwith, Warriors of the Cloisters: The Central Asian Origins of Science in the Medieval World. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):201–204, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/680378.

Mills:2015:EMC

[Mil15b] M. Anthony Mills. Explicating Meyerson: The critique of positivism and historical Épistémologie. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):318–347, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/682423.

Milkov:2016:WDP

[Mil16] Nikolay Milkov. Walter Dubislav’s philosophy of science and mathematics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):96–116, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Miller:2020:BRM

[Mil20] David Marshall Miller. Book review: Michael J. Sauter, The Spatial Reformation: Euclid between Man, Cosmos, and God. Philadelphia: University of Pennsylvania Press, 2019. Pp. 327. $89.95 (cloth). ISBN 978-0-812-25066-4. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):601–605, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Mirus:2011:BRT

[Mir11] Christopher V. Mirus. Book review: Timothy McGrew, Marc Alspector-Kelly, and Fritz Allhoff, eds., Philosophy of Science: An Historical Anthology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):132–135, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/658287.

Mirowski:2017:BRP

[Mir17] Philip Mirowski. Book review: Paul Erickson, The World the Game Theorists Made. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):160–163, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Misak:2013:AP

[Mis13] Cheryl J. (Cheryl J.) Misak. The American pragmatists. The Oxford history of philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-923120-6 (hardcover). xvi + 286 pp. LCCN B832 .M57 2013.

Misak:2015:KJW

[Mis15] Cheryl Misak. Klein on James on the will to believe. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):118–128, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/680372.

Mormann:2013:IIN

[MK13] Thomas Mormann and Mikhail Katz. Infinitesimals as an issue of neo-Kantian philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):236–280, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/671348.

Machamer:2009:DCM

[MM09] Peter K. Machamer and J. E. McGuire. Descartes’s changing mind. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13889-3. xi + 258 pp. LCCN B1878.M55 M33 2009.

Marren:2023:PMD

[MM23] Marina Marren and Kevin Marren. Philosophical method of Dioscorides’s De Materia Medica. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1): 180–198, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Monagle:2015:BRD

[Mon15] Clare Monagle. Book review: David Bloch, John of Salisbury on Aristotelian Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):387–389, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/682381.

Mormann:2017:PFA

[Mor17] Thomas Mormann. Philipp Frank’s Austro–American logical empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):56–87, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Mormann:2024:BRC

[Mor24] Thomas Mormann. Book review: Christian Damböck, Günther Sandner, and Meike G. Werner, eds., Logischer Empirismus, Lebensreform und die deutsche Jugendbewegung: Logical Empiricism, Life Reform, and the German Youth Movement. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14 (2):613–616, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Mosley:2014:BRB

[Mos14] Adam Mosley. Book review: André Goddu, Copernicus and the Aristotelian Tradition: Education, Reading, and Philosophy in Copernicus’s Path to Heliocentrism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1): 199–203, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-

6240 (electronic). URL http://www.jstor.org/stable/10.1086/
672155.

Milkov:2013:BGP

[MP13] Nikolay Milkov and Volker Peckhaus. The Berlin Group and the Philosophy of Logical Empiricism, volume 273 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-5484-1 (hardcover), 94-007-5485-X (e-book). ISSN 0068-0346. x + 332 pp. LCCN Q175.32.A24 .M384 2011; Q175 .B47

2013. URL https://link.springer.com/book/10.1007/978-94-
007-5485-0.

Montminy:2018:BRK

[MP18] David Montminy and François Papale. Book review: Kevin C. Elliott, A Tapestry of Values: an Introduction to Values in Science. Oxford: Oxford University Press, 2017. Pp. 224. $105.00 (cloth); $35.00 (paper). HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):500–504, Fall 2018. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Morris-Reich:2017:IRHi

[MRR17] Amos Morris-Reich and Dirk Rupnow, editors. Ideas of ‘Race’ in the History of the Humanities. Palgrave Critical Studies of Antisemitism and Racism. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-49952-1, 3-319-49953-X (e-book). xiii + 337 + 12 pp. LCCN D203.2-475. URL http://link.

springer.com/10.1007/978-3-319-49953-6.

Muntersbjorn:2012:BRA

[Mun12] Madeline Muntersbjorn. Book review: Alexander Marr, Between Raphael and Galileo: Mutio Oddi and the Mathematical Culture of Late Renaissance Italy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):391–394, Fall 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/666969.

Muntersbjorn:2016:BRS

[Mun16] Madeline Muntersbjorn. Book review: Snezana Lawrence and Mark McCartney, eds., Mathematicians and Their Gods: Interactions between Mathematics and Religious Beliefs. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2): 333–336, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Murphy:2024:ACS

[Mur24] Claire Murphy. The apostle of “Common Sense”: The historical roots of Duhem’s distinction between physics and metaphysics. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(2):554–571, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

McKaughan:2018:HPS

[MV18] Daniel J. (Daniel Jon) McKaughan and Holly R. VandeWall, editors. The History and Philosophy of Science: a Reader. Bloomsbury Academic, London, UK, 2018. ISBN 1-4742-3272-8 (paperback), 1-4742-3273-6 (hardcover), 1-4742-3274-4 (ePub e-book), 1-4742-3275-2 (PDF ebook). xxiii + 1073 pp. LCCN Q175 .H497 2018.

Methot:2024:SSI

[MV24] Pierre-Olivier Méthot and Florence Vienne. Special section introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):454–462, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Muller-Wille:2012:CHH

[MWR12] Staffan Müller-Wille and Hans-Jörg Rheinberger. A cultural history of heredity. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-54570-9 (hardcover). xiii + 323 pp. LCCN

QH438.5 .M8513 2012. URL http://catalogue.bnf.fr/ark:/
12148/cb43636679x.

Nasim:2016:BRA

[Nas16] Omar W. Nasim. Book review: Alan G. Gross and Joseph E. Harmon, Science from Sight to Insight: How Scientists Illustrate Meaning. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):168–171, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Nemeth:2012:BRJ

[Nem12] Elisabeth Nemeth. Book reviews: Jean Leroux, Aux sources du Cercle de Vienne. Jean Leroux, L’empirisme logique en débat. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):147–154, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/664462.

Nemeth:2016:ELE

[Nem16] Elisabeth Nemeth. Embedding logical empiricism into the history of epistemology: Eino Kaila on human knowledge. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):148–157, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Neuber:2018:HKR

[Neu18] Matthias Neuber. Helmholtz, Kaila, and the representational theory of measurement. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):409–431, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Neuber:2020:TFA

[Neu20] Matthias Neuber. Two forms of American critical realism: Perception and reality in Santayana/Strong and Sellars. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):76–105, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Newton:2014:BRW

[New14] Alexandra Newton. Book review: Wayne Waxman, Kant’s Anatomy of the Intelligent Mind. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):375–378, Fall 2014.

CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/677321.

Nguyen:2023:CNC

[Ngu23] Jen Nguyen. Contextualizing Newton and Clarke’s “Argument from Quantity”. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):1–23, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Nicolaidis:2011:SEO

[Nic11] Efthymios Nicolaïdis. Science and Eastern orthodoxy: from the Greek fathers to the age of globalization. Medicine, science, and religion in historical context. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0298-X (hardcover), 1-4214-0426-5 (e-book). xviii + 252 + 16 pp. LCCN BX342.9.S35 N55 2011.

Nickles:2012:BRM

[Nic12] Thomas Nickles. Book review: Matthew Lund, N. R. Hanson: Observation, Discovery, and Scientific Change. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):364–368, Fall 2012. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/664731.

Niemann:2014:KPT

[Nie14] Hans-Joachim Niemann. Karl Popper and the two new secrets of life: including Karl Popper’s Medawar lecture 1986 and three related texts. Mohr Siebeck, Tübingen, Germany, 2014. ISBN 3-16-153207-4 (paperback), 3-16-153455-7 (e-book). vii + 157 pp. LCCN QH360.5 .N54 2014.

Nicholls:2012:TUN

[NL12] Angus James Nicholls and Martin Liebscher, editors. Thinking the unconscious: Nineteenth-Century German thought. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-41176-9 (paperback). ix + 329 pp. LCCN B2741 T443 2012.

Noble:2019:LDP

[Nob19] Christopher P. Noble. Leibniz on the divine preformation of souls and bodies. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):327–342, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Nordmann:2018:BRM

[Nor18] Alfred Nordmann. Book review: Michael Heidelberger, Helmut Pulte, and Gregor Schiemann, eds., Hermann von Helmholtz: Philosophische und Populärwissenschaftliche Schriften. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):495–497, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Nobre:2022:QDS

[NP22] Bruno Nobre and João Carlos Onofre Pinto. The quest for the dynamic structure of reality: Xavier Zubiri, phenomenology, and quantum mechanics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):22–42, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Nachtomy:2014:LSE

[NS14] Ohad Nachtomy and Justin E. H. Smith, editors. The life sciences in early modern philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-998731-9 (hardcover). xiii + 256 pp. LCCN QH331 .L538 2014.

Nye:2011:MPH

[Nye11] Mary Jo Nye. Michael Polanyi and his generation: origins of the social construction of science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-61063-2 (hardcover), 0-226-61065-9 (ebook). xxi + 405 pp. LCCN B945.P584 N94 2011.

Naaman-Zauderer:2010:DDT

[NZ10] Noa Naaman-Zauderer. Descartes’ Deontological Turn: Reason, Will, and Virtue in the Later Writings. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-511-85878-7, 0-521-76330-4. xii + 224 pp. LCCN B1878.W5N33 2010.

Oberheim:2014:BRV

[Obe14] Eric Oberheim. Book review: Vasso Kindi and Theodore Arabatzis, eds. Kuhn’s “The Structure of Scientific Revolutions” Revisited. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):157–161, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/674981.

Olen:2016:FPM

[Ole16] Peter Olen. From formalism to psychology: Metaphilosophical shifts in Wilfrid Sellars’s early works. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):24–63, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Omodeo:2021:HAF

[Omo21] Pietro Daniel Omodeo. Heavenly animation as the foundation for Fracastoro’s homocentrism: Aristotelian–Platonic eclecticism beyond the School of Padua. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):585–603, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Olen:2017:PTC

[OS17] Peter Olen and Carl Sachs, editors. Pragmatism in Transition: Contemporary Perspectives on C. I. Lewis. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-52862-9 (hardcover), 3-319-52863-7 (e-book). viii + 222 pp. LCCN B945.L454 P73 2017.

Ott:2009:CLN

[Ott09]  Walter R. Ott.  Causation and laws of nature in early

modern philosophy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-157140-7 (ebook), 0-19-957043-4. xii + 260 pp. LCCN BD541 .O83

2009. URL http://www.oxfordscholarship.com/oso/public/
content/philosophy/9780199570430/toc.html.

Padovani:2011:BRH

[Pad11] Flavia Padovani. Book review: Hans Reichenbach, ed., The Concept of Probability in the Mathematical Representation of Reality. HO-POS: Journal of the International Society for the History of Philosophy of Science, 1(2):344–347, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/661509.

Palmer:2015:BRH

[Pal15] Eric Palmer. Essay review: How to succeed in science while really, really trying: The central European savant of the Mid-Eighteenth Century: André Holenstein, Hubert Steinke, and Martin Stuber, eds. Scholars in Action: The Practice of Knowledge and the Figure of the Savant in the 18th Century. 2 vols. Leiden: Brill, 2013.

Pp. xxx + 932. $318.00 (cloth). HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):167–173, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/680375.

Parsons:2011:BRF

[Par11] Keith M. Parsons. Book review: Friedel Weinert, Copernicus, Darwin, and Freud: Revolutions in the History and Philosophy of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):351–355, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/660792.

Patton:2011:RE

[Pat11] Lydia Patton. Reconsidering experiments. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1 (2):209–226, Fall 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/660167.

Patton:2014:PSH

[Pat14] Lydia Patton. Philosophy, science and history: a guide and reader. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2014. ISBN 0-415-89831-5 (paperback), 0-415-89830-7 (hardcover), 0-203-80245-4 (e-book), 1-136-62689-1 (e-book). 469 pp. LCCN Q175 .P455 2014.

Patton:2017:EN

[Pat17] Lydia Patton. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):??, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Patton:2018:M

[Pat18] Lydia Patton. In memoriam. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):??, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Patton:2020:EN

[Pat20] Lydia Patton. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):??, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Patton:2021:ENa

[Pat21a] Lydia Patton. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):??, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Patton:2021:ENb

[Pat21b] Lydia Patton. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):??, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Patton:2021:ENc

[Pat21c] Lydia Patton. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):??, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Patton:2024:EN

[Pat24] Lydia Patton. Editor’s note. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):??, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pearce:2014:BRB

[Pea14] Trevor Pearce. Book review: Cheryl Misak, The American Pragmatists. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):172–176, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/674978.

Pearce:2018:PFR

[Pea18] Trevor Pearce. “Protoplasm feels”: The role of physiology in Charles Sanders Peirce’s evolutionary metaphysics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1): 28–61, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pearce:2019:BRM

[Pea19a] Trevor Pearce. Book review: Michel Bourdeau, Mary Pickering, and Warren Schmaus, eds., Love, Order, and Progress: The Science, Philosophy, and Politics of Auguste Comte. Pittsburgh: University of Pittsburgh Press, 2018. Pp. xiv + 402. $49.95 (cloth). ISBN 0-8229-4522-3. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):419–423, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pearson:2019:BRJ

[Pea19b] James Pearson. Book review: James McElvenny, Language and Meaning in the Age of Modernism: C. K. Ogden and His Contemporaries. Edinburgh: Edinburgh University Press, 2018. Pp. 200. $110.00 (cloth). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):193–198, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pecere:2021:SSW

[Pec21] Paolo Pecere. “Stahl was often closer to the truth”: Kant’s second thoughts on animism, monadology, and hylozoism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):660–678, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pence:2013:BRS

[Pen13] Charles H. Pence. Book review: Staffan Müller-Wille and Hans-Jörg Rheinberger, A Cultural History of Heredity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):168–172, Spring 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/668913.

Pence:2017:BRR

[Pen17] Charles H. Pence. Book review: Robert J. Richards and Michael Ruse, Debating Darwin. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):397–400, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pence:2018:SJF

[Pen18] Charles H. Pence. Sir John F. W. Herschel and Charles Darwin: Nineteenth-century science and its methodology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):108–140, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Petit:2016:SAC

[Pet16] Annie Petit. Le système d’Auguste Comte. Librairie Philosophique J. Vrin, Paris, France, 2016. ISBN 2-7116-2699-7. 389 pp.

Peterman:2018:ERC

[Pet18] Alison Peterman. Essay review: Canonizing Cavendish. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):191–197, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Psillos:2019:PMA

[PG19] Stathis Psillos and Eirini Goudarouli. Principles of motion and the absence of laws of nature in Hobbes’s natural philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):93–119, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pickave:2016:BRN

[Pic16] Martin Pickavé. Book review: Neil Lewis and Rega Wood, eds., In Aristotelis De generatione et corruptione. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):181–184, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Preston:2008:KBS

[PK08] John Preston and Thomas S. Kuhn. Kuhn’s The structure of scientific revolutions: a reader’s guide. Continuum reader’s guides. Continuum, London, UK, 2008. ISBN 0-8264-9375-0, 0-8264-9376-9. viii + 126 pp. LCCN Q175.K953 P74

2008. URL http://www.loc.gov/catdir/enhancements/fy1116/
2007044434-b.html;   http://www.loc.gov/catdir/
enhancements/fy1213/2007044434-d.html;
http://www.loc.gov/catdir/toc/ecip083/2007044434.html.

Plutynski:2011:FPA

[Plu11] Anya Plutynski. Four problems of abduction: A brief history. HO-POS: Journal of the International Society for the History of Philosophy of Science, 1(2):227–248, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/660746.

Podgorny:2014:BRA

[Pod14] Irina Podgorny. Book review: Alex Levine and Adriana Novoa, ‘!Darwinistas! The Construction of Evolutionary Thought in Nineteenth Century Argentina. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):179–182, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/675435.

Poole:2010:JAA

[Poo10] William Poole. John Aubrey and the advancement of learning. Bodleian Library, Oxford, UK, 2010. ISBN 1-85124-319-4 (paperback). 111 pp. LCCN DA93.A8 P66 2010.

Popa:2014:I

[Pop14a] Tiberiu Popa. Introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):271, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/677565.

Popa:2014:SMM

[Pop14b] Tiberiu Popa. Scientific method in meteorology IV. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):306–334, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/677566.

Popa:2016:BRD

[Pop16] Tiberiu Popa. Book review: David Ebrey, ed., Theory and Practice in Aristotle’s Natural Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):354–357, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Powers:2012:ICH

[Pow12] John C. Powers. Inventing chemistry: Herman Boerhaave and the reform of the chemical arts. Synthesis. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-67760-5 (hardcover), 1-280-12635-3, 0-226-67762-1 (e-book). viii + 260 pp. LCCN QD15 .P69 2012.

Pravica:2013:BRH

[Pra13] Sandra Pravica. Book review: Hans-Jörg Rheinberger, On Historicizing Epistemology: An Essay. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):345–348, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/670131.

Preston:2017:API

[Pre17] Aaron Preston, editor. Analytic Philosophy: an Interpretive History. Routledge, Taylor and Francis Group, New York, NY, USA, 2017. ISBN 1-138-80078-3 (hardcover), 1-138-80079-1 (paperback), 1-315-73305-6. viii + 289 pp. LCCN B808.5 .A523 2017.

Potters:2023:WNB

[PS23] Jan Potters and Massimiliano Simons. We have never been “New Experimentalists”: On the rise and fall of the turn to experimentation in the 1980s. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):91–119, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Purnell:2015:BRH

[Pur15] Carolyn Purnell. Book review: Henry Martyn Lloyd, ed., The Discourse of Sensibility: The Knowing Body in the Enlightenment. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):198–201, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/680377.

Radler:2013:NCD

[Rad13] Jan Radler. Neurath’s congestions, depth of intention, and precization: Arne Naess and his Viennese heritage. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):59–90, Spring 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/667986.

Raghuramaraju:2019:BRJ

[Rag19] A. Raghuramaraju. Book review: Jobin M. Kanjirakkat, Gordon McOuat, and Sundar Sarukkai, eds., Science and Narratives of Nature: East and West. London: Routledge, 2015. Pp. xi + 337. $175.00 (cloth); $51.95 (paper). ISBN 978-1-138-90089-9 (cloth); ISBN: 978-0-815-37349-0 (paper). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(2):407– 410, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Raposo:2020:BRO

[Rap20] Pedro M. P. Raposo. Book review: Omar W. Nasim, Observing by Hand: Sketching the Nebulae in the Nineteenth Century. Chicago:

University of Chicago Press, 2013. Pp. 296. $48.00 (cloth). ISBN 978-0-226-08437-4. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):347–349, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rates:2024:BRT

[Rat24] Bruno Rates. Book review: Tano S. Posteraro, Bergson’s Philosophy of Biology: Virtuality, Tendency, and Time. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):606–610, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rayman:2014:CED

[Ray14] Joshua Rayman. Crossing the epistemological divide: Foucault, Barthes, and neo-Kantianism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):217–240, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/677036.

Richards:2016:KSS

[RD16] Robert J. (Robert John) Richards and Lorraine Daston, editors. Kuhn’s Structure of Scientific Revolutions at fifty: reflections on a science classic. University of Chicago Press, Chicago, IL, USA, 2016. ISBN 0-226-31703-X (hardcover), 0-226-31720-X (paperback), 0-226-31717-X (e-book). 202 pp. LCCN Q175 .K84 2016.

Redding:2024:HAA

[Red24] Paul Redding. Hegel’s (anticipated) answer to Peirce’s stalled critique of Cantor’s analytic continuum. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2): 479–507, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Reichenbach:2008:CPM

[REG08] Hans Reichenbach, Frederick Eberhardt, and Clark N. Glymour, editors. The concept of probability in the mathematical representation of reality, volume 3 of Full circle. Open Court, Chicago, IL, USA, 2008. ISBN 0-8126-9609-3 (paperback). xi + 154 pp. LCCN B2779

.R4513 2008. URL http://www.loc.gov/catdir/enhancements/
fy0738/2007021145-d.html;
http://www.loc.gov/catdir/toc/ecip0718/2007021145.html.

Regier:2021:HMG

[Reg21] Jonathan Regier. A hot mess: Girolamo Cardano, the Inquisition, and the soul. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):547–563, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Reichenbach:2006:CPM

[Rei06] Hans Reichenbach. The concept of probability in the mathematical representation of reality. Open Court, Chicago, IL, USA, 2006. ISBN 0-8126-9609-3 (paperback). xi + 154 pp. LCCN B2779 .R45 2007.

Reisch:2014:BRP

[Rei14a] George Reisch. Book review: Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca Lemov, Thomas Sturm, and Michael D. Gordin, How Reason Almost Lost Its Mind: The Strange Career of Cold War Rationality. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):358–361, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/677213.

Reisch:2014:BRB

[Rei14b] George A. Reisch. Book review: Andrew Jewett, Science, Democracy, and the American University: From the Civil War to the Cold War. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):150–153, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/675399.

Reiners:2020:OSM

[Rei20] Stefan Reiners. “Our science must establish itself”: On the scientific status of Lazarus and Steinthal’s Völkerpsychologie. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):234–253, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Remes:2013:BRB

[Rem13] Pauliina Remes. Book review: Lloyd P. Gerson, ed., The Cambridge History of Philosophy in Late Antiquity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3 (2):376–380, Fall 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/670130.

Rendl:2018:BRR

[Ren18] Lois Marie Rendl. Book review: Rudolf Carnap, L’espace: Une contribution à la théorie de la science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8 (1):236–240, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rusu:2024:BRG

[RES^{+}24] Doina-Cristina Rusu, Mehmet Elgin, Elliott Sober, Thomas Uebel, and Daniel Schwartz. Book review: Gideon Manning, ed., Testimonies: States of Mind and States of Body in the Early Modern Period. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):654–657, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Reynolds:2015:BRL

[Rey15] Sarah Jozina Reynolds. Book review: Lydia Patton, ed., Philosophy, Science, and History: A Guide and Reader. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):352–356, Fall 2015. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/682715.

Rheinberger:2010:HEE

[RF10] Hans-Jörg Rheinberger and David Fernbach. On historicizing epistemology: an essay. Cultural memory in the present. Stanford University Press, Stanford, CA, USA, 2010. ISBN 0-8047-6288-0 (hardcover), 0-8047-6289-9 (paperback), 0-8047-7420-X (e-book). 114 pp. LCCN Q175.32.K45 R48513 2010.

Ritter:2010:KTJ

[RH10] Johann Wilhelm Ritter and Jocelyn Holland. Key texts of Johann Wilhelm Ritter (1776–1810) on the science and art of nature, volume 16(13) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-18367-1 (hardcover). xiv + 713 pp. LCCN Q113 .R47213 2010.

Rheinberger:2010:ECT

[Rhe10] Hans-Jorg Rheinberger. An Epistemology of the Concrete: Twentieth-Century Histories of Life. Duke University Press, Durham, NC, USA, 2010. ISBN 0-8223-4575-7. xix + 330 pp. LCCN QH331 .R4513 2010.

Rheinberger:2024:RFJ

[Rhe24] Hans-Jörg Rheinberger. Remarks on François Jacob’s concept of integron. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):483–491, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Richards:2008:TSL

[Ric08] Robert J. Richards. The tragic sense of life: Ernst Haeckel and the struggle over evolutionary thought. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-71214-1 (hardcover), 0-226-71216-8 (paper). xx + 551 + 8 pp. LCCN QH31.H2 R53 2008.

Richards:2010:SPP

[Ric10] Richard A. Richards. The species problem: a philosophical analysis. Cambridge studies in philosophy and biology. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19683-3 (hardcover). x + 236 pp. LCCN QH83 .R53 2010.

Richardson:2012:OEH

[Ric12] Alan Richardson. Occasions for an empirical history of philosophy of science: American philosophers of science at work in the 1950s and 1960s. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):1–20, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/664596.

Richardson:2013:BRT

[Ric13] Alan Richardson. Book review: Thomas S. Kuhn, ed., The Structure of Scientific Revolutions. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):151–154, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/668208.

Richardson:2015:BRN

[Ric15] Alan Richardson. Book review: Nikolay Milkov and Volker Peckhaus, eds., The Berlin Group and the Philosophy of Logical Empiricism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):174–177, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/680434.

Righini:1966:ADC

[Rig66] Guglielmo Righini, editor. Atti del convegno sulle macchie solari = Proceedings of the Meeting on Sunspots, volume 2(2) of Pubblicazioni del Comitato nazionale per le manifestazioni celebrativo del IV centenario della nascita di Galileo Galilei. G. Barbèra, Firenze, Italia, 1966. ISSN 1564-1964. LCCN QB525 .M43 1964.

Rocke:2010:IRK

[Roc10] Alan J. Rocke. Image and reality: Kekulé, Kopp, and the scientific imagination. Synthesis. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-72332-1. xxvi + 375 pp. LCCN QD248 .R63

2010. URL http://www.loc.gov/catdir/enhancements/fy1012/
2009025749-b.html;   http://www.loc.gov/catdir/
enhancements/fy1012/2009025749-d.html;   http://www.
loc.gov/catdir/enhancements/fy1012/2009025749-t.html.

RodriguezGonzalez:2024:BRM

[Rod24] Mariano Rodríguez González. Book review: Mattia Riccardi, Nietzsche’s Philosophical Psychology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):229– 232, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rogers:2015:BRP

[Rog15] Brian Rogers. Book review: Penelope Maddy, The Logical Must: Wittgenstein on Logic. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):360–364, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/682714.

Romizi:2012:VCS

[Rom12] Donata Romizi. The Vienna Circle’s “scientific world–conception”: Philosophy of science in the political arena. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(2):205–242, Fall 2012. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/666659.

Romagni:2024:MKM

[Rom24a] Domenica Romagni. “To measure by a known measure“: Kepler’s geometrical epistemology in the Harmonices Mundi libri V. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(1):103–133, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Romizi:2024:EZP

[Rom24b] Donata Romizi. Edgar Zilsel’s politically engaged philosophy of science (1916–1932). HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):323–360, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Roques:2019:BRH

[Roq19] Magali Roques. Book review: Henrik Lagerlund and Benjamin Hill, eds., The Routledge Companion to Sixteenth Century Philosophy. New York: Routledge, 2017. Pp. 660. $250.00 (cloth); $58.00 (ebook). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):186–189, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rosenkranz:2013:BRS

[Ros13] Ze’ev Rosenkranz. Book review: Steven Gimbel, Einstein’s Jewish Science: Physics at the Intersection of Politics and Religion. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):160–164, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/668914.

Rossini:2022:NOC

[Ros22] Paolo Rossini. The networked origins of Cartesian philosophy and science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):97–120, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rothman:2017:BRM

[Rot17] Aviva Rothman. Book review: Miguel Á. Granada, Patrick J. Boner, and Dario Tessicini, eds., Unifying Heaven and Earth: Essays in the History of Early Modern Cosmology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):169–173, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rot:2024:BQC

[Rot24] Avraham Rot. Between the quest for certainty and intolerance of uncertainty: Hugo Dingler’s way to the forefront of the Deutsche Physik movement, 1900–1937. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):413–452, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rowbottom:2015:BRM

[Row15] Darrell P. Rowbottom. Book review: Mélanie Frappier, Letitia Meynell, and James R. Brown, eds., Thought Experiments in Science, Philosophy, and the Arts. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):348–352, Fall 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/682375.

Rozemond:2013:BRR

[Roz13] Marleen Rozemond. Book review: Roger Ariew, Descartes among the Scholastics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):186–190, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/667763.

Russell:2007:RHB

[Rus07] Paul Russell. The riddle of Hume’s Treatise: skepticism, naturalism, and irreligion. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-511033-1 (hardcover). xvi + 424 pp.

LCCN B1489 .R87 2007. URL http://catdir.loc.gov/catdir/
enhancements/fy0724/2006052536-b.html;   http://catdir.
loc.gov/catdir/enhancements/fy0724/2006052536-d.html.

Ruse:2009:PAD

[Rus09] Michael Ruse, editor. Philosophy after Darwin: classic and contemporary readings. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13553-3 (hardcover), 0-691-13554-1 (paperback). ix + 580 pp. LCCN BJ71 .P49 2009.

Russo:2019:BRK

[Rus19] Federica Russo. Book review: Kevin C. Elliott and Ted Richards, eds., Exploring Inductive Risk: Case Studies of Values in Science. New York: Oxford University Press, 2017. Pp. xiv + 277. $99.00 (cloth); $40.00 (paper). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):179–182, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rusu:2021:ACE

[Rus21] Doina-Cristina Rusu. Anne Conway’s exceptional vitalism: Material spirits and active matter. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):528–546, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Rutkin:2021:CCR

[Rut21] H. Darrel Rutkin. A cosmological controversy in the Renaissance: Marsilio Ficino’s and Giovanni Pico della Mirandola’s contrasting views on the animation of the heavens. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11 (2):604–620, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ryckman:2013:BRN

[Ryc13] Thomas Ryckman. Book review: Norman Sieroka, Umgebungen: Symbolischer Konstruktivismus im Anschluss an Hermann Weyl und Fritz Medicus. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):164–168, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/669491.

Rynasiewicz:2014:BRA

[Ryn14] Robert Rynasiewicz. Book review: Adrian Bardon, A Brief History of the Philosophy of Time. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):165–168, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/675456.

Salter:2012:BRL

[Sal12] Alan Salter. Book review: Lorraine Daston and Elizabeth Lunbeck, eds., Histories of Scientific Observation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2 (1):196–200, Spring 2012. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/663620.

Sandoz:2016:BQQ

[San16] Raphaël Sandoz. Beyond quantities and qualities: Frege and Jevons on measurement. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):212–238, Fall 2016. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sankey:2022:BRK

[San22] Howard Sankey. Book review: K. Brad Wray, ed., Interpreting Kuhn: Critical Essays. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):311–314, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sarasohn:2010:NPM

[Sar10] Lisa T. Sarasohn. The natural philosophy of Margaret Cavendish: reason and fancy during the scientific revolution, volume 128(2) of The Johns Hopkins University studies in historical and political science. The Johns Hopkins University Press, Baltimore, MD, USA, 2010. ISBN 0-8018-9443-3 (hardcover). ISSN 0075-3904. xi + 25 pp.

LCCN Q130 .S266 2010. URL http://catalogue.bnf.fr/ark:/
12148/cb42478595q.

Sargent:2013:I

[Sar13] Rose-Mary Sargent. Introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):135–136, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/669953.

Sarkar:2016:I

[Sar16] Sahotra Sarkar. Introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):239–241, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sargent:2017:RYN

[Sar17a] Rose-Mary Sargent. Book review: Richard Yeo, Notebooks, English Virtuosi, and Early Modern Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1): 167–169, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sarkar:2017:BRH

[Sar17b] Sahotra Sarkar. Book review: Hans-Joachim Niemann, Karl Popper and the Two New Secrets of Life. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):156– 160, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sauer:2017:BRJ

[Sau17] Tilman Sauer. Book review: Jimena Canales, The Physicist and the Philosopher: Einstein, Bergson, and the Debate That Changed Our Understanding of Time. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):163–167, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sangiacomo:2020:DIM

[SB20] Andrea Sangiacomo and Daan Beers. Divide et Impera: Modeling the relationship between canonical and noncanonical authors in the early modern natural philosophy network. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10 (2):365–413, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Streijger:2010:JBQ

[SBT10] Michiel Streijger, Paul J. J. M. Bakker, and J. M. M. H. Thijssen, editors. John Buridan: Quaestiones super libros De generatione et corruptione Aristotelis: a critical edition with an introduction, volume 17(14) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-18504-6 (hardcover). vii + 269 pp. LCCN Q151.A63 B872 2010.

Schmaltz:2007:DC

[Sch07] Tad M. Schmaltz. Descartes on Causation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-532794-2. xii + 237 pp. LCCN B1878.C3 S26 2007.

Schmid:2010:FFN

[Sch10] Stephan Schmid. Finalursachen in der frühen Neuzeit: Eine Untersuchung der Transformation teleologischer Erklärungen. (German) [Final causes in the early modern period: A study of the transformation of teleological explanations]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 2010. ISBN 3-11-024665-1, 3-11-024666-X. xii + 410 pp. LCCN BD543 .S34 2010.

Schliesser:2011:NCP

[Sch11a] Eric Schliesser. Newton’s challenge to philosophy: A programmatic essay. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):101–128, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/658906.

Schmaus:2011:SSC

[Sch11b] Warren Schmaus. Science and the social contract in Renouvier. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):73–100, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/658370.

Schliesser:2013:BRP

[Sch13] Eric Schliesser. Book review: Paul Russell, The Riddle of Hume’s Treatise: Skepticism, Naturalism, and Irreligion. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):172–175, Spring 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/669546.

Schech:2014:BRC

[Sch14a] Elay Schech. Book review: Chunglin Kwa, Styles of Knowing: A New History of Science from Ancient Times to the Present. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):161–165, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/675384.

Schmaltz:2014:ECH

[Sch14b] Tad M. Schmaltz, editor. Efficient Causation: the History of a Concept. Oxford philosophical concepts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-978218-0 (hardcover), 0-19-978217-2 (paperback). xiv + 372 pp. LCCN BD531.5 .E34 2014.

Schmaltz:2016:EMC

[Sch16] Tad M. Schmaltz. Early Modern Cartesianisms: Dutch and French Constructions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-049522-7 (hardcover). ix + 382 pp. LCCN B1875 .S328 2016.

Schmaus:2017:HPC

[Sch17a] Warren Schmaus. Henri Poincaré and Charles Renouvier on conventions; or, how science is like politics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2): 182–198, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Schwartz:2017:CIF

[Sch17b] Daniel Schwartz. Crucial instances and Francis Bacon’s quest for certainty. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):130–150, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Schickore:2018:LLT

[Sch18] Jutta Schickore. Larry Laudan’s typology for historical methodology and the historical and experimental turns in philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):87–107, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Schmaus:2021:BRS

[Sch21] Warren Schmaus. Book review: Stefano Bordoni, When Historiography Met Epistemology: Sophisticated Histories and Philosophies of Science in French-Speaking Countries in the Second Half of the Nineteenth Century. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):488–492, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Schickore:2024:SDE

[Sch24] Jutta Schickore. “Dare Explanations” (Wagerklärungen): Hypothetical thinking in late eighteenth- and early nineteenth-century German philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):387– 412, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Smith:2021:STM

[SDR^{+}21] Roger Smith, Boris Demarest, Jonathan Regier, Charles Wolfe, Warren Schmaus, Olivier Rey, and Anastasios Brenner. The senses of touch and movement and the argument for active powers. HO-POS: Journal of the International Society for the History of Philosophy of Science, 11(2):679–699, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Seger:2014:BRW

[Seg14] Laura M. Seger. Book review: William H. Krieger, ed., Science at the Frontiers: Perspectives on the History and Philosophy of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):365–367, Fall 2014. CODEN ????

ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/677539.

Sepkoski:2012:RFR

[Sep12] David Sepkoski. Rereading the fossil record: the growth of paleobiology as an evolutionary discipline. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-74855-3 (hardcover), 0-226-74858-8 (e-book). 432 pp. LCCN QE719.8 .S47 2012.

Sepetyi:2024:RCP

[Sep24] Dmytro Sepetyi. Rationalism critical and pancritical: What did Popper and Bartley disagree about? HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2): 572–602, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sereni:2021:BRE

[Ser21] Andrea Sereni. Book review: Erich H. Reck and Georg Schiemer, eds., The Prehistory of Mathematical Structuralism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):412–415, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Stuart:2021:MHP

[SF21] Michael T. Stuart and Yiftach Fehige. Motivating the history of the philosophy of thought experiments. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11 (1):212–221, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sgarbi:2012:ATR

[Sga12] Marco Sgarbi. The Aristotelian tradition and the rise of British empiricism: logic and epistemology in the British Isles (1570–1689), volume 32 of Studies in history and philosophy of science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4950-3, 94-007-4951-1 (e-book). xii + 259 pp. LCCN B1131 .S43 2013.

Sgarbi:2017:WDR

[Sga17] Marco Sgarbi. What does a Renaissance Aristotelian look like? From Petrarch to Galilei. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(2):226–245, Fall 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Shaw:2020:PEB

[Sha20] Jamie Shaw. The problem of the empirical basis in the Popperian tradition: Popper, Bartley, and Feyerabend. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):524–561, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Shaheen:2021:LTS

[Sha21] Jonathan L. Shaheen. The life of the thrice sensitive, rational and wise animate matter: Cavendish’s animism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(2):621–641, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Shea:2011:BRG

[She11] William R. Shea. Book review: Galileo Galilei and Christoph Scheiner, On Sunspots. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):173–175, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/658489.

Sidzinska:2024:CES

[Sid24] Maja Sidzińska. Cogito, Ergo Sumus? the pregnancy problem in Descartes’s philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):209–240, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sieroka:2010:USK

[Sie10] Norman Sieroka. Umgebungen: Symbolischer Konstruktivismus im Anschluss an Hermann Weyl und Fritz Medicus. (German) [Surroundings: Symbolic Constructivism in connection with Hermann Weyl and Fritz Medicus], volume 8 of Legierungen. Chronos, Zürich, Switzerland, 2010. ISBN 3-0340-1006-0 (hardcover). 411 pp. LCCN B809.13 .S54 2010.

Simon:2011:BRB

[Sim11] Jonathan Simon. Book review: Bernadette Bensaude-Vincent and William R. Newman, eds., The Artificial and the Natural: An Evolving Polarity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):333–337, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/660848.

Sinclair:2019:BRP

[Sin19] Robert Sinclair. Book review: Peter Olen and Carl Sachs, eds., Pragmatism in Transition: Contemporary Perspectives on C. I. Lewis. Basingstoke: Palgrave Macmillan, 2017. Pp. viii + 222. $109.99 (cloth); $84.99 (e-book). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):201– 205, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sinclair:2022:BRQ

[Sin22] Robert Sinclair. Book review: Quentin Kammer, Jean-Philippe Narboux, and Henri Wagner, eds., C. I. Lewis: The A Priori and the Given. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):315–319, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Slowik:2011:BRT

[Slo11a] Edward Slowik. Book review: Tad M. Schmaltz, Descartes on Causation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):165–169, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/658854.

Slowik:2011:NPS

[Slo11b] Edward Slowik. Newton, the parts of space, and the holism of spatial ontology. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):249–272, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http:/

/www.jstor.org/stable/10.1086/660923.

Slowik:2024:BRR

[Slo24] Edward Slowik. Book review: Richard T. W. Arthur, Leibniz on Time, Space, and Relativity. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):243–246, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Smeenk:2014:BRJ

[Sme14] Chris Smeenk. Book review: Jed Z. Buchwald and Mordechai Feingold, Newton and the Origin of Civilization. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):383–387, Fall 2014. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/677248.

Smith:2011:BRD

[Smi11a] Justin E. H. Smith. Book review: Daniel Garber, Leibniz: Body, Substance, Monad. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):153–157, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/656672.

Smith:2011:DML

[Smi11b] Justin E. H. Smith. Divine machines: Leibniz and the sciences of life. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14178-9 (hardcover). xii + 380 pp. LCCN Q143.L472 S65 2011.

URL http://www.loc.gov/catdir/toc/fy11pdf04/2010053174.
html.

Smith:2024:BRS

[Smi24] James Andrew Smith. Book review: Sean Morris, ed., The Philosophical Project of Carnap and Quine. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14 (1):260–263, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Snapper:2011:BRW

[Sna11] John W. Snapper. Book review: W. D. Hart, The Evolution of Logic. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):337–340, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/661232.

Southgate:2013:KCL

[Sou13] Henry Michael Southgate. Kant’s critique of Leibniz’s rejection of real opposition. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(1):91–134, Spring 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/667898.

Spink:2018:BRT

[Spi18] Aaron Spink. Book review: Tad M. Schmaltz, Early Modern Cartesianisms: Dutch and French Constructions. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):229–232, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sheinbaum:2018:BRJ

[SR18] Mariana Imaz Sheinbaum and Paul A. Roth. Book review: Jörn Rüsen, Evidence and Meaning: a Theory of Historical Studies. HO-POS: Journal of the International Society for the History of Philosophy of Science, 8(2):489–492, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Schmaus:2021:SII

[SR21] Warren Schmaus and Olivier Rey. Special issue introduction. HO-POS: Journal of the International Society for the History of Philosophy of Science, 11(2):421–427, Fall 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Sorell:2009:SEM

[SRK09] Tom Sorell, G. A. J. (Graham Alan John) Rogers, and Jill Kraye, editors. Scientia in early modern philosophy: Seventeenth-Century thinkers on demonstrative knowledge from first principles, volume 24 of Studies in history and philosophy of science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-3076-X (hardcover). xv + 139 pp. LCCN B801 .S38 2010.

Staley:2016:BRA

[ST16] Kent Staley and Heráclio Tavares. Book review: Allan Franklin, Shifting Standards: Experiments in Particle Physics in the Twentieth Century. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):158–162, Spring 2016. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Stan:2011:BRM

[Sta11] Marius Stan. Book review: Michela Massimi, ed., Kant and Philosophy of Science Today. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):364–367, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/661208.

Stachel:2016:POS

[Sta16] John Stachel. Poincaré and the origins of Special Relativity. HO-POS: Journal of the International Society for the History of Philosophy of Science, 6(2):242–256, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Stengers:2010:C

[Ste10] Isabelle Stengers. Cosmopolitics, volume 1. University of Minnesota Press, Minneapolis, MN, USA, 2010. ISBN 0-8166-5686-X (hardcover), 0-8166-5687-8 (paperback). viii + 299 pp. LCCN Q125 .S742613 2010.

Sterrett:2013:BRD

[Ste13] S. G. Sterrett. Book review: Daniela Bailer-Jones, Scientific Models in Philosophy of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):334–337, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/670128.

Steigerwald:2019:BRJ

[Ste19a] Joan Steigerwald. Book review: John H. Zammito, The Gestation of German Biology: Philosophy and Physiology from Stahl to Schelling: Chicago: University of Chicago Press, 2017. Pp. 523. $45.00 (cloth). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):205–208, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Steizinger:2019:BRA

[Ste19b] Johannes Steizinger. Book review: Amos Morris-Reich and Dirk Rupnow, eds., Ideas of ‘Race’ in the History of the Humanities. Cham: Palgrave Macmillan, 2017. Pp. xiii + 337. $109.00 (cloth); $85.00 (e-book). HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):182–185, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Steizinger:2020:VCA

[Ste20] Johannes Steizinger. From Völkerpsychologie to cultural anthropology: Erich Rothacker’s philosophy of culture. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):308–328, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Stone:2013:EQQ

[Sto13] A. Douglas Stone. Einstein and the quantum: the quest of the valiant Swabian. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-13968-7 (hardcover), 1-4008-4834-2 (ebook), 1-299-87423-1 (e-book). x + 332 pp. LCCN QC16.E5

S76 2013. URL http://www.npr.org/2013/11/01/242356997/
einsteins-real-breakthrough-quantum-theory.

Soler:2015:SIC

[STP15] Lena Soler, Emiliano Trizio, and Andrew Pickering, editors. Science as it could have been: discussing the contingency/inevitability problem. University of Pittsburgh Press, Pittsburgh, PA, USA, 2015. ISBN 0-8229-4445-6 (hardcover). x + 462 pp. LCCN Q125 .S4327 2015.

Strick:2018:DLC

[Str18] James Strick. Book review: Daryn Lehoux, Creatures Born of Mud and Slime: The Wonder and Complexity of Spontaneous Generation. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):492–494, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Stuchlik:2011:FNN

[Stu11a] Joshua Stuchlik. Felicitology: Neurath’s naturalization of ethics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):183–208, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/659399.

Stump:2011:APF

[Stu11b] David J. Stump. Arthur Pap’s functional theory of the A Priori. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):273–290, Fall 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/659342.

Stump:2011:BRM

[Stu11c] David J. Stump. Book review: Michael Heidelberger and Gregor Schiemann, eds., The Significance of the Hypothetical in the Natural Sciences. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):129–132, Spring 2011. CO-DEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL

http://www.jstor.org/stable/10.1086/656510.

Stump:2015:CCP

[Stu15a] David J. Stump. Conceptual change and the philosophy of science: alternative interpretations of the a priori, volume 16 of Routledge studies in the philosophy of science. Routledge, Taylor and Francis Group, New York, NY, USA, 2015. ISBN 1-138-89013-8 (hardcover), 1-315-71258-X (e-book). xviii + 176 pp. LCCN Q175 .S934 2015.

Stump:2015:I

[Stu15b] David J. Stump. Introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):70–71, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/680370.

Stump:2016:LSP

[Stu16] David J. Stump. A life in science, philosophy, and the public domain: Three biographies of Poincaré. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2): 309–318, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Stuart:2021:TSS

[Stu21] Michael T. Stuart. Telling stories in science: Feyerabend and thought experiments. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):262–281, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Tabb:2011:BRJ

[Tab11] Kathryn Tabb. Book review: Jonathan Hodge and Gregory Radick, eds., The Cambridge Companion to Darwin. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(2):355–359, Fall 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/661291.

Tauber:2009:SQM

[Tau09] Alfred I. Tauber. Science and the quest for meaning. Baylor University Press, Waco, TX, USA, 2009. ISBN 1-60258-210-6 (hardcover). xi + 255 pp. LCCN Q175 .T2245 2009.

Teresa:2010:GBE

[Ter10] Casteläao Teresa. Gaston Bachelard et les études critiques de la science. (French) [Gaston Bachelard and critical studies of science].

Ouverture philosophique. Harmattan, Paris, France, 2010. ISBN 2-296-11719-8. 212 pp. LCCN B2430.B254 C37 2010. Preface by Jean-Jacques Wunenburger.

Tester:2016:NKS

[Tes16] Steven Tester. Nature, knowledge, and scientific theories in G. C. Lichtenberg’s Reflections on Physics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2): 185–211, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Teufel:2015:BRJ

[Teu15] Thomas Teufel. Book review: Jennifer Mensch, Kant’s Organicism: Epigenesis and the Development of Critical Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(1):190–194, Spring 2015. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.jstor.org/

stable/10.1086/678186.

Textor:2021:MNM

[Tex21] Mark Textor. Mach’s neutral monism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11 (1):143–165, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Ternullo:2024:PCS

[TF24] Claudio Ternullo and Isabella Fascitiello. Peano’s conception of a single infinite cardinality. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):241–260, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Therrien:2020:ACR

[The20] Valérie Lynn Therrien. The Axiom of Choice and the road paved by Sierpiński. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):504–523, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Timmins:2013:WWK

[Tim13] Adam Timmins. Why was Kuhn’s structure more successful than Polanyi’s personal knowledge? HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):306–317, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/671347.

Timmins:2019:BHP

[Tim19] Adam Timmins. Between history and philosophy of science: The relationship between Kuhn’s Black-Body Theory and Structure. HO-POS: Journal of the International Society for the History of Philosophy of Science, 9(2):371–387, Fall 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Toader:2013:CFS

[Toa13] Iulian D. Toader. Concept formation and scientific objectivity: Weyl’s turn against Husserl. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3(2):281–305, Fall 2013. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/670650.

Toader:2021:WDW

[Toa21] Iulian D. Toader. Why did Weyl think that Emmy Noether made algebra the Eldorado of axiomatics? HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1): 122–142, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Toscano:2009:ADM

[Tos09] Maria Toscano. Gli archivi del mondo: antiquaria, storia naturale e collezionismo nel secondo Settecento. (Italian) [The archives of the world: antiques, natural history, and collecting in the late Eighteenth Century], volume 21 of Le voci del museo. Edifir, Firenze, Italia, 2009. ISBN 88-7970-427-3 (paperback). 350 pp. LCCN AM342 .T67 2009.

Tuboly:2018:ERL

[Tub18] Ádám Tamás Tuboly. Essay review: Logical empiricism in international context. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):198–204, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Tuboly:2020:BRG

[Tub20] Adam Tamas Tuboly. Book review: George A. Reisch, The Politics of Paradigms: Thomas S. Kuhn, James B. Conant, and the Cold War “Struggle for Men’s Minds”. Albany, NY: SUNY Press, 2019. Pp. xlv + 456. $95.00 (cloth). ISBN 978-14-38-47367-3. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):605–608, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Tuboly:2021:ISM

[Tub21] Adam Tamas Tuboly. To the icy slopes in the melting pot: Forging logical empiricisms in the context of American pragmatisms. HO-POS: Journal of the International Society for the History of Philosophy of Science, 11(1):27–71, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Turner:2011:BRU

[Tur11] Stephen Turner. Book review: Uljana Feest, ed., Historical Perspectives on Erklären and Verstehen. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1): 135–139, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-

6240 (electronic). URL http://www.jstor.org/stable/10.1086/
658288.

Turnbull:2024:BRK

[Tur24] Nick Turnbull. Book review: Kyung-Man Kim, Bourdieu’s Philosophy and Sociology of Science: a Critical Appraisal. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2):627–630, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Uebel:2007:ECV

[Ueb07] Thomas E. (Thomas Ernst) Uebel. Empiricism at the crossroads: the Vienna Circle’s protocol-sentence debate, volume 4 of Full circle. Open Court, Chicago, IL, USA, 2007. ISBN 0-8126-9640-9 (paperbackback). xviii + 518 pp. LCCN B824.6 .U325 2007. URL http:

//catdir.loc.gov/catdir/toc/ecip082/2007040959.html.

Uebel:2011:BFC

[Ueb11] Thomas Uebel. Beyond the formalist criterion of cognitive significance: Philipp Frank’s later antimetaphysics. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):47–72, Spring 2011. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/658918.

Uebel:2016:BRC

[Ueb16] Thomas Uebel. Book review: Christoph Limbeck-Lilienau and Friedrich Stadler, Der Wiener Kreis: Texte und Bilder zum Logischen Empirismus. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(2):323–325, Fall 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Uebel:2017:PHH

[Ueb17] Thomas Uebel. Philosophy of history and history of philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 7(1):1–30, Spring 2017. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Uebel:2018:BRK

[Ueb18] Thomas Uebel. Book review: Karl Sigmund, Exact Thinking in Demented Times: The Vienna Circle and Its Epic Quest for the Foundations of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):471–475, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Uebel:2021:BRM

[Ueb21a] Thomas Uebel. Book review: Matt LaVine, Race, Gender and the History of Early Analytic Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1): 324–327, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Uebel:2021:RGN

[Ueb21b] Thomas Uebel. Rejecting the given: Neurath and Carnap on methodological solipsism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):1–26, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Vaesen:2021:FNL

[Vae21] Krist Vaesen. French neopositivism and the logic, psychology, and sociology of scientific discovery. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):183– 200, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

VanDyck:2011:BRH

[Van11] Maarten Van Dyck. Book review: Helen Hattab, Descartes on Forms and Mechanisms. HOPOS: Journal of the International Society for the History of Philosophy of Science, 1(1):157–161, Spring 2011. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/658289.

VandenBerg:2013:KPS

[Van13] Hein Van den Berg. Kant on proper science: biology in the critical philosophy and the opus postumum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-7139-8, 94-007-7140-1. xii + 283 pp. LCCN B2798 .V27 2014; Q175 .V33

2014. URL http://www.loc.gov/catdir/enhancements/fy1318/
2013949296-b.html;   http://www.loc.gov/catdir/
enhancements/fy1318/2013949296-d.html;   http://www.
loc.gov/catdir/enhancements/fy1318/2013949296-t.html.

Vatansever:2018:KRH

[Vat18] Saniye Vatansever. Kant’s response to Hume in the second analogy: a critique of Gerd Buchdahl’s and Michael Friedman’s accounts. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):310–346, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

vanBerkel:2013:IBM

[vB13] Klaas van Berkel. Isaac Beeckman on matter and motion: mechanical philosophy in the making. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0936-4 (paperback), 1-4214-0961-5 (e-book). viii + 265 pp. LCCN Q125.2 .B4713 2013.

vanDongen:2010:EU

[vD10] Jeroen van Dongen. Einstein’s unification. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-88346-6. x + 213 pp.

LCCN QC173.6 .D66 2010. URL http://assets.cambridge.org/
97805218/83467/cover/9780521883467.jpg;   http://www.
loc.gov/catdir/enhancements/fy1009/2010017820-b.html;
http://www.loc.gov/catdir/enhancements/fy1009/
2010017820-d.html;   http://www.loc.gov/catdir/
enhancements/fy1009/2010017820-t.html.

Verhaegh:2020:ARL

[Ver20a] Sander Verhaegh. The American reception of logical positivism: First encounters, 1929–1932. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):106–142, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Verhaegh:2020:BRS

[Ver20b] Sander Verhaegh. Book review: Sean Morris, Quine, New Foundations, and the Philosophy of Set Theory. Cambridge University Press, 2018. Pp. x + 209. $105.00 (cloth). ISBN 978-1-107-15250-2. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):340–343, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Verburgt:2024:SMI

[Ver24] Lukas M. Verburgt. Scientific method, induction, and probability: The Whewell–De Morgan debate on Baconianism, 1830s–1850s. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):134–163, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Vinci:2015:SGK

[Vin15] Thomas C. Vinci. Space, geometry, and Kant’s transcendental deduction of the categories. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-938116-X (hardcover). xii + 251 pp. LCCN B2779 .V56 2015.

Viola:2013:SLE

[Vio13] Enrico Viola. The specificity of logical empiricism in the Twentieth-Century history of scientific philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 3 (2):191–209, Fall 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/670135.

Viola:2024:BRR

[Vio24] Tullio Viola. Book review: Richard Kenneth Atkins, Peirce on Inference: Validity, Strength, and the Community of Inquirers. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(2):644–648, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Virmajoki:2022:DCP

[Vir22] Veli Virmajoki. In defense of causal presentism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):68–96, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Vrahimis:2020:SSP

[Vra20] Andreas Vrahimis. Scientism, social praxis, and overcoming metaphysics: a debate between logical empiricism and the Frankfurt School. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(2):562–597, Fall 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

vanStrien:2015:BRM

[vS15] Marij van Strien. Book review: María de Paz and Robert DiSalle, eds., Poincaré, Philosopher of Science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5 (1):183–187, Spring 2015. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/680380.

vanStrien:2022:VCA

[vS22] Marij van Strien. The Vienna Circle against quantum speculations. HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(2):359–394, Fall 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Wagner:2014:WI

[Wag14] Roy Wagner. Wronski’s infinities. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(1):26–61, Spring 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (elec-

tronic). URL http://www.jstor.org/stable/10.1086/675402.

Ward:2024:WWC

[War24] Zina B. Ward. William Whewell, cluster theorist of kinds. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(2):362–386, ???? 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Watkins:2013:DOH

[Wat13] Eric Watkins, editor. The divine order, the human order, and the order of nature: historical perspectives. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-993440-1, 0-19-993441-X (e-book). xxviii + 240 pp. LCCN B105.O7 D58

2013. URL http://www.oxfordscholarship.com/view/10.1093/
acprof:oso/9780199934409.001.0001/acprof-9780199934409.

Waxman:2013:KAI

[Wax13] Wayne Waxman. Kant’s anatomy of the intelligent mind. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-932831-5, 0-19-936934-8 (e-book). xx + 582 pp. LCCN B2798 .W44 2014.

Webster:2024:GWH

[Web24] Colin Webster. Galen, Writings on Health: Thrasybulus and Health (De sanitate tuenda). HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):239–243, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Weinert:2008:CDF

[Wei08] Friedel Weinert. Copernicus, Darwin, Freud: revolutions in the history and philosophy of science. Blackwell, Oxford, UK, 2008. ISBN 1-4051-8184-2 (hardcover), 1-4051-8183-4 (paperback). xi + 1 + 284 pp. LCCN Q175 .W49 2009.

Wells:2023:SPS

[Wel23] Aaron Wells. Science and the principle of sufficient reason: Du Châtelet contra Wolff. HOPOS: Journal of the International Society for the History of Philosophy of Science, 13(1):24–53, ???? 2023. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Westphal:2015:CRL

[Wes15] Kenneth R. Westphal. Causal realism and the limits of empiricism: Some unexpected insights from Hegel. HOPOS: Journal of the International Society for the History of Philosophy of Science, 5(2):281–317, Fall 2015. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/682374.

West:2024:PSM

[Wes24] Peter West. Philosophy is not a science: Margaret Macdonald on the nature of philosophical theories. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(2): 527–553, Fall 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Wolfe:2012:BRL

[WG12] Charles T. Wolfe and Benjamin Goldberg. Book review: Luuc Kooijmans, Death Defied: The Anatomy Lessons of Frederik Ruysch.

HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):177–182, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/663304.

Williams:1925:EL

[Wil25] Henry Horace Williams. The evolution of logic. [The author], Chapel Hill, NC, USA, 1925. ix–xiv, 181 pp. LCCN BC50 .W5.

Wilson:2008:EOM

[Wil08]  Catherine  Wilson.  Epicure-

anism at the Origins of Modernity. Clarendon Press, New York, NY, USA, 2008. ISBN 0-19-923881-2. x + 304 pp. LCCN B512 .W55

2008. URL http://www.loc.gov/catdir/enhancements/fy0907/
2008015450-b.html;   http://www.loc.gov/catdir/
enhancements/fy0907/2008015450-d.html;   http://www.
loc.gov/catdir/enhancements/fy0907/2008015450-t.html.

Williams:2020:SFF

[Wil20] Jessica J. Williams. “The shape of a four-footed animal in general”: Kant on empirical schemata and the system of nature. HOPOS: Journal of the International Society for the History of Philosophy of Science, 10(1):1–23, Spring 2020. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Wray:2016:IJB

[Wra16] K. Brad Wray. The influence of James B. Conant on Kuhn’s Structure of Scientific Revolutions. HOPOS: Journal of the International Society for the History of Philosophy of Science, 6(1):1–23, Spring 2016. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Wray:2019:KHC

[Wra19] K. Brad Wray. Kuhn, the history of chemistry, and the philosophy of science. HOPOS: Journal of the International Society for the History of Philosophy of Science, 9(1):75–92, Spring 2019. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Wray:2021:KCR

[Wra21] K. Brad Wray. Kuhn and the contemporary realism/antirealism debates. HOPOS: Journal of the International Society for the History of Philosophy of Science, 11(1):72–92, Spring 2021. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Wuest:2013:YTP

[Wue13] Amy Wuest. Yes, there is a problem: What is to be done about the climate for women in philosophy? HOPOS: Journal of the International Society for the History of Philosophy of Science, 3 (1):146–150, Spring 2013. CODEN ???? ISSN 2152-5188 (print),

2156-6240 (electronic). URL http://www.jstor.org/stable/10.
1086/669958.

Wuthrich:2014:BRT

[Wüt14] Christian Wüthrich. Book review: Tian Yu Cao, From Current Algebra to Quantum Chromodynamics: A Case for Structural Realism. HOPOS: Journal of the International Society for the History of Philosophy of Science, 4(2):368–371, Fall 2014. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic). URL http://www.

jstor.org/stable/10.1086/677388.

Yeo:2014:NEV

[Yeo14] Richard R. Yeo. Notebooks, English virtuosi, and early modern science. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-10656-X (hardcover), 0-226-10673-X (e-book). xviii + 398 pp. LCCN Q127.G4 Y47 2014.

Zammito:2018:BRA

[Zam18a] John H. Zammito. Book review: Andrea Gambarotto, Vital Forces, Teleology, and Organization: Philosophy of Nature and the Rise of Biology in Germany. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(2):497–500, Fall 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Zammito:2018:GGB

[Zam18b] John H. Zammito. The gestation of German biology: philosophy and physiology from Stahl to Schelling. University of Chicago Press, Chicago, IL, USA, 2018. ISBN 0-226-52079-X (hardcover), 0-226-52082-X (e-book). 523 pp. LCCN QH320.G3 Z355 2018.

Zamboni:2024:BRR

[Zam24a] Francesco Omar Zamboni. Book review: Riccardo Strobino, Avicenna’s Theory of Science: Logic, Metaphysics, Epistemology. HO-POS: Journal of the International Society for the History of Philosophy of Science, 14(1):263–267, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Zammito:2024:BRC

[Zam24b] John H. Zammito. Book review: Catherine Wilson, Kant and the Naturalistic Turn of 18th Century Philosophy. HOPOS: Journal of the International Society for the History of Philosophy of Science, 14(1):250–253, Spring 2024. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Zupko:2012:BRM

[Zup12] Jack Zupko. Book review: Michiel Streijger, Paul J. J. M. Bakker, and Johannes M. M. H. Thijssen, eds., John Buridan, Quaestiones super libros De generatione et corruption Aristotelis: A Critical Edition with an Introduction. HOPOS: Journal of the International Society for the History of Philosophy of Science, 2(1):192–195, Spring 2012. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

URL http://www.jstor.org/stable/10.1086/663305.